### v2

# Counting rational curves and standard complex structures on hyperKähler ALE 4-manifolds

Abstract.
All hyperKähler ALE 4-manifolds with a given non-trivial finite group $1\not=\Gamma<{\rm SU}(2)$ at infinity are parameterized by an open dense subset of a real linear space of dimension $3{\rm rank}\,\Phi$. Here, $\Phi$ denotes the root system associated with $\Gamma$ via the McKay correspondence. Such manifolds are diffeomorphic to the minimal resolution of a Kleinian singularity. By using the period map of the twistor space, we specify those points in the parameter space at which the hyperKählerian family of complex structures includes the complex structure of the minimal resolution. Furthermore, we count the rational curves lying on each hyperKähler ALE 4-manifold. As an application, we prove that the twistor space of any hyperKähler ALE cannot be Kählerian. In particular, we strengthen some results of Kronheimer (J. Differential Geom., 29(3):665–683, 1989) and provide examples of non-compact and non-Kählerian twistor spaces.
Keywords hyperKähler ALE, Kleinian singularity, twistor space, root system, rational curve.
2000 MR Subject Classification 53C28, 17B22
B.X. is supported in part by the Project of Stable Support for Youth Team in Basic Research Field, CAS (Grant No. YSBR-001) and NSFC (Grant Nos. 12271495 and 12071449).
Y.L. is the corresponding author.
1. introduction
1.1. Background
A hyperKähler structure on a Riemannian manifold $\left(M,g\right)$ is, by definition, a triple $(I,J,K)$ of three parallel complex structures on $M$ that satisfies the quaternion relation $IJ=-JI=K$. The set $\left\{aI+bJ+cK\,|\,(a,b,c)\in\mathbb{S}^{2}\right\}$ is called the hyperKählerian family of complex structures on $(M,g)$. We say a 4-dimensional Riemannian manifold $(M,g)$ is asymptotically locally Euclidean (ALE) if there exist a compact set $K\subset M$, a finite subgroup $1\not=\Gamma<\mathrm{SO}(4)$, and a diffeomorphism $\Psi:M\backslash K\to(\mathbb{R}^{4}\backslash B_{R})/\Gamma$, where $B_{R}$ is the standard ball of radius $R$ in $\mathbb{R}^{4}$, such that for each non-negative integer $k$, there holds
$|\nabla^{k}(\Psi_{*}g-g_{\rm Euc})(x)|=O(|x|^{-4-k})\,,\qquad x\in(\mathbb{R}^{4}\backslash B_{R})/\Gamma\qquad{\rm and}\qquad|x|\to+\infty.$
Kronheimer [8, 9] classified all hyperKähler ALE 4-manifolds. He constructed explicitly for each finite group $1\not=\Gamma<\mathrm{SU}(2)$ a complete family of hyperKähler ALEs with group $\Gamma$ at infinity. We provide a brief overview as follows. Let’s fix a non-trivial finite group $\Gamma<\mathrm{SU}(2)$.
Denote $Q$ as the natural two-dimensional complex representation of $\Gamma$ and $R$ as the regular representation. Define a linear space
$M(\Gamma):=\mathrm{Hom}_{\Gamma}\bigl{(}R,Q\otimes R\bigr{)}$
equipped with a natural hyperKähler structure when we identify $Q=\mathbb{C}^{2}\simeq\mathbb{H}$. This space $M(\Gamma)$ is acted upon by a Lie group $\mathrm{U}(\Gamma)$, which preserves the hyperKähler structure on $M(\Gamma)$. Denote by $Z$ the dual of the center of the Lie algebra $\mathfrak{U}(\Gamma)$ of $\mathrm{U}(\Gamma)$. There exists a natural isomorphism
$\rho:Z\xrightarrow{\sim}\mathfrak{h},$
where $\mathfrak{h}$ is the Cartan subalgebra associated with $\Gamma$, see subsection 2.1 for $\mathfrak{h}$. With respect to the $\mathrm{U}(\Gamma)$-action, we can associate a hyperKähler moment map
$\mu_{\mathbb{H}}:M(\Gamma)\to\mathfrak{U}(\Gamma)^{*}\otimes\mathbb{R}^{3}$
and consider the hyperKähler quotient $X_{\zeta}:=\mu_{\mathbb{H}}^{-1}(\zeta)/\mathrm{U}(\Gamma)$ for each $\zeta\in Z\otimes\mathbb{R}^{3}$. For details on hyperKähler quotient, see [7]. These quotient manifolds may not be smooth. However, for $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$, $X_{\zeta}$ is a smooth ALE 4-manifold equipped with the quotient metric $g_{\zeta}$, where
$(Z\otimes\mathbb{R}^{3})^{\circ}:=(Z\otimes\mathbb{R}^{3})\backslash\bigcup_{\theta\in\Phi}(D_{\theta}\otimes\mathbb{R}^{3})\,,\qquad D_{\theta}:=\mathrm{Ker}(\theta\circ\rho)\subset Z$
and the union is taken over all roots in the root system $\Phi$ associated with $\mathfrak{h}$. By the property of hyperKähler quotient, the smooth quotient Riemannian manifold $\left(X_{\zeta},g_{\zeta}\right)$ admits a natural hyperKähler structure. We denote by $I_{\zeta},J_{\zeta},K_{\zeta}$ the three compatible complex structures on $X_{\zeta}$, and by $\omega_{\zeta}^{k}$, $k\in\{1,2,3\}$ the corresponding Kähler forms. The ordering of $I_{\zeta},J_{\zeta},K_{\zeta}$ is inherited from the ordering of complex structures on the hyperKähler manifold $M(\Gamma)$.
Furthermore, Kronheimer showed that each smooth manifold $X_{\zeta}$ for $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$ is diffeomorphic to the minimal resolution of $\mathbb{C}^{2}/\Gamma$. We denote the minimal resolution of the $\mathbb{C}^{2}/\Gamma$ by $\eta:\widetilde{\mathbb{C}^{2}/\Gamma}\to\mathbb{C}^{2}/\Gamma$ (see Definition 2.1). The set $\eta^{-1}(0)$ is called the exceptional locus, it is a union of $\mathrm{rank}\,\Phi$ rational curves. We call the complex structure on the minimal resolution $\widetilde{\mathbb{C}^{2}/\Gamma}$ the standard complex structure, which we denote by $I_{\mathrm{std}}$.
1.2. Main Results
For $\zeta=(\zeta^{1},0,0)\in(Z\otimes\mathbb{R}^{3})^{\circ}$, Kronheimer [8, Proposition 3.10] showed that the complex manifold $X_{\zeta}$, with respect to the complex structures $I_{\zeta}$, is isomorphic to $\widetilde{\mathbb{C}^{2}/\Gamma}$. Our first main result characterizes all such $\zeta$ as follows. Denote by $\mathrm{rank}(\zeta)$ the dimension of $\mathrm{span}_{\mathbb{R}}\left\{\zeta^{1},\zeta^{2},\zeta^{3}\right\}$. It is clear that $1\leq\mathrm{rank}(\zeta)\leq 3$ for each $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$.
Theorem 1.1.
Let $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$. The hyperKählerian family of complex structures on $(X_{\zeta},g_{\zeta})$ includes $I_{\rm std}$ if and only if ${\rm rank}(\zeta)=1$, i.e. $\zeta$ is of the form $\left(x\zeta^{1},y\zeta^{1},z\zeta^{1}\right)$ for some non-zero vector $(x,y,z)\in\mathbb{R}^{3}\setminus\{0\}$ and $\zeta^{1}\in Z\setminus\bigcup_{\theta\in\Phi}D_{\theta}$. Under this context, the hyperKählerian family of $X_{\zeta}$ includes exactly two complex structures which are equvalent to $I_{\rm std}$, and all the other complex structures are mutually equivalent and isomorphic to a hypersurface in $\mathbb{C}^{3}$.
The twistor space [7, section 3.F] of the hyperKähler manifold $\left(X_{\zeta},g_{\zeta}\right)$ is a complex threefold, denoted by $\mathrm{Z}(X_{\zeta})$, equipped with a holomorphic projection $\pi:\mathrm{Z}(X_{\zeta})\to\mathbb{P}^{1}$. For $u\in\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$ we define a complex structure $I_{\zeta}^{u}:=aI_{\zeta}+bJ_{\zeta}+cK_{\zeta}$, where
[…]
Note equation (1.1) provides an identification between $\mathbb{P}^{1}$ and $\mathbb{S}^{2}$. The fiber $\pi^{-1}(u)$ over $u$ is a complex submanifold of $\mathrm{Z}(X_{\zeta})$ that is isomorphic to the complex manifold $X_{\zeta}$ with respect to the complex structure $I_{\zeta}^{u}$. Our next result is:
Theorem 1.2.
Fix a non-trivial finite subgroup $\Gamma$ in $\mathrm{SU(2)}$. Recall the linear space $Z$ and root system $\Phi$ associated with $\Gamma$. Given $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$, we could count rational curves fiberwise lying in the twistor space $\pi:\mathrm{Z}(X_{\zeta})\to\mathbb{P}^{1}$ as follows:
  • (i)
    If $\mathrm{rank(\zeta)}=1$, then there hold
    (1.2) $\sharp\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}=2,$
    and
    (1.3) $\sum_{u\in\mathbb{P}^{1}}\sharp\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}(u)\right\}=2r,$
    where $r:=\mathrm{rank}\,\Phi$.
  • (ii)
    If $\mathrm{rank(\zeta)}=2$, then there hold
    (1.4) $4\leq\sharp\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}\leq\sharp\Phi,$
    and
    (1.5) $2r\leq\sum_{u\in\mathbb{P}^{1}}\sharp\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}(u)\right\}\leq\sharp\Phi.$
  • (iii)
    If $\mathrm{rank(\zeta)}=3$, then there hold
    (1.6) $3\leq\sharp\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}\leq\sharp\Phi,$
    and
    (1.7) $2r-1\leq\sum_{u\in\mathbb{P}^{1}}\sharp\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}(u)\right\}\leq\sharp\Phi.$
Remark 1.3.
The upper bound of (1.4) and (1.6) is obtained by Kronheimer [9, pp.693]. All the lower bounds are new results. However, it is hard to find sharp lower bounds. The difficulty comes from combinatorial problem which would be demonstrate in section 4. We plan to do some computer calculations and discuss the sharpness in our next paper.
By considering all $\zeta\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$, we can define a map as:
$Q_{1}:\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$ $\to\mathbb{Z}$
$\zeta$$\mapsto\sharp\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}$
We proved:
Theorem 1.4.
$Q_{1}$ is lower semi-continuous, i.e. there holds
$\liminf_{\zeta\to\zeta_{0}}Q_{1}(\zeta)\geq Q_{1}(\zeta_{0})\,\,,\,\,\,\forall\zeta_{0}\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}.$
Remark 1.5.
However, the semi-continuity of the number of rational curves is hard to figure out since the lower bound of this number in case (iii) of Theorem 1.2 is $2r-1<2r$.
As a corollary of both Theorem 1.1 and Theorem 1.2, we can generalize some results of Hitchin [6] about the non-compact non-Kählerian twistor spaces:
Corollary 1.6.
For any finite group $1\not=\Gamma<\mathrm{SU(2)}$ and any $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$, the twistor space $\mathrm{Z}(X_{\zeta})$ is not Kählerian.
1.3. Outline
In Section 2, we recall some basic facts about Kleinian singularities and the structure of their semi-universal deformations. The basic tool in our proofs is the period map of the twistor space of hyperKähler ALE, as discussed in subsection 2.3 and subsection 2.4. In Section 3, we prove Theorem 1.1 and Theorem 1.2 in Case (i). In Section 4, we prove the remainder of Theorem 1.2. In fact, we abstract the Theorem 1.2 to problems about certain types of decompositions of an irreducible root system, which also represent new questions in the study of root systems and Lie algebras. In Section 5, we prove the lower semi-continuity of $Q_{1}$. In the final section, we prove Corollary 1.6 based on known results of Hitchin [6].
1.4. Some Basic Notations
  • $\left(X_{\zeta},I_{\zeta}^{u}\right)$: The complex manifold with underlying differential manifold $X_{\zeta}$ and complex structure $I_{\zeta}^{u}$.
  • $\mathfrak{h}^{\mathbb{C}}$: The complexification of the real Cartan subalgebra $\mathfrak{h}$.
  • $\mathbb{P}(V)$ or $\mathbb{P}V$: The projective space associated with the complex vector space $V$.
  • $\sharp P$: The cardinality of the set $P$.
  • $P_{1}\sqcup P_{2}$: The disjoint union of two sets $P_{1}$ and $P_{2}$.
  • $\mathrm{i}$: The imaginary unit, $\sqrt{-1}$.
2. The Kleinian singularities
There are some important relations between hyperKähler ALE and Kleinian singularities proved by Kronheimer [8, 9]. In this section, we recall some basic facts about the Kleinian singularities and make some preparation for the proofs of our main theorems. In subsection 2.1, we discuss the correspondence between Kleinian singularities and root systems of type ADE. In subsection 2.2, we summarize known results about the structures of the semi-universal deformations of Kleinian singularities and their simultaneous resolutions. In subsection 2.3, we will recall the relations between hyperKähler ALE and Kleinian singularity, e.g., the twistor space of a hyperKähler ALE can be seen as a simultaneous resolution of a deformation of singularity, see details in Proposition 2.12. We also discuss the properties about the period map of the twistor space, which is a basic tool in our proofs, see Lemma 2.13 and its corollary. In subsection 2.4, based on Kronheimer’s results, we establish an isomorphism (Lemma 2.16) between hyperKähler ALEs in the parameter space $Z\otimes\mathbb{R}^{3}$, which is the key of our proofs of Theorem 1.1 in Section 3.
Let $\Gamma$ be a non-trivial finite subgroup of $\mathrm{SU}(2)$, the singular complex surface $\mathbb{C}^{2}/\Gamma$ is called a Kleinian singularity .
2.1. Kleinian Singularities and Root Systmes
Definition 2.1.
[1, Chapter III] A minimal resolution of a Kleinian singularity $\mathbb{C}^{2}/\Gamma$ is a smooth complex surface $\widetilde{\mathbb{C}^{2}/\Gamma}$ together with a proper holomorphic map $\eta:\widetilde{\mathbb{C}^{2}/\Gamma}\to\mathbb{C}^{2}/\Gamma$ such that:
  • The restriction, $\eta:\widetilde{\mathbb{C}^{2}/\Gamma}\setminus\eta^{-1}(0)\to\mathbb{C}^{2}/\Gamma\setminus{0}$, is an isomorphism,
  • $\widetilde{\mathbb{C}^{2}/\Gamma}$ contains no $(-1)$-curves.
The set $\eta^{-1}(0)$ is called the exceptional locus of $\widetilde{\mathbb{C}^{2}/\Gamma}$.
Any Kleinian singularity admits a unique minimal resolution, denoted by $\eta:\widetilde{\mathbb{C}^{2}/\Gamma}\to\mathbb{C}^{2}/\Gamma$. The exceptional locus is a union of rational curves $\eta^{-1}(0)=\Sigma_{1}\cup\cdots\cup\Sigma_{r}$, where each $\Sigma_{i}\simeq\mathbb{P}^{1}$. Topologically, these two-spheres $\Sigma_{i}$, for $i=1,\cdots,r$, generate the second homology group $H_{2}(\widetilde{\mathbb{C}^{2}/\Gamma},\mathbb{Z})$. We associate a vertex with each $\Sigma_{i}$, and connect two vertices with an edge if the corresponding curves intersect. Thus, for each group $\Gamma$, we obtain a dual diagram $\Delta(\Gamma)$ by resolving $\mathbb{C}^{2}/\Gamma$. It is well known that $\Delta(\Gamma)$ is a Dynkin diagram of type ADE. Then associating $\Gamma$ the diagram $\Delta(\Gamma)$ induces a bijection between sets[1] :
[…]
$\Gamma\,\mapsto$ $\,\,\Delta(\Gamma)$
Thus we can say the Kleinian singularity $\mathbb{C}^{2}/\Gamma$ is of Type A, D or E if the Dynkin diagram $\Delta(\Gamma)$ is of Type A, D or E.
Denote by $\mathfrak{h}$ the real Cartan algebra associated with the Dynkin diagram $\Delta(\Gamma)$, and by $\Phi$ the associated root system of $\mathfrak{h}$. The vertices of $\Delta(\Gamma)$ correspond to simple roots $\theta_{1},\cdots,\theta_{r}\,\in\mathfrak{h}^{*}$. We denote the set of simple roots by $\Delta$, which forms a root base of the root system $\Phi$. Returning to the minimal resolution $\widetilde{\mathbb{C}^{2}/\Gamma}$, the two-spheres $\Sigma_{i}$, for $i=1,\cdots,r$, correspond to the simple roots $\theta_{i}$, for $i=1,\cdots,r$. The fact that these two-spheres generate $H_{2}(\widetilde{\mathbb{C}^{2}/\Gamma},\mathbb{Z})$ provides an embedding of the root system $\Phi\hookrightarrow H_{2}(\widetilde{\mathbb{C}^{2}/\Gamma},\mathbb{Z})$. Hence, there are natural isomorphisms:
$H_{2}(\widetilde{\mathbb{C}^{2}/\Gamma},\mathbb{R})\simeq\mathfrak{h}^{*}\,,\,H^{2}(\widetilde{\mathbb{C}^{2}/\Gamma},\mathbb{C})\simeq\mathfrak{h}^{\mathbb{C}}.$
The Kleinian singularity $\mathbb{C}^{2}/\Gamma$ is said to be of type $\Phi$ at the singular point 00. It is inherently local.
2.2. Deformations of Kleinian Singularities
Definition 2.2.
[4, pp.7] A complex space is a topological space $X$ equipped with a structure sheaf $\mathcal{O}_{X}$, which satisfies the following properties:
  • (i)
    every point $x\in X$ has a neighbourhhod $U$ such that $U=D\cap Z$ where $D$ is a domain in $\mathbb{C}^{n}$ for some integer $n$ and $Z$ is the zero set of some holomorphic functions $f_{1},\cdots,f_{k}$ on $D$.
  • (ii)
    And the restriction of the structure sheaf $\mathcal{O}_{X}$ to $U$, denoted by $\mathcal{O}_{X}(U)$, is isomorphic to the restriction of the quotient sheaf $\mathcal{O}_{D}/I$ to $U$, where $I$ is the ideal generated by $f_{1},\cdots,f_{k}$.
Morphisms between complex spaces are called holomorphic maps.
Definition 2.3.
[11] A deformation of a complex space $X^{0}$ is a flat holomorphic map $\pi:\mathcal{X}\to T$ between two complex spaces $\mathcal{X}$ and $T$, together with an isomorphism $\xi:X^{0}\xrightarrow{\sim}\pi^{-1}(t_{0})$, where $t_{0}$ is a point in $T$. We call $t_{0}$ the distinguished point. This data is expressed by the following commutative diagram:
} ${\left\{t_{0}\right\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Here $\pi$ is a flat holomorphic map, and $\xi$ is an embedding. We say a deformation is a $\mathbb{C}^{*}$-deformation if the above diagram admits a $\mathbb{C}^{*}$-equivariant action. This means there exist respectively $\mathbb{C}^{*}$-actions on complex spaces $X^{0},\mathcal{X}$ and $T$ such that $\left\{t_{0}\right\}$ is a fixed point in $T$ and the morphisms $\xi$ and $\pi$ are equivariant.
Lemma 2.4.
[8, Section 4] There exists a $\mathbb{C}^{*}$-deformation of the Kleinian singularity $\mathbb{C}^{2}/\Gamma$, denoted by $f:\mathcal{Y}\to\mathcal{V}$, which satisfies:
  • (1)
    $\mathcal{V}$ can be identified with $\mathfrak{h}^{\mathbb{C}}/W$, where $\mathfrak{h}^{\mathbb{C}}$ is the complexification of $\mathfrak{h}$. The distinguished point is $0\in\mathfrak{h}^{\mathbb{C}}/W$, meaning $f^{-1}(0)\simeq\mathbb{C}^{2}/\Gamma$.
  • (2)
    The branch locus $\mathcal{D}\subset\mathcal{V}$ is the projection of the kernels of the roots in $\Phi$ under the quotient map $\mathfrak{h}^{\mathbb{C}}\to\mathfrak{h}^{\mathbb{C}}/W$.
  • (3)
    $\mathcal{D}$ is the set of $v\in\mathcal{V}$ for which the fiber $f^{-1}(v)$ is singular.
We call the deformation $f:\mathcal{Y}\to\mathcal{V}$ the semi-universal deformation of $\mathbb{C}^{2}/\Gamma$.
Lemma 2.5.
[10, 11] By the ADE classification of Kleinian singularities stated in subsection 2.1, the space $\mathcal{Y}$ in the semi-universal deformation $f:\mathcal{Y}\to\mathcal{V}$ is a hypersurface in a vector space $\mathbb{C}^{3}\crossproduct\mathbb{C}^{r}$ given by one of the following equations:
$\mathrm{Type}\,A_{r}:\,\,$ $xy+z^{r+1}+t_{1}z^{r-1}+t_{2}z^{r-2}+\cdots+t_{r-1}z+t_{r}=0;$
$\mathrm{Type}\,D_{r}:\,\,$ $x^{2}+y^{2}z+z^{r-1}+t_{1}z^{r-2}+\cdots+t_{r-1}+2t_{r}y=0;$
$\mathrm{Type}\,E_{6}:\,\,$ $x^{2}+y^{3}+z^{4}+t_{1}z^{2}+t_{2}z+t_{3}z+y(t_{4}z^{2}+t_{5}z+t_{6})=0;$
$\mathrm{Type}\,E_{7}:\,\,$ $x^{2}+y^{3}+yz^{3}+y(t_{1}z+t_{2})+t_{3}z^{4}+\cdots+t_{6}z+t_{7}=0;$
$\mathrm{Type}\,E_{8}:\,\,$ $x^{2}+y^{3}+z^{5}+y(t_{1}z^{3}+\cdots+t_{4})+t_{5}z^{3}+\cdots+t_{8}=0,$
and the projection $f$ is simply the projection to the parameter space $\{(t_{1},\cdots,t_{r})\in\mathbb{C}^{r}\}\simeq\mathcal{V}$. Thus, each fiber of $f$ is a complex surface (possibly singular) in the affine space $\mathbb{C}^{3}$.
We denote by $\tilde{f}:\tilde{\mathcal{Y}}\to\mathfrak{h}^{\mathbb{C}}$ the deformation which is the pullback of the semi-universal deformation by the natural quotient map $q:\mathfrak{h}^{\mathbb{C}}\to\mathfrak{h}^{\mathbb{C}}/W\simeq\mathcal{V}$, i.e.,
(2.1) ${\tilde{\mathcal{Y}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{f}}$${\mathcal{Y}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${\mathfrak{h}^{\mathbb{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{q}$${\mathcal{V}.}$
Then, we have the explicit description about the singularities in the singular fibers of $\tilde{f}$:
Lemma 2.6.
[10, Lemma 1, p.94] For any $\lambda\in\mathfrak{h}^{\mathbb{C}}$, denote $\Phi_{\lambda}:=\{\theta\in\Phi\,|\,\theta(\lambda)=0\}$, which is a root subsystem of the root system $\Phi$. We have:
  • (1)
    The fiber $\tilde{\mathcal{Y}}_{\lambda}$ has singularities if and only if $\lambda\in\bigcup_{\theta\in\Phi}\mathrm{Ker}\theta$.
  • (2)
    The fiber $\tilde{\mathcal{Y}}_{\lambda}$ has $m$ Kleinian singularities of type $\Phi_{\lambda,1},\cdots,\Phi_{\lambda,m}$ at $m$ points respectively if and only if $\Phi_{\lambda}=\Phi_{\lambda,1}\sqcup\cdots\sqcup\Phi_{\lambda,m}$ is the orthogonal decomposition of $\Phi_{\lambda}$ into irreducible root subsystems.
As a consequence, any singular fiber $\tilde{\mathcal{Y}}_{\lambda}$ is a singular complex surface that has only finite isolated singularities.
Definition 2.7.
[1, Section 6. Chapter III] A minimal resolution of a singular complex surface $X$ with only isolated singularities at a finite set $S$ is a smooth complex surface $\widetilde{X}$ together with a proper holomorphic map $\eta:\widetilde{X}\to X$ such that:
  • The restriction to the smooth part, $\eta:\widetilde{X}\setminus\eta^{-1}(S)\to X\setminus S$, is an isomorphism.
  • $\tilde{X}$ contains no $(-1)$-curves.
And we call the pre-image of singularites $\eta^{-1}(S)$ the exceptional locus of $\widetilde{X}$.
Lemma 2.8.
[1, Chapter III] Every normal surface with finitely many singularities admits a unique minimal resolution. In particular, each fiber $\mathcal{Y}_{q\left(\lambda\right)}\simeq\tilde{\mathcal{Y}}_{\lambda}$ of is normal and admits a unique minimal resolution.
Lemma 2.9.
[8, 10] There exists a commutative diagram of holomorphic maps:
${\mathfrak{Y}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$$\scriptstyle{\hat{f}}$${\mathcal{Y}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${\mathfrak{h}^{\mathbb{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{q}$${\mathcal{V}.}$
such that:
  • $\hat{f}$ is a holomorphic map between smooth complex manifolds.
  • The restriction of $\psi$ to each fiber $\mathfrak{Y}_{\lambda}$ of $\hat{f}$ is the minimal resolution of the fiber $\mathcal{Y}_{q(\lambda)}$ of $f$, where $\lambda\in\mathfrak{h}^{\mathbb{C}}$.
We call the map $\hat{f}$ is a simultaneous resolution of the semi-universal deformation $f$.
Lemma 2.10.
Any rational curve in the fiber $\mathfrak{Y}_{\lambda}$ of $\hat{f}$, $\lambda\in\mathfrak{h}^{\mathbb{C}}$, must lie in the exception locus of $\mathfrak{Y}_{\lambda}$.
Proof.
By Lemma 2.5, $\mathcal{Y}_{q\left(\lambda\right)}$ is a hypersurface in $\mathbb{C}^{3}$. Since any holomorphic map $\mathbb{P}^{1}\hookrightarrow\mathcal{Y}_{q\left(\lambda\right)}\subset\mathbb{C}^{3}$ must be constant, $\mathcal{Y}_{q\left(\lambda\right)}$ contains no rational curves. Suppose a rational curve in $\mathfrak{Y}_{\lambda}$ is given by a holomorphic map $\mathbb{P}^{1}\hookrightarrow\mathfrak{Y}_{\lambda}$, then its composition with the resolution $\mathfrak{Y}_{\lambda}\to\mathcal{Y}_{q\left(\lambda\right)}\subset\mathbb{C}^{3}$ is a constant map. Hence by the definition of minimal resolution, the rational curve must lie in the exceptional locus of $\mathfrak{Y}_{\lambda}$. ∎
Corollary 2.11.
For any $\lambda\in\mathfrak{h}^{\mathbb{C}}$, recall that $\Phi_{\lambda}=\Phi_{1,\lambda}\sqcup\cdots\sqcup\Phi_{m,\lambda}$ represents the type of singularities in Lemma 2.6. Then we have the counting of rational curves in $\mathfrak{Y}_{\lambda}$:
$\sharp\left\{\mathrm{rational\,\,curves\,\,in\,\,}\mathfrak{Y}_{\lambda}\right\}=\mathrm{rank}\,\Phi_{\lambda}.$
Proof.
By Lemma 2.10, the number of rational curves in $\mathfrak{Y}_{\lambda}$ is the sum of the numbers of rational curves in the exceptional locus of each singularity, hence is given by the sum of the ranks of the types $\Phi_{1,\lambda},\cdots,\Phi_{m,\lambda}$ of these singularities in $\mathcal{Y}_{q\left(\lambda\right)}\simeq\tilde{\mathcal{Y}}_{\lambda}$. Since the decomposition is orthogonal, we have:
$\mathrm{rank}\ \Phi_{\lambda}=\mathrm{rank}\ \Phi_{1,\lambda}+\cdots+\mathrm{rank}\ \Phi_{m,\lambda}.$
2.3. Period Map of Twistor Space
For any non-trivial finite group $\Gamma<\mathrm{SU}(2)$. Using the same notations as in the introduction 1.1, let $\zeta=(\zeta^{1},\zeta^{2},\zeta^{3})\in(Z\otimes\mathbb{R}^{3})^{\circ}$. Recall we denote by $I_{\zeta},J_{\zeta},K_{\zeta}$ the three compatible complex structures on $X_{\zeta}$, the associated Kähler forms by $\omega_{\zeta}^{k}$, $k\in\{1,2,3\}$ and by $\mathrm{Z}(X_{\zeta})$ the twistor space of the hyperKähler manifold $\left(X_{\zeta},g_{\zeta}\right)$. Recall the holomorphic projection $\pi:\mathrm{Z}(X_{\zeta})\to\mathbb{P}^{1}$.
Proposition 2.12.
[9, pp.691-692] Related to $\mathrm{Z}(X_{\zeta})$, there is a $\mathbb{C}^{*}$-deformation of $\mathbb{C}^{2}/\Gamma$:
$\phi:Y\to\mathbb{C}^{2}.$
where the $\mathbb{C}^{*}$-action on $\mathbb{C}^{2}$ is the natural diagonal action and $\phi^{-1}(0)\simeq\mathbb{C}^{2}/\Gamma$. We define a map $\pi^{s}:Z^{s}\to\mathbb{P}^{1}$ by:
${\left(Y\setminus\phi^{-1}(0)\right)/\mathbb{C}^{*}\quad\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$        $\scriptstyle{\phi}$${\mathrm{Z}^{s}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{s}}$${\left(\mathbb{C}^{2}\setminus{0}\right)/\mathbb{C}^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{P}^{1}.}$
Then there exists a commutative diagram below:
(2.2) ${\mathrm{Z}(X_{\zeta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\chi}$$\scriptstyle{\pi}$${\mathrm{Z}^{s}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{s}}$${\mathbb{P}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{P}^{1},}$
such that the restriction of $\chi$ to each fiber of $\pi$ provides the minimal resolution of corresponding fiber of $\pi^{s}$. We call the projection $\pi:\mathrm{Z}(X_{\zeta})\to\mathbb{P}^{1}$ a simultaneous resolution of the map $\pi^{s}$.
Being the twistor space of a hyperKähler manifold, $\mathrm{Z}(X_{\zeta})$ carries a twisted vertical 2-form [7] $\Omega\in H^{0}(\mathrm{Z}(X_{\zeta}),\,\Lambda^{2}T_{F}^{*}\otimes\pi^{*}\mathcal{O}_{\mathbb{P}^{1}}(2))$ defined on the open set $\mathbb{C}$ of $\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$ by
$\Omega(u)=\omega^{2}_{\zeta}+\mathrm{i}\omega^{3}_{\zeta}+u\left(2\omega^{1}_{\zeta}\right)-u^{2}\left(\omega^{2}_{\zeta}-\mathrm{i}\omega^{3}_{\zeta}\right)$
[…]
Note that $\Omega(u)$ is a holomorphic two-form on $(X_{\zeta},I_{\zeta}^{u})$, where $I_{\zeta}^{u}$ is defined in the paragraph of (1.1). We call the cohomology class $[\Omega(u)]$ the period of $(X_{\zeta},I_{\zeta}^{u})$. Taking the cohomology class of $\Omega$ gives, under the identification $H^{2}(X_{\zeta},\mathbb{C})=\mathfrak{h}^{\mathbb{C}}$, an element of $H^{0}(\mathbb{P}^{1},\mathfrak{h}^{\mathbb{C}}\otimes\mathcal{O}_{\mathbb{P}^{1}}(2))$ and a map when lifted to $\mathbb{C}^{2}$:
(2.3) $\widetilde{p_{\zeta}}:\quad\mathbb{C}^{2}$ $\to\mathfrak{h}^{\mathbb{C}}$
$(z_{1},z_{2})$ $\mapsto z_{1}^{2}[\omega^{2}_{\zeta}+\mathrm{i}\omega^{3}_{\zeta}]+2z_{1}z_{2}[\omega^{1}_{\zeta}]-z_{2}^{2}[\omega^{2}_{\zeta}-\mathrm{i}\omega^{3}_{\zeta}].$
[…]
Lemma 2.13.
[9, pp.693] Recall the semi-universal deformation $f:\mathcal{Y}\to\mathcal{V}$ of $\mathbb{C}^{2}/\Gamma$ in Lemma 2.4. We have that the deformation $\phi$ is the pullback of the semi-universal deformation by the period map $p_{\zeta}$, i.e. we have the following diagram:
${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${\mathcal{Y}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${\mathbb{C}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{\zeta}}$${\mathcal{V}.}$
The singular fibers of (the pullback of) semi-universal deformation $f$ and the types of their singularities are characterized by Lemma 2.6. Combining these with Lemma 2.13 and Proposition 2.12, we have:
Corollary 2.14.
Recall the holomorphic map $\pi:\mathrm{Z}(X_{\zeta})\to\mathbb{P}^{1}$. Given $u\in\mathbb{P}^{1}$, then:
  • (1)
    The complex submanifold $\pi^{-1}(u)$ is isomorphic to the minimal resolution of the fiber $\phi^{-1}(\tilde{u})$ of $\phi$, where $\tilde{u}$ is any lifting of $u$ with respect to the natural quotient map $\mathbb{C}^{2}\setminus 0\to\mathbb{P}^{1}$.
  • (2)
    $\pi^{-1}(u)$ contains a rational curve if and only if the period $\widetilde{p_{\zeta}}(\tilde{u})\in\mathfrak{h}^{\mathbb{C}}$ lies in the kernel of some root in $\Phi$.
  • (3)
    The number of rational curves in $\pi^{-1}(u)$ is given by $\mathrm{rank}\,\Phi_{\lambda}$, where $\lambda:=\widetilde{p_{\zeta}}(\tilde{u})$.
Proof.
The statement (1) is immediately obtained from Proposition 2.12. By Lemma 2.13, the fiber $\phi^{-1}(\tilde{u})$ is isomorphic to the fiber $f^{-1}(p_{\zeta}(\tilde{u}))$ hence isomorphic to the fiber $\tilde{f}^{-1}(\widetilde{p_{\zeta}}(\tilde{u}))$. Thus, statement (2) is obtained from the Lemma 2.6. And statement (3) is a straightforward consequence of Corollary 2.11. ∎
2.4. An Isomorphism Between HyperKähler ALEs
Lemma 2.15.
[8, Proposition 4.1] Given $\zeta=\left(\zeta^{1},\zeta^{2},\zeta^{3}\right)\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$. There exists a linear isomorphism
$\sigma:Z\xrightarrow{\sim}H^{2}(X_{\zeta}^{\mathrm{Diff}},\mathbb{R})(\simeq\mathfrak{h}),$
such that $\sigma(\zeta^{k})=[\omega^{k}_{\zeta}]$, for $k=1,2,3.$ Consequently, the complexification $\sigma_{\mathbb{C}}:Z\otimes\mathbb{C}\to H^{2}(X_{\zeta}^{\mathrm{Diff}},\mathbb{C})$ is also an isomorphism.
Lemma 2.16.
Without loss of generality, we assume that $u\neq\infty$ and denote $u=u_{1}+\mathrm{i}u_{2}\in\mathbb{C}\hookrightarrow\mathbb{P}^{1}$. Recall the notations for complex structures: $I_{\zeta}^{u}=aI_{\zeta}+bI_{\zeta}+cI_{\zeta}$. For $\zeta=\left(\zeta^{1},\zeta^{2},\zeta^{3}\right)\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$, let
$\tilde{\zeta}^{2}$ $:=2u_{1}\zeta^{1}+(1-u_{1}^{2}+u_{2}^{2})\zeta^{2}-2u_{1}u_{2}\zeta^{3},$
$\tilde{\zeta}^{3}$ $:=2u_{2}\zeta^{1}-2u_{1}u_{2}\zeta^{2}+(1+u_{1}^{2}-u_{2}^{2})\zeta^{3}.$
Then the complex manifold $\left(X_{\zeta},I_{\zeta}^{u}\right)$ is isomorphic to the complex manifold $\left(X_{\tilde{\zeta}},I_{\tilde{\zeta}}\right)$, where $\tilde{\zeta}=\left(\tilde{\zeta}^{1},\tilde{\zeta}^{2},\tilde{\zeta}^{3}\right)$ for any $\tilde{\zeta}^{1}\in Z$ such that $\tilde{\zeta}\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$.
Proof.
For the complex manifold $\left(X_{\zeta},I_{\zeta}^{u}\right)$, there is a holomorphic two form $\Omega(u):=2u\,\omega_{\zeta}^{1}+\left(1-u^{2}\right)\omega_{\zeta}^{2}+\mathrm{i}\left(1+u^{2}\right)\omega_{\zeta}^{3}$. By the isomorphism in Lemma 2.15, the period $[\Omega(u)]$ of $\left(X_{\zeta},I_{\zeta}^{u}\right)$ can also be given by an element in $Z\otimes\mathbb{C}$ as follows:
$\sigma_{\mathbb{C}}^{-1}\,[\Omega(u)]=\zeta^{2}+\mathrm{i}\zeta^{3}+2u\zeta^{1}-u^{2}\left(\zeta^{2}-\mathrm{i}\zeta^{3}\right).$
By the definitions of $\tilde{\zeta}^{2}$ and $\tilde{\zeta}^{3}$, the period of $(X_{\tilde{\zeta}},I_{\tilde{\zeta}})$ in $Z\otimes\mathbb{C}$ is given by $\tilde{\zeta}^{2}+\mathrm{i}\tilde{\zeta}^{2}$ which also equals to $\zeta^{2}+\mathrm{i}\zeta^{3}+2u\zeta^{1}-u^{2}\left(\zeta^{2}-\mathrm{i}\zeta^{3}\right)$. Hence $\left(X_{\zeta},I_{\zeta}^{u}\right)$ and $(X_{\tilde{\zeta}},I_{\tilde{\zeta}})$ have the same period. Denote $u^{\prime}\in\mathbb{P}^{1}$ the corresponding point of $(1,0,0)\in\mathbb{S}^{2}$, and $\tilde{u}^{\prime}\in\mathbb{C}^{2}$ any lifting of $u^{\prime}$. Then,
$\widetilde{p_{\zeta}}(\tilde{u})=\widetilde{p_{\tilde{\zeta}}}(\tilde{u}^{\prime})=:\lambda_{0}.$
As in the proof of Corollary 2.14, $\left(X_{\zeta},I_{\zeta}^{u}\right)$ is the minimal resolution of the fiber $\tilde{f}^{-1}(\lambda_{0})$ of $\tilde{f}$. Similarly, $(X_{\tilde{\zeta}},I_{\tilde{\zeta}})$ is also the minimal resolution of the fiber $\tilde{f}^{-1}(\lambda_{0})$ of $\tilde{f}$. Since the minimal resolution of a normal surface is unique, see Lemma 2.8, they are isomorphic to each other.
Remark 2.17.
When $u=\infty$, then $I_{\zeta}^{u}=-I_{\zeta}$. And the complex manifold $\left(X_{\zeta},-I_{\zeta}\right)$ is isomorphic to $\left(X_{\tilde{\zeta}},I_{\tilde{\zeta}}\right)$, where $\tilde{\zeta}:=\left(\zeta^{1},-\zeta^{2},\zeta^{3}\right)$.
3. The hyperKähler ALEs as minimal resolution of Kleinian singularities
In this section, we begin the proofs of our main theorems. In subsection 3.1, we prove Theorem 1.1. In subsection 3.2, we prove the Case (i) in Theorem 1.2. And the rest part of Theorem 1.2 will be proved in section 4.
3.1. Proof of Theorem 1.1:
(if part). Suppose $I_{\mathrm{std}}\in\{aI_{\zeta}+bJ_{\zeta}+cK_{\zeta}\mid(a,b,c)\in\mathbb{S}^{2}\}$ for some given $\zeta\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$. Assume first that $I_{\mathrm{std}}\simeq I_{\zeta}$, i.e., $\left(X_{\zeta},I_{\zeta}\right)\simeq\widetilde{\mathbb{C}^{2}/\Gamma}$. If $(\zeta^{2},\zeta^{3})\neq 0$, then by Lemma 2.15 the cohomology class of the holomorphic 2-form on $\left(X_{\zeta},I_{\zeta}\right)$ is non-zero i.e., $[\omega_{\mathbb{C}}]=[\omega_{2}]+[\mathrm{i}\omega_{3}]=\sigma(\zeta^{2}+\mathrm{i}\zeta^{3})\neq 0$. Since the second homology group $H_{2}(X_{\zeta},\mathbb{Z})$ is generated by the two-spheres $\Sigma_{i},i=1,\cdots,r$ in exceptional locus of $\widetilde{\mathbb{C}^{2}/\Gamma}$, so there exists some two-sphere $[\Sigma_{i_{0}}]\in H_{2}(X_{\zeta},\mathbb{Z})$ such that $\omega_{\mathbb{C}}[\Sigma_{i_{0}}]\neq 0$. Consequently, the exceptional curve $\Sigma_{i_{0}}$ cannot be holomorphic on $\left(X_{\zeta},I_{\zeta}\right)$, contradiction. Therefore, those $\zeta$ for which $X_{\zeta}$ with respect to the first complex structure $I_{\zeta}$ is isomorphic to $\widetilde{\mathbb{C}^{2}/\Gamma}$ must be of the form $(\zeta^{1},0,0)$. As stated in the introduction, Kronheimer proved that if $\zeta=(\zeta^{1},0,0)$ then $\left(X_{\zeta},I_{\zeta}\right)\simeq\widetilde{\mathbb{C}^{2}/\Gamma}$. Therefore, $\left(X_{\zeta},I_{\zeta}\right)\simeq\widetilde{\mathbb{C}^{2}/\Gamma}$ if and only if $\zeta=(\zeta^{1},0,0)$.
Now, assume that $I_{\mathrm{std}}\simeq I_{\zeta}^{u}$ for some $u\in\mathbb{P}^{1}$ (without loss of generality we always assume $u\in\mathbb{C}$). And choose a $\tilde{\zeta}$ = $\left(\tilde{\zeta^{1}},\tilde{\zeta^{2}},\tilde{\zeta^{3}}\right)$ which is defined as in Lemma 2.16. Then $\left(X_{\zeta},I_{\zeta}^{u}\right)\simeq\left(X_{\tilde{\zeta}},I_{\tilde{\zeta}}\right)$. Since $I_{\mathrm{std}}\simeq I_{\zeta}^{u}$ i.e., $\left(X_{\tilde{\zeta}},I_{\tilde{\zeta}}\right)\simeq\widetilde{\mathbb{C}^{2}/\Gamma}$, it follows from the above argument that $\tilde{\zeta}^{2}=\tilde{\zeta}^{3}=0$, that is
(3.1) $\begin{cases}2u_{1}\zeta^{1}+(1-u_{1}^{2}+u_{2}^{2})\zeta^{2}-2u_{1}u_{2}\zeta^{3}=0\\ 2u_{2}\zeta^{1}-2u_{1}u_{2}\zeta^{2}+(1+u_{1}^{2}-u_{2}^{2})\zeta^{3}=0.\\ \end{cases}$
By checking the rank of these equations, we have $\mathrm{rank}(\zeta)=1$.
(only if part). Given $\zeta\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$ such that $\mathrm{rank}(\zeta)=1$, i.e., $\zeta=(x\zeta^{1},y\zeta^{1},z\zeta^{1})$ for $\zeta^{1}\in Z\setminus\bigcup_{\theta\in\Phi}D_{\theta}$ and some non-zero vector $v=(x,y,z)\in\mathbb{R}^{3}\setminus\{0\}$. From the above argument that $\left(X_{\zeta},I_{\zeta}\right)\simeq\widetilde{\mathbb{C}^{2}/\Gamma}$ if and only if $\zeta=(\zeta^{1},0,0)$, we know that to find $u$ such that $I_{\mathrm{std}}\simeq I_{\zeta}^{u}$ we need to solve the system of equations in (3.1) for $(u_{1},u_{2})\in\mathbb{R}^{2}$. The system of equations (3.1) holds if and only if the period of $\left(X_{\zeta},I_{\zeta}^{u}\right)$ is zero. In other words, if we lift $u\in\mathbb{P}^{1}$ to $(z_{1},z_{2})\in\mathbb{C}^{2}$, we must solve for $\left(z_{1},z_{2}\right)\in\mathbb{C}^{2}\setminus\{0\}$ from equation:
$z_{1}^{2}\left(y\zeta^{1}+\mathrm{i}z\zeta^{1}\right)+2z_{1}z_{2}x\zeta^{1}-z_{2}^{2}\left(y\zeta^{1}-\mathrm{i}z\zeta^{1}\right)=0.$
Since $\zeta^{1}\neq 0$, this reduces to solving $(z_{1},z_{2})\in\mathbb{C}^{2}\setminus\{0\}$ from:
$z_{1}^{2}\left(y+\mathrm{i}z\right)+2z_{1}z_{2}x-z_{2}^{2}\left(y-\mathrm{i}z\right)=0.$
A non-trivial solution exists if and only if the complex numbers $y+\mathrm{i}z$, $2x$, and $-y+\mathrm{i}z$ satisfy:
(3.2) $w_{1}\left(y+\mathrm{i}z\right)+w_{2}(2x)+w_{3}\left(-y+\mathrm{i}z\right)=0$
for some complex numbers $(w_{1},w_{2},w_{3})\in\mathbb{C}^{3}\setminus\{0\}$ such that $w_{1}\cdot w_{3}=w_{2}^{2}$. In terms of projective space:
$\mathbb{P}^{2}=\left\{\left[w_{1}:w_{2}:w_{3}\right]\mid\left(w_{1},w_{2},w_{3}\right)\in\mathbb{C}^{3}\setminus 0\right\},$
it means that the complex projective line defined by (3.2) intersects with the smooth quadratic curve defined by $w_{1}\cdot w_{3}=w_{2}^{2}$ in $\mathbb{P}^{2}$. Since such an intersection always occurs, a solution $(z_{1},z_{2})$ is guaranteed to exist. This completes the first statement in Theorem 1.1. Now we prove the remaining part.
First, we note that the period of $\left(X_{\zeta},I^{u}_{\zeta}\right)$, for which $I_{\zeta}^{u}\simeq I_{\mathrm{std}}$, is equal to zero, since the type of its exceptional locus is exactly $\Phi$, by Lemma 2.6 its period must be zero. Since $\mathrm{rank}(\zeta)=1$, the image of the period map $p_{\zeta}$ is a one-dimensional subspace of $\mathfrak{h}^{\mathbb{C}}$. The image of $p_{\zeta}$ cannot lie in the branch locus $\mathcal{D}$ [9, p.693-p.694]. By Lemma 2.4, $\mathcal{D}$ is a union of linear subspaces, so $\mathrm{Im}p_{\zeta}$ intersects $\mathcal{D}$ only at 0. Hence, by Lemma 2.13 the fiber $\phi^{-1}(s)$ of $\phi$ for any $s\in\mathrm{Im}p_{\zeta}\backslash 0$ is smooth. Since the deformation $\phi$ is a also $\mathbb{C}^{*}$-deformation, the $\mathbb{C}^{*}$-action implies that $\phi^{-1}(s)$ and $\phi^{-1}(\gamma s),\,\gamma\in\mathbb{C}^{*}$ are isomorphic as complex manifolds, we conclude that the complex structures in the hyperKählerian family other than $I_{\mathrm{std}}$ are all equivalent. By Lemma 2.5, the complex surface $\phi^{-1}(s)$ is a smooth hypersurface in $\mathbb{C}^{3}$. Therefore, if $\left(X_{\zeta},I^{u}_{\zeta}\right)$ contains rational curves, then $I_{\zeta}^{u}\simeq I_{\mathrm{std}}$. We will show the number of complex structures that contains rational curves is $2$ in next subsections. ∎
3.2. Proof of Case (i) of Theorem 1.2
By Theorem 1.1, the number of $u\in\mathbb{P}^{1}$ for which $\left(X_{\zeta},I_{\zeta}^{u}\right)$ contains rational curves is exactly the number of $u\in\mathbb{P}^{1}$ such that $I_{\zeta}^{u}\simeq I_{std}$. By the proof of Theorem 1.1, it is the number of intersection points of the projective line given by equation (3.2) and the conic curve given by $w_{1}\cdot w_{3}=w_{2}^{2}$. By elementary calculations, they are tangent with each other if and only if the three real number $x,y,z$ satisfy
$x^{2}+(y+\mathrm{i}z)(y-\mathrm{i}z)=0.$
The only solution is $(x,y,z)=(0,0,0)$ and the corresponding $\zeta$ is equal to zero, which will contradict to our assumption $\mathrm{rank}(\zeta)=1$. Hence the projective line of the form (3.2) and the conic curve always intersect transitively, hence the intersection number is always 2. Thus, we obtain the equation (1.2):
(3.3) $\sharp\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}=2,$
and equation (1.3) follows immediately from the fact that the number of rational curves in $\widetilde{\mathbb{C}^{2}/\Gamma}$ is $r$:
[…]
Suppose $\Gamma=\mathbb{Z}_{2}$. Then $\mathrm{dim}\,Z=1$. Hence $\mathrm{rank}(\zeta)=1$ for all $\zeta\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$. Therefore, by Theorem 1.1 and Theorem 1.2 Case (i), we obtain :
Corollary 3.1.
In the hyperKählerian family of any hyperKähler ALE with the group $\mathbb{Z}_{2}$ at infinity, there exist exactly two complex structures which is isomorphism to the minimal resolution of $\mathbb{C}^{2}/\mathbb{Z}_{2}$.
Remark 3.2.
The further properties of its twistor space is discussed Example 6.2 in section 5.
4. Counting rational curves
In this section, we will complete the proof of Theorem 1.2. Recall there are three cases for $\zeta$: (i) $\mathrm{rank}(\zeta)$ = 1; (ii) $\mathrm{rank}(\zeta)=2$; (ii) $\mathrm{rank}(\zeta)=3$. We have proved the Theorem 1.2 in Case (i) in section 3. We will prove Theorem 1.2 for Case (ii) and Case (iii) from subsection 4.2 to subsection 4.3. In subsection 4.4, we provide a proof of Corollary 1.6. As a technical preparation of the proof of Case (ii), subsection 4.1 will introduce a type of decomposition of irreducible root system and present some useful combinatorial results that are also of interest to us.
Before detailing the following subsections, let’s briefly discuss our strategy for proving Theorem 1.2. According to Corollary 2.14 and Lemma 2.6, the image of period map $\tilde{p_{\zeta}}$ gives the information of the complex structures containing rational curves and the number of rational curves. Therefore, to bound the number in equation (1.4):
$\sharp\left\{u\in\mathbb{P}^{1}\mid\pi^{-1}(u)\text{ contains a rational curve}\right\}$
and the number in equation (1.5):
$\sharp\sum_{u\in\mathbb{P}^{1}}\sharp\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}(u)\right\}$
we need to analyze the ’position’ of the image of $\widetilde{p_{\zeta}}$. The ’position’ is determined by its intersection with the kernels of roots, which will induce a decomposition of the root system $\Phi$. The decomposition showed in subsection 4.1 is a simple one which is designed for Case (ii). It is also a simple version for Case (iii). We will follow (a little) different approaches to consider these two cases separately. The reader may skip the technical details in subsection 4.1 and see the proof for Case (ii) in subsection 4.2 and the the proof for Case (iii) in subsection 4.3.
4.1. A Decomposition of An Irreducible Root System
For a non-trivial finite group $\Gamma\leq\mathrm{SU}(2)$, we recall the notations $\mathfrak{h}^{\mathbb{C}},\Phi$, and $\Delta$ denote the complex Cartan subalgebra, the root system, and the simple roots, respectively. See subsection 2.1. We always assume that $\mathrm{rank}\,\Phi\geq 2$, i.e., $\Gamma\neq\mathbb{Z}_{2}$. We choose a two-dimensional subspace $L$ in $\mathfrak{h}^{\mathbb{C}}$, which does not lie in the kernel of any root in $\Phi$. Hence, the restriction of each root $\theta\in\Phi\subset\mathrm{Hom}(\mathfrak{h},\mathbb{R})\hookrightarrow\mathrm{Hom}(\mathfrak{h}^{\mathbb{C}},\mathbb{C})$ to $L$ has a one-dimensional kernel denoted by $l_{\theta}$. The choice of the subspace $L$ induces a map:
$\mu_{L}:\Phi$ $\to\mathbb{P}(L),$
$\theta$ $\mapsto[l_{\theta}].$
Since $\Phi$ is a finite set, the image of the map $\mu_{L}$ is also finite, which can be expressed as $\mathrm{Im}\mu_{L}=\{p_{1},\cdots,p_{s}\}$. Hence there exists a type of decomposition of $\Phi$ into disjoint union of root subsystems as:
(4.1) $\Phi$ $=\Phi_{1}\sqcup\Phi_{2}\sqcup\cdots\sqcup\Phi_{s},$
$\Delta$ $=\Delta_{1}\sqcup\Delta_{2}\sqcup\cdots\sqcup\Delta_{s},$
where $\Phi_{k}:=\mu_{L}^{-1}\{p_{k}\}$ and $\Delta_{k}:=\Phi_{k}\cap\Delta$ for $k=1,\ldots,s$. Each $\Phi_{k}$ is a root subsystem in $\Phi$ that satisfies $\Phi_{k}\bigcap\mathrm{span}_{\mathbb{R}}\Phi_{k}=\Phi_{k}$. It is important to note that this decomposition cannot be orthogonal and $\Delta_{k}$ may not form a root base of the root subsystem $\Phi_{k}$. Since $\bigcap_{\theta\in\Phi}\mathrm{Ker}\,\theta=0\in\mathfrak{h}$, we have $\bigcap_{\theta\in\Phi}l_{\theta}=0\in V$. Hence $s\neq 1$, moreover,
Lemma 4.1.
$3\leq s\leq\frac{1}{2}\sharp\Phi$.
Proof.
The upper bound is straightforward. Now, we prove the lower bound by contradiction. Assume $s=2$. Let $\mathrm{Im}\mu_{L}=\{[l_{\theta_{1}}],[l_{\theta_{2}}]\}$ for two distinct roots $\theta_{1},\theta_{2}$. The corresponding decompositions are:
$\Phi$ $=\Phi_{1}\sqcup\Phi_{2},$
$\Delta$ $=\Delta_{1}\sqcup\Delta_{2}.$
Neither $\Delta_{k}$ for $k=1,2$, is empty. For example, if $\Delta_{1}=\emptyset$, then $\Delta_{2}=\Delta$, hence $\Phi_{2}=\Phi$, which implies $\Phi_{1}=\emptyset$. Since $\Delta$ consists of the simple roots of an irreducible root system $\Phi$, there exist $\alpha_{1}\in\Delta_{1}$ and $\alpha_{2}\in\Delta_{2}$ such that $(\alpha_{1},\alpha_{2})\neq 0$ in the ambient Euclidean space of $\Phi$. Hence, either $\alpha_{1}+\alpha_{2}\in\Phi$ or $\alpha_{1}-\alpha_{2}\in\Phi$. Assume $\alpha_{1}+\alpha_{2}\in\Phi$. Since $l_{\alpha_{1}+\alpha_{2}}=l_{\theta_{1}}=l_{\alpha_{1}}$, it follows that $l_{\alpha_{1}}=l_{\alpha_{2}}$. Consequently, $l_{\alpha_{1}}=l_{\alpha_{2}}$ implies $l_{\theta_{1}}=l_{\theta_{2}}$, which contradicts the condition that $l_{\theta_{1}}\neq l_{\theta_{2}}$. Therefore, for $k=1,2$, $l_{\alpha_{1}+\alpha_{2}}\neq l_{\theta_{k}}$, which means $\alpha_{1}+\alpha_{2}\notin\Phi_{k},k=1,2$, leading to a contradiction. ∎
Define further $r:=\mathrm{rank}\,\Phi$, $r_{k}:=\mathrm{rank}\,\Phi_{k}$, and $\delta_{k}=\sharp\Delta_{k}$. Obviously, $\delta_{k}\leq r_{k}$, and $r\leq\sum_{k=1}^{s}\,r_{k}$. Moreover, we have :
Lemma 4.2.
(4.2) $r+1\leq\sum_{k=1}^{s}\,r_{k}\leq\frac{1}{2}\sharp\Phi.$
Proof.
The upper bound is obvious because increasing the number of root subsystems in the decompositions increases the sum. We now prove the lower bound by contradiction. Assume $r=\sum_{k=1}^{s}r_{k}$. Since $r=\sum_{k=1}^{s}\delta_{k}$ and $\delta_{i}\leq r_{k}$, it follows that for each $k=1,\cdots,s$ we have $r_{k}=\delta_{k}$, which implies $\mathrm{rank}\,\Phi_{k}=\sharp\Delta_{k}$. Therefore, $\Delta_{k}$ forms a basis for the root subsystem $\Phi_{k}$. There exists the maximal root $\theta_{max}$ of $\Phi$ with respect to $\Delta$ by the irreducibility of $\Phi$, and $\theta_{max}=\sum_{\alpha\in\Delta}k_{\alpha}\alpha$ with $k_{\alpha}>0$ for any $\alpha\in\Delta$. Hence, $\theta_{max}$ cannot lie in any root subsystem $\Phi_{k}$ since $\Delta_{k}$ is a proper subset of $\Phi$, which leads to a contradiction. ∎
For the purpose of producing a better lower bound $12r-$ 1 than Lemma 4.2, we introduce a certain space of decompositions of the irreducible root system $\Phi$ as:
Definition 4.3.
We call an expression $\Phi=\Phi_{1}\sqcup\cdots\sqcup\Phi_{s}$ a decomposition of type-$1$ of $\Phi$ if the following conditions hold:
  • (1)
    $s\neq 0$.
  • (2)
    For each $k\in\left\{1,\cdots,s\right\}$, the $\Phi_{k}$ is a proper and non-empty subset of $\Phi$. Moreover, it is a closed root subsystem of $\Phi$, where closed means:
    $\Phi_{k}=\Phi\cap\mathrm{span}_{\mathbb{R}}\,\Phi_{k}.$
  • (3)
    Each $\Phi_{k}$ is disjoint from the others.
Denote $\mathfrak{D}^{1}\left(\Phi\right)$ as the space of all decompositions of type-$1$ of $\Phi$.
Remark 4.4.
Obvious the decomposition above which is induced from a two-dimensional subspace $L$ is a decomposition of type-1. But not all decompositions are induced from a two-dimensional subspace $L$ as before, so the space $\mathfrak{D}^{1}\left(\Phi\right)$ is somewhat coarse.
For any decomposition $\Phi=\Phi_{1}\sqcup\cdots\sqcup\Phi_{s}$ of type-$1$, define $r_{k}:=\mathrm{rank}\,\Phi_{k}$. Denote $f_{1}\left(\Phi\right)$ the infimum of the sum $\sum_{k=1}^{s}r_{k}$ for all the decompositions in $\mathfrak{D}^{1}\left(\Phi\right)$, i.e.,
$f_{1}\left(\Phi\right):=\inf_{\mathfrak{D}^{1}\left(\Phi\right)}\left\{\sum_{k=1}^{s}r_{k}\right\}.$
Since the proof of Lemma 4.2 only depend on the irreducibility of $\Phi$ and condition (1)(2), we immediately obtain $f_{1}\left(\Phi\right)\geq r+1$. Moreover, there is a better lower bound:
Proposition 4.5.
For $\Phi(n)=A_{n\geq 2},\,D_{n\geq 3}$ or $E_{6},E_{7},E_{8}$, i.e. $\Phi(n)$ is the root system of type $ADE$ with rank $n\geq 2$, we have estimate:
$2n-1\leq f_{1}(\Phi(n)).$
This can be proved by induction on $n$. As a preparation of its proof, we need a Lemma first. Let $\Phi(n+1)$ be of type ADE with rank equals to $n+1$ in an ambient Euclidean space $\mathrm{E}^{n+1}$. Then the root system $A_{n}$ can be naturally embedded into $\Phi(n+1)$. We always fix an embedding which we will choose in the proof of the following lemma. We call the set $S:=\Phi(n+1)\setminus A_{n}$ the set of extra roots of $\Phi(n+1)$ with respect to $A_{n}$ Then:
Lemma 4.6.
The set of extra roots $S$ span $\mathrm{E}^{n+1}$, that is, $\mathrm{dim}_{\mathbb{R}}\,\mathrm{span}_{\mathbb{R}}S=n+1$.
Proof.
The proof involves explicitly calculating for each type of root systems.
[…]
    $\alpha_{1}=\epsilon_{1}-\epsilon_{2}$
    $\cdots$
    $\alpha_{i}=\epsilon_{i}-\epsilon_{i+1}$
    $\cdots$
    $\alpha_{n}=\epsilon_{n}-\epsilon_{n+1}$
    $\alpha_{n+1}=\epsilon_{n+1}-\epsilon_{n+2}$
    Then we can write the positive roots as linear combinations of these simple roots as:
    $\epsilon_{i}-\epsilon_{j}$ $=\sum_{i\leq k<j}\alpha_{k}$ $1\leq i<j\leq n+2;$
    The first $n$ simple roots $\alpha_{1},\cdots,\alpha_{n}$ determine the embedding of $A_{n}\hookrightarrow A_{n+1}$. The ’extra’ positive roots with respect to $A_{n}$ are: $S^{+}:=\left\{\epsilon_{i}-\epsilon_{n+2},1\leq i\leq n+1\right\}$. By elementary computation, we see:
    $\mathrm{span}_{\mathbb{R}}\,S^{+}=\mathrm{E}.$
    Thus, $\dim\mathrm{span}_{\mathbb{R}}S=n+1$.
[…]
    $\alpha_{1}=\epsilon_{1}-\epsilon_{2}$
    $\cdots$
    $\alpha_{i}=\epsilon_{i}-\epsilon_{i+1}$
    $\cdots$
    $\alpha_{n}=\epsilon_{n}-\epsilon_{n+1}$
    $\alpha_{n+1}=\epsilon_{n}+\epsilon_{n+1}$
    Then we draw the Dynkin diagram of $D_{n+1}$ together with an embedding of $A_{n}$ as:
[…]
    We also write the positive roots as linear combinations of these simple roots:
    $\epsilon_{i}-\epsilon_{j}$ $=\sum_{i<k<j}\alpha_{k}$ $1\leq i<j\leq n+1;$
    $\epsilon_{i}+\epsilon_{j}$ $=\sum_{i\leq k<j}\alpha_{k}+2\left(\sum_{j\leq k<n}\alpha_{k}\right)+\alpha_{n}+\alpha_{n+1}$ $1\leq i<j<n+1;$
    $\epsilon_{i}+\epsilon_{n+1}$ $=\sum_{i\leq k\leq n-1}\alpha_{k}+\alpha_{n+1}$ $1\displaystyle 1\leq i<n+$ 1 .
    The simple roots $\alpha_{1},\cdots,\alpha_{n}$ determine an embedding of $A_{n}\hookrightarrow D_{n+1}.$ The ’extra’ (positive) roots with respect to $A_{n}$ are: $S^{+}:=\left\{\epsilon_{i}+\epsilon_{j},1\leq i<j\leq n+1\right\}$. By elementary computation, we see:
    $\mathrm{span}_{\mathbb{R}}\,S=\mathrm{E}.$
    Thus, $\dim\mathrm{span}_{\mathbb{R}}S=n+1$.
  • For $\Phi(n+1)=E_{6},E_{7},E_{8}$. We only show here the case $\Phi(n+1)=E_{6}$, since the other cases are basically the same. The Dynkin diagram of $E_{6}$ together with an embedding of $A_{5}$ is
[…]
    where we choose a root base $\left\{\alpha_{1},\cdots,\alpha_{6}\right\}$. The simple roots $\alpha_{1},\alpha_{3},\cdots,\alpha_{6}$ determine an embedding of $A_{5}\hookrightarrow E_{6}$. We denote a root $a\alpha_{1}+b\alpha_{2}+c\alpha_{3}+d\alpha_{4}+e\alpha_{5}+f\alpha_{6}$ in the root system $E_{6}$ by a diagram $\begin{pmatrix}&&b&&\\ a&c&d&e&f\end{pmatrix}$. Then we cite from [2] that the following roots are part of the extra roots of $E_{6}$ with respect to $A_{5}$:
    $\begin{pmatrix}&&1&&\\ 0&1&2&1&0\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&1&2&1&0\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 0&1&2&1&1\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&2&2&1&0\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&1&2&1&1\end{pmatrix}$
    $\begin{pmatrix}&&1&&\\ 0&1&2&2&1\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&2&2&1&1\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&1&2&2&1\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&2&2&2&1\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&2&3&2&1\end{pmatrix}$
    $\begin{pmatrix}&&2&&\\ 1&2&3&2&1\end{pmatrix}$
    As an observation, the $\mathbb{R}$-span of these ’extra’ roots of $E_{6}$ with respect to $A_{5}$ contains all the simple roots $\alpha_{1},\cdots,\alpha_{6}$. For example, the difference of the first two roots above is $\alpha_{1}$, hence $\alpha_{1}$ lies in the span of extra roots. Thus, $\dim\mathrm{span}_{\mathbb{R}}S=6$.
Proof of Proposition 4.5: Since $\Phi(n+1)$ is a finite set, the set of all decompositions is also finite. Hence we can take a decomposition of $\Phi(n+1)$
$\Phi(n+1)=\Phi_{1}\sqcup\cdots\sqcup\Phi_{s}\quad,\quad r_{k}:=\mathrm{rank}\,\Phi_{k}.$
such that $\sum_{k=1}^{s}r_{k}=f_{1}(\Phi(n+1))$. By deleting the extra roots, we obtain:
(4.3) $A_{n}=\Phi_{1}^{\prime}\sqcup\cdots\sqcup\Phi_{s}^{\prime}\quad,\quad\Phi_{k}^{\prime}:=\Phi_{k}\setminus S\,,\,r_{k}^{\prime}=\mathrm{rank}\,\Phi_{k}^{\prime}.$
By definition, each set $\Phi_{k}^{\prime}$ for $k=1,\cdots,s$ is either empty or a root subsystem of $A_{n}$. Set $s^{\prime}:=\left\{k\mid\Phi_{k}^{\prime}\neq\emptyset\right\}$, then we have two possibilities:
  • $s^{\prime}=1$. Then $\Phi_{k_{0}}=A_{n}$ for some $k_{0}$ and the other $\Phi_{k}$, $k\neq k_{0}$, contains only extra roots. By Lemma 4.6:
    $f_{1}\left(\Phi(n+1)\right)$ $=\sum_{k=1}^{s}r_{k}$
    $=r_{k_{0}}+\sum_{k\neq k_{0}}r_{k}$
    $\geq n+n+1=2n+1.$
  • $s^{\prime}>1$. Then the disjoint union (4.3) is a decomposition of $A_{n}$ of type-1. By Lemma 4.6, these extra roots cannot be put into only one root subsystem, hence there exist at least two root subsystems $\Phi_{k_{i}},i=1,2$ such that $r_{k_{i}}\geq r_{k_{i}}^{\prime}+1$. Therefore,
    $f_{1}\left(\Phi(n+1)\right)$ $=\sum_{k=1}^{s}r_{k}$
    $\geq\sum_{k=1}^{s}r_{k}^{\prime}+2$
    $\geq f_{1}\left(A_{n}\right)+2.$
Now we can prove the lower bound $12n-$ 1 for $A_{n}$ in the proposition by induction on $n$:
Lemma 4.2 implies that $f_{1}\left(A_{2}\right)\geq 3=2\cdot 2-1$. And the assumption $f_{1}\left(A_{n}\right)\geq 2n-1$ implies that either $f_{1}\left(A_{n+1}\right)\geq 2n+1$ or $f_{1}\left(A_{n+1}\right)\geq f_{1}\left(A_{n}\right)+2\geq 2n-1+2=2n+1$. Thus, by induction we have $f_{1}\left(A_{n}\right)\geq 2n-1$. And then it is a repeated argument to show $f_{1}\left(\Phi_{n}\right)\geq 2n-1$.
4.2. Case (ii)
$\mathrm{rank}(\zeta)=2$. By this condition, one of the $\zeta^{i}$s can be represented as a linear combination of the other two. Without loss of generality, assume $\zeta^{1}\in\mathrm{span}\{\zeta^{2},\zeta^{3}\}$. Recall that the period map is given by:
$\widetilde{p_{\zeta}}:\mathbb{C}^{2}$ $\to\mathfrak{h}^{\mathbb{C}},$
$(z_{1},z_{2})$ $\mapsto z_{1}^{2}[\omega^{2}_{\zeta}+\mathrm{i}\omega^{3}_{\zeta}]+2z_{1}z_{2}\omega_{\zeta}^{1}-z_{2}^{2}[\omega^{2}_{\zeta}-\mathrm{i}\omega^{3}_{\zeta}]$
$\quad=z_{1}^{2}(\zeta^{2}+\mathrm{i}\zeta^{3})+2z_{1}z_{2}\zeta^{1}-z_{2}^{2}(\zeta^{2}-\mathrm{i}\zeta^{3}).$
Since $\zeta^{1}\in\mathrm{span}\{\zeta^{2},\zeta^{3}\}$, there exist complex numbers $c,\bar{c}$ such that $\zeta^{1}=c\left(\zeta^{2}+\mathrm{i}\zeta^{3}\right)+\bar{c}\left(\zeta^{2}-\mathrm{i}\zeta^{3}\right)$. It is straightforward to verify that $\mathrm{Im}\,\widetilde{p_{\zeta}}$ is a two-dimensional linear subspace in $\mathfrak{h}^{\mathbb{C}}$. After choosing a basis for $\mathrm{Im}\,\widetilde{p_{\zeta}}$ as $\left\{\left(\zeta^{2}+\mathrm{i}\zeta^{3}\right),\left(\zeta^{2}-\mathrm{i}\zeta^{3}\right)\right\}$, we obtain the following map for $\widetilde{p_{\zeta}}$:
$p_{\zeta}^{\prime}:\mathbb{P}(\mathbb{C}^{2})\simeq\mathbb{P}^{1}\xrightarrow{\quad}$ $\,\mathbb{P}^{1}\simeq\mathbb{P}(\mathrm{Im}\,\widetilde{p_{\zeta}})$
$[z_{1}:z_{2}]\,\mapsto\,$ $\,[w_{1}:w_{2}]:=\left[z_{1}^{2}+2cz_{1}z_{2}:-z_{2}^{2}+2\bar{c}z_{1}z_{2}\right].$
This $p_{\zeta}^{\prime}$ is a degree two branched covering between Riemann spheres, with exactly two branched points. Each one-dimensional subspace in $\mathrm{Im}\,\widetilde{p_{\zeta}}$ can be represented by a point $[w_{1}:w_{2}]$ in its projective space $\mathbb{P}(\mathrm{Im}\,\widetilde{p_{\zeta}})$. Since the map $p_{\zeta}^{\prime}$ is of degree two, each point in $\mathbb{P}(\mathrm{Im}\,\widetilde{p_{\zeta}})$ has preimage with two points in $\mathbb{P}(\mathbb{C}^{2})$. The two points are distinct if $[w_{1}:w_{2}]$ is not a branched value, and coincide if it is.
Since $\mathrm{rank}(\zeta)=2$, $\mathrm{Im}\,\widetilde{p_{\zeta}}$ is a two-dimensional linear subspace in $\mathfrak{h}^{\mathbb{C}}$. According to the argument in [9, p.693-694], which is true for any $\zeta\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$, the image of the period map $\widetilde{p_{\zeta}}$ of the twistor space $\mathrm{Z}(X_{\zeta})$ cannot be entirely contained in the kernel of any root. Therefore, the two-dimensional subspace $\mathrm{Im}\,\widetilde{p_{\zeta}}$ satisfies the condition for $L$ as discussed in subsection 4.1. We thus set $L:=\mathrm{Im}\,\widetilde{p_{\zeta}}$, which, as detailed in subsection 4.1, induces the following decompositions:
$\Phi$ $=\Phi_{1}\sqcup\Phi_{2}\sqcup\cdots\sqcup\Phi_{s},$
$\Delta$ $=\Delta_{1}\sqcup\Delta_{2}\sqcup\cdots\sqcup\Delta_{s},$
where $s\geq 3$ and $2r-1\leq\sum_{k=1}^{s}r_{k}\leq\frac{1}{2}\sharp\Phi$. We denote $l_{k}$, $k=1,\ldots,s$, as the one-dimensional subspaces in $L$ associated with $\Phi_{k}$, as in subsection 4.1. According to Corollary 2.14, the equation (1.4) reads:
$\sharp\left\{u\in\mathbb{P}^{1}\mid\pi^{-1}(u)\text{ contains a rational curve}\right\}=\sharp\left\{(p_{\zeta}^{\prime})^{-1}\{[l_{1}],\ldots,[l_{s}]\}\right\}.$
The upper bound $\sharp\Phi$ has already been established in [9, Lemma 2.7]. We now proceed to establish the lower bound.
Each one-dimensional subspace $l_{k}$ , for $k=1,\ldots,s$, corresponds to two points $(p_{\zeta}^{\prime})^{-1}[l_{k}]=\{u_{k}^{1},u_{k}^{2}\}$ in the hyperKählerian family of complex structures $\mathbb{P}^{1}$, where $u_{k}^{1}=u_{k}^{2}$ if and only if $[l_{k}]$ is $[1:0]$ or $[0:1]$. Since $s\geq 3$, we have:
(4.4) $\sharp\left\{(p_{\zeta}^{\prime})^{-1}\{[l_{1}],\ldots,[l_{s}]\}\right\}\geq 1+1+2=4.$
Thus, the desired lower bound of equation (1.4) is established. Each complex manifold $\left(X_{\zeta},u_{k}^{i}\right)$, for $i=1,2$, is the minimal resolution of the fiber of $\tilde{f}$ over any point of $l_{\theta_{k}}\setminus\{0\}\subset\mathrm{Im}(p_{\zeta})\subset\mathfrak{h}^{\mathbb{C}}$ in the deformation $\tilde{\mathcal{Y}}$, as depicted in diagram 2.1. By Corollary 2.11, the fiber of $\tilde{f}$ over any nonzero point of $l_{\theta_{k}}$ contains $r_{k}(=\mathrm{rank}\,\Phi_{k})$ rational curves. Hence, inequality (1.5) follows from Proposition 4.5:
$2r\leq$ $\sum_{k=1}^{s}r_{k}\cdot\sharp\{u_{k}^{1},u_{k}^{2}\}=\sum_{k=1}^{s}\sharp\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}\{u_{k}^{1},u_{k}^{2}\}\right\}$
$=$ $\sum_{u\in\mathbb{P}^{1}}\sharp\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}(u)\right\}\leq\sharp\Phi.$
So far, Case (ii) in Theorem 1.2 has been proved.
4.3. Case (iii)
$\mathrm{rank}(\zeta)=3$. We first describe the strategy of our proof. In this case, $\mathrm{Im}\,\widetilde{p_{\zeta}}$ is a hypersurface in the three-dimensional space $V_{\zeta}:=\mathrm{span}_{\mathbb{C}}\left\{[\omega_{\zeta}^{1}],[\omega_{\zeta}^{2}],[\omega_{\zeta}^{3}]\right\}=\mathrm{span}_{\mathbb{C}}\left\{\zeta^{1},\zeta^{2},\zeta^{3}\right\}\subset\mathfrak{h}^{\mathbb{C}}$. If we choose $\left\{\left(\zeta^{2}+\mathrm{i}\zeta^{3}\right),2\zeta^{1},\left(-\zeta^{2}+\mathrm{i}\zeta^{3}\right)\right\}$ as a basis for $V_{\zeta}$, then the period map $\widetilde{p_{\zeta}}$ induces an bijective map denoted by:
$p_{\zeta}^{\prime}:\mathbb{P}(\mathbb{C}^{2})\simeq\mathbb{P}^{1}\xrightarrow{1:1}$ $\mathbb{P}(\mathrm{Im}\,\widetilde{p_{\zeta}})\subset\mathbb{P}(V_{\zeta})\simeq\mathbb{P}^{2}$
$[z_{1}:z_{2}]\mapsto$ $[x:y:z]:=[z_{1}^{2}:z_{1}z_{2}:z_{2}^{2}].$
$\mathbb{P}(\mathrm{Im}\,\widetilde{p_{\zeta}})\subset\mathbb{P}\left({V_{\zeta}}\right)$ is a smooth conic given by the equation $\{xy=z^{2}\}\subset\mathbb{P}^{2}$. Denote this conic by $C$ for simplicity. Since the image $\mathrm{Im}\,\widetilde{p_{\zeta}}$ and then $V_{\zeta}$ cannot lie entirely in kernel of any root $\theta\in\Phi$, we define $H_{\theta}$ to be the kernel of $\theta$ restricted to $V_{\zeta}$, it is a two-dimensional subspace and $\mathbb{P}H_{\theta}\simeq\mathbb{P}^{1}$ is a projective line in $\mathbb{P}\left({V_{\zeta}}\right)$. Each projective line $\mathbb{P}H_{\theta}$ intersects $C$ either transversely at two distinct points or tangentially at one point. We call an intersection point $\gamma$ a tangential point if any $\mathbb{P}H_{\theta}$, $\theta\in\Phi$, passing through $\gamma$ tangents to $C$ at $\gamma$. Otherwise, we call an intersection point $\gamma$ a transversal point, i.e., there exists some root $\theta\in\Phi$ such that the projective line $\mathbb{P}H_{\theta}$ intersects $C$ transversely at $\gamma$. Denote the set of all intersection points by:
$\bigcup_{\theta\in\Phi}\,\,\mathbb{P}H_{\theta}\cap C=\{\gamma_{1},\ldots,\gamma_{s},\gamma_{s+1},\ldots,\gamma_{s+t}\}$
where $\gamma_{k}$ and $\gamma_{s+j}$ (with indices $k=1,\ldots,s$; $j=1,\ldots,t$) are distinct points on $C$, and for each $k=1,\ldots,s$, the point $\gamma_{k}$ is a transversal point, and for each $j=1,\ldots,t$, the point $\gamma_{s+j}$ is a tangential point. Hence, the integer $s$ or $t$ is the number of transversal points or the tangential points respectively. Note that $s$ or $t$ could be zero at this stage. By the injectivity of $p_{\zeta}^{\prime}$, each point $\gamma_{k}$ or $\gamma_{s+j}$ in $C$ has a unique preimage, which corresponds to a complex structure in the hyperKählerian family of $X_{\zeta}$. Denote the preimage of $\gamma_{k}$ by $u_{k}$ and the preimage of $\gamma_{s+j}$ by $u_{s+j}$. Then equation (1.6) reads:
$s+t=\sharp\,\{u_{1},\cdots,u_{s+t}\}=\sharp\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\text{ contains a rational curve}\right\}$
We obtain the following decomposition (which are not necessarily all disjoint unions):
(4.5) $\Phi$ $=\Phi_{1}\cup\cdots\cup\Phi_{s}\cup\Phi_{s+1}\cup\cdots\cup\Phi_{s+t},$
$\Delta$ $=\Delta_{1}\cup\cdots\cup\Delta_{s}\cup\Delta_{s+1}\cup\cdots\cup\Delta_{s+t},$
where for indices $k=1,\ldots,s$; $j=1,\ldots,t$, we define:
$\Phi_{k}$ :=$:=\{\theta\in\Phi\mid\gamma_{k}\in\mathbb{P}H_{\theta}\cap C\}\,,\quad\Delta_{k}:=\Phi_{k}\cap\Delta;$
$\Phi_{s+j}$ :=$:=\{\theta\in\Phi\mid\gamma_{s+j}\in\mathbb{P}H_{\theta}\cap C\}\,,\quad\Delta_{s+j}:=\Phi_{s+j}\cap\Delta.$
Each set $\Phi_{k}$ or $\Phi_{s+j}$ is a root subsystem and satisfies the following conditions:
(4.6) $\Phi\bigcap\mathrm{span}_{\mathbb{R}}\Phi_{k}=\Phi_{k}\,,\text{ or}\quad\Phi\bigcap\mathrm{span}_{\mathbb{R}}\Phi_{s+j}=\Phi_{s+j}.$
The following is a basic fact:
Lemma 4.7.
The root subsystems $\Phi_{s+j}$, for $j=1,\cdots,t$, are mutually disjoint. The set $\bigcup_{k=1}^{s}\Phi_{k}$ is disjoint from the set $\bigcup_{j=1}^{t}\Phi_{s+j}$.
Proof.
For any $j\in\{1,\cdots,t\}$, by the definition of $\gamma_{s+j}$, any line $\mathbb{P}H_{\theta}$ through the point $\gamma_{s+j}$ must be tangent to $\mathbb{P}(\mathrm{Im}\,p_{\zeta})$ at $\gamma_{s+j}$. Since $C$ is a conic curve, the line $\mathbb{P}H_{\theta}$ intersects $C$ only at $\gamma_{s+j}$. Therefore, the root subsystems $\Phi_{s+j}$, for $j=1,\cdots,t$, are mutually disjoint. Moreover, a line $\mathbb{P}H_{\theta}$ passing through a point $\gamma_{s+j}$ cannot intersect any other point $\gamma_{k}$, for $k=1,\cdots,s$. Thus, the set $\bigcup_{k=1}^{s}\Phi_{k}$ is disjoint from $\bigcup_{j=1}^{t}\Phi_{s+j}$. ∎
Denote $r_{k}:=\mathrm{rank}\,\Phi_{k}$, $\delta_{k}=\sharp\Delta_{k}$, and $r_{s+j}:=\mathrm{rank}\,\Phi_{s+j}$, $\delta_{s+j}=\sharp\Delta_{s+j}$. Then the equation (1.7) reads:
$\sum_{k=1}^{s}r_{k}+\sum_{j=1}^{t}r_{s+j}=\sum_{u\in\mathbb{P}^{1}}\sharp\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}(u)\right\}.$
The upper bound for equation (1.7) is easy to see: we can further decompose (4.5) to a disjoint union
$\Phi_{1}\cup\cdots\cup\Phi_{s}=\Phi_{11}\sqcup\cdots\sqcup\Phi_{ss}\sqcup\cdots\sqcup\Phi_{k_{1}k_{2}}\sqcup\cdots.$
where
$\Phi_{kk}$ :=$:=\left\{\theta\in\Phi\mid\mathbb{P}H_{\theta}\text{ tangents at }\gamma_{k}\right\}$ $k=1,\cdots,s$
$\Phi_{k_{1}k_{2}}$ :=$:=\left\{\theta\in\Phi\mid\mathbb{P}H_{\theta}\text{pass through }\gamma_{k_{1}}\text{ and }\gamma_{k_{2}}\right\}$ $1\leq k_{1}<k_{2}\leq s.$
Note that $r_{k}\leq r_{kk}+\sum_{k\neq k^{\prime}}r_{kk^{\prime}}$, hence $\sum_{k=1}^{s}r_{k}\leq 2\sum_{k,k^{\prime}}r_{kk^{\prime}}$. Then
$\sum_{k=1}^{s}r_{k}+\sum_{j=1}^{t}r_{s+j}\leq 2\sum_{k,k^{\prime}}r_{kk^{\prime}}+\sum_{j=1}^{t}r_{s+j}\leq 2\left(\sum_{k,k^{\prime}}r_{kk^{\prime}}+\sum_{j=1}^{t}r_{s+j}\right)\leq\sharp\Phi,$
the last inequality comes from the upper bound in Lemma 4.2 since it is arised as disjoint union of root subsystems. This argument also apply to the upper bound of equation (1.6) which already has been proved by Kronheimer [9].
We will obtain the lower bound $3$ for equation (1.6) in Theorem 1.2 by dividing the problem into three cases: (1) $s=0$, i.e., all the intersection points are tangential point; (2) $t=0$, i.e., all the intersection points are transversal point; (3) $s\neq 0$ and $t\neq 0$, i.e., there exist a transversal point and a tangential point. After discussing these three cases separately, in the end of subsection 4.3, we introduce another type of decomposition of $\Phi$ which is designed as an analogue of the one in subsection 4.1 to obtain the lower bound $12r-$ 1 of equation (1.7).
4.3.1. $s=0$
We show that this is impossible.
Lemma 4.8.
Given the decomposition (4.5), the number $s$ cannot be zero.
Proof.
Suppose $s=0$. Then, according to Lemma 4.7, the decomposition is a disjoint union:
$\Phi$ $=\Phi_{0+1}\sqcup\cdots\sqcup\Phi_{0+t},$
$\Delta$ $=\Delta_{0+1}\sqcup\cdots\sqcup\Delta_{0+t}.$
The proof of Lemma 4.1 applies to this case, which implies that $t\geq 3$. Then by the irreducibility of $\Phi$, we can assume without loss of generality that $\Phi_{0+1}\neq\emptyset$, $\Phi_{0+2}\neq\emptyset$, and that $(\Phi_{0+1},\Phi_{0+2})\neq 0$. Hence, there exist roots $\alpha\in\Phi_{0+1}$ and $\beta\in\Phi_{0+2}$ such that either $\alpha+\beta\in\Phi$ or $\alpha-\beta\in\Phi$. Assume $\alpha+\beta\in\Phi_{0+3}$. Since $\mathbb{P}H_{\alpha}\cap\mathbb{P}H_{\beta}=\mathbb{P}(H_{\alpha}\cap H_{\beta})$, we have:
$\mathbb{P}H_{\alpha}\cap\mathbb{P}H_{\beta}=\mathbb{P}H_{\alpha}\cap\mathbb{P}H_{\alpha+\beta}=\mathbb{P}H_{\beta}\cap\mathbb{P}H_{\alpha+\beta}.$
This implies that the three distinct lines $\mathbb{P}H_{\alpha}$, $\mathbb{P}H_{\beta}$, and $\mathbb{P}H_{\alpha+\beta}$ intersect at a common point in the projective plane $\mathbb{P}(V_{\zeta})$. However, by definition, these lines are all tangent to the conic curve $C$. Thus, we obtain an impossible configuration. Therefore, $s\neq 0$. ∎
4.3.2. $t=0$
We have the following lemma:
Lemma 4.9.
Given the decomposition (4.5), if $t=0$, then $s\geq 3$.
Proof.
Suppose $t=0$. Then $s\neq 0$. Since $s$ is the number of transversal points, it must hold that $s\geq 2$. For the sake of contradiction, assume $s=2$, and we can write the set of all intersection points as $\left\{\gamma_{1},\gamma_{2}\right\}$. An intersection point of $C$ with any line $\mathbb{P}H_{\theta},\theta\in\Phi$ can only be $\gamma_{1}$ or $\gamma_{2}$. The decompositions of $\Phi$ and $\Delta$ are:
$\Phi$ $=\Phi_{1}\cup\Phi_{2},$
$\Delta$ $=\Delta_{1}\cup\Delta_{2}.$
Moreover, $\Phi$ and $\Delta$ can be decomposed further into three disjoint root subsystems:
$\Phi$ $=\Phi_{1}^{\prime}\sqcup\Phi_{2}^{\prime}\sqcup\Phi_{3}^{\prime},$
$\Delta$ $=\Delta_{1}^{\prime}\sqcup\Delta_{2}^{\prime}\sqcup\Delta_{3}^{\prime},$
where $\Phi_{1}^{\prime}$ contains all the $\theta$ which the line $\mathbb{P}H_{\theta}$ tangents to $C$ at $\gamma_{1}$, $\Phi_{2}^{\prime}$ contains all the $\theta$ which the line $\mathbb{P}H_{\theta}$ tangents to $C$ at $\gamma_{2}$, and $\Phi_{3}^{\prime}$ contains all the $\theta$ which the line $\mathbb{P}H_{\theta}$ pass through both $\gamma_{1}$ and $\gamma_{2}$. By definition, $\Phi_{1}^{\prime}$, $\Phi_{2}^{\prime}$, and $\Phi_{3}^{\prime}$ are all root subsystems of $\Phi$, and $\Phi_{3}^{\prime}\neq\emptyset$. Since $\bigcap_{\theta\in\Phi}H_{\theta}=0$, there hold $\Phi_{1}^{\prime}\neq\emptyset$ and $\Phi_{2}^{\prime}\neq\emptyset$. By the irreducibility of $\Phi$, the three root subsystem $\Phi_{i}^{\prime}$, $i=1,2,3$, can not mutually orthogonal, without loss of generality, we assume $\alpha\in\Phi_{1}^{\prime}$ and $\beta\in\Phi_{2}^{\prime}$ such that $\alpha+\beta\in\Phi$. Then, $\alpha+\beta\in\Phi_{3}^{\prime}$. We use again the intersection properties:
$\mathbb{P}H_{\alpha}\cap\mathbb{P}H_{\alpha+\beta}=\mathbb{P}H_{\beta}\cap\mathbb{P}H_{\alpha+\beta}=\mathbb{P}H_{\alpha}\cap\mathbb{P}H_{\beta}.$
This implies that the three distinct lines $\mathbb{P}H_{\alpha}$, $\mathbb{P}H_{\beta}$, and $\mathbb{P}H_{\alpha+\beta}$ intersect at a common point in $\mathbb{P}(V_{\zeta})$. However, $\mathbb{P}H_{\alpha}$ is tangent to $\mathbb{P}(\mathrm{Im}\widetilde{p_{\zeta}})$ at $\gamma_{1}$, $\mathbb{P}H_{\beta}$ is tangent to $\mathbb{P}(\mathrm{Im}\widetilde{p_{\zeta}})$ at $\gamma_{2}$, and $\mathbb{P}H_{\alpha+\beta}$ intersects transversely at both $\gamma_{1}$ and $\gamma_{2}$. This configuration is impossible in the complex plane $\mathbb{P}^{2}$. Thus $s\neq 2$.
Hence, we can obtain the lower bound of equation (1.6) in the case of $t=0$:
$3\leq s=\sharp\,\{u_{k},k=1,\cdots,s\}=\sharp\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\text{ contains a rational curve}\right\}$
4.3.3. $s\neq 0$ and $t\neq 0$
This means $s\geq 2$ and $t\geq 1$. We obtain immediately the estimate of (1.6):
$3\leq s+t=$$\sharp\,\{u_{k},u_{s+j};k=1,\cdots,s\,,\,j=1,\cdots,t\}=\sharp\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\text{ contains a rational curve}\right\}.$
Moreover we can prove:
Lemma 4.10.
Given the decomposition (4.5), if $s=2$, then $t\geq 2$.
Proof.
Suppose $s=2$. By Lemma 4.9, $t\neq 0$.
We first show that the union $\Phi_{1}\cup\Phi_{2}$ forms a root subsystem of $\Phi$. To establish this, we only need to verify that for any $\alpha,\beta\in\Phi_{1}\cup\Phi_{2}$, if $\alpha+m\beta\in\Phi$ for some integer $m$, then $\alpha+m\beta\in\Phi_{1}\cup\Phi_{2}$. We prove this by contradiction. Suppose there exist $\alpha,\beta\in\Phi_{1}\cup\Phi_{2}$ and an integer $m$ such that $\alpha+m\beta\in\Phi$ but $\alpha+m\beta\notin\Phi_{1}\cup\Phi_{2}$. Thus, $\alpha+m\beta$ must belong to $\Phi_{2+j_{0}}$ for some $j_{0}\in\{1,\dots,t\}$. Since $\Phi_{1}\cap\Phi_{2}$ is a root subsystem, $\alpha$ and $\beta$ cannot both belong to $\Phi_{1}\cap\Phi_{2}$. Since we can exchange $\alpha$ and $\beta$ , without loss of generality, we can only consider the following two types of the positions of $\alpha$ and $\beta$:
  • (P1)
    $\alpha\in\Phi_{1}\setminus\Phi_{2}$ and $\beta\in\Phi_{2}\setminus\Phi_{1}$. This case is impossible because it results in an impossible configuration where the lines $\mathbb{P}H_{\alpha}$, $\mathbb{P}H_{\beta}$, and $\mathbb{P}H_{\alpha+\beta}$ intersect at a common point, and they are all tangent to $C$ in projective plane.
  • (P2)
    $\alpha\in\Phi_{1}\cap\Phi_{2}$ and $\beta\in\Phi_{2}\setminus\Phi_{1}$. This case is impossible because the line $\mathbb{P}H_{\alpha}$ passes through two points $\gamma_{1}$ and $\gamma_{2}$, while the line $\mathbb{P}H_{\beta}$ is tangent to $C$ at point $\gamma_{2}$. Consequently, the line $\mathbb{P}H_{\alpha+m\beta}$ must also pass through $\gamma_{2}$ which is a transversal point. This contradicts the condition that $\alpha+m\beta\in\Phi_{2+j_{0}}$.
Since all cases lead to contradictions, $\alpha+m\beta$ must lie in $\Phi_{1}\cup\Phi_{2}$. Hence, $\Phi_{1}\cup\Phi_{2}$ is a root subsystem. Since both $\Phi_{1}$ and $\Phi_{2}$ satisfy the condition (4.6), there holds:
(4.7) $\Phi\bigcap\,\,\mathrm{span}_{\mathbb{R}}\left\{\Phi_{1}\cup\Phi_{2}\right\}=\Phi_{1}\cup\Phi_{2}.$
Now, we prove $t\geq 2$. Assume for contradiction that $t=1$. Then the decomposition reads:
$\Phi$ $=\Phi_{1}\cup\Phi_{2}\sqcup\Phi_{2+1},$
$\Delta$ $=\Delta_{1}\cup\Delta_{2}\sqcup\Delta_{2+1}.$
Since $\Phi_{1}\cup\Phi_{2}$ is a root subsystem, the irreducibility of $\Phi$ implies that $\Phi_{1}\cup\Phi_{2}$ cannot be orthogonal to $\Phi_{2+1}$. Therefore, there exist $\alpha\in\Phi_{1}\cup\Phi_{2}$ and $\theta\in\Phi_{2+1}$ such that $(\alpha,\theta)\neq 0$. Assume $\alpha+\theta\in\Phi$. Since the conditions (4.6) and (4.7) hold, we have that $\alpha+\theta$ cannot belong to either $\Phi_{1}\cup\Phi_{2}$ or $\Phi_{2+1}$. Thus, we obtain a contradiction. Therefore, $t\geq 2$.
So far, we have proved that if $\mathrm{rank}(\zeta)=3$, then
$3\leq\sharp\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}.$
Now we aim to obtain the lower bound of equation (1.7):
$2r-1\leq\sum_{k=1}^{s}r_{k}+\sum_{j=1}^{t}r_{s+j}\leq\sum_{u\in\mathbb{P}^{1}}\sharp\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}(u)\right\}.$
For the purpose of producing the lower bound $12r-$ 1, we introduce another space of decompositions of the irreducible root system $\Phi$ as follows.
Definition 4.11.
We call an expression $\Phi=\Phi_{1}\cup\cdots\cup\Phi_{s}\sqcup\Phi_{s+1}\sqcup\cdots\sqcup\Phi_{s+t}$ a decomposition of type-$2$ of $\Phi$ if the following conditions hold:
  • (1)
    $s\geq 0$, $t\geq 0$ and $s+t\neq 1$.
  • (2)
    For each $i\in\left\{1,\cdots,s+t\right\}$, the subset $\Phi_{i}$ is a non-empty closed root subsystem of $\Phi$, where closed means:
    $\Phi_{i}=\Phi\cap\mathrm{span}_{\mathbb{R}}\,\Phi_{i}.$
  • (3)
    The root subsystems $\Phi_{s+j}$, for $j=1,\cdots,t$, are mutually disjoint. The set $\bigcup_{k=1}^{s}\Phi_{k}$ is disjoint from the set $\bigcup_{j=1}^{t}\Phi_{s+j}$.
Denote $\mathfrak{D}^{2}\left(\Phi\right)$ the space of all decompositions of type-$2$ of $\Phi$.
For any decomposition $\Phi=\Phi_{1}\cup\cdots\cup\Phi_{s}\sqcup\Phi_{s+1}\sqcup\cdots\sqcup\Phi_{s+t}$ of type-$2$, define $r_{i}:=\mathrm{rank}\,\Phi_{i}$. Denote $f_{2}\left(\Phi\right)$ as the infimum of the sum $\sum_{i=1}^{s+t}r_{i}$ for all the decompositions in $\mathfrak{D}^{2}\left(\Phi\right)$, i.e.,
$f_{2}\left(\Phi\right):=\inf_{\mathfrak{D}^{2}\left(\Phi\right)}\left\{\sum_{i=1}^{s+t}r_{i}\right\}.$
Similar to Lemma 4.2, condition (1) together with the existence of the maximal root guarantees $f_{2}\left(\Phi\right)\geq r+1$.
Remark 4.12.
Obviously, for $r=\mathrm{rank}\,\Phi\geq 3$, the decomposition (4.5) induced by the conic curve $C$ is a decomposition of type-$2$ of $\Phi$. Hence $f_{2}\left(\Phi\right)$ is a lower bound for equation (1.7). If $s=0$, then the decomposition is also of type-$1$, i.e., $\mathfrak{D}^{1}\left(\Phi\right)\subset\mathfrak{D}^{2}\left(\Phi\right)$. However, the space $\mathfrak{D}^{2}\left(\Phi\right)$ is too coarse to produce a better lower bound for equation (1.6).
Similar with Proposition 4.5, we have:
Proposition 4.13.
For $\Phi(n)=A_{n\geq 2},\,D_{n\geq 3}$ or $E_{6},E_{7},E_{8}$, i.e., $\Phi(n)$ is the root system of type $ADE$ with rank $r\geq 2$, we have estimates:
$2n-1\leq f_{2}(\Phi(n)).$
Remark 4.14.
Note that Lemma 4.6 is also valid in this situation since it is a property of root systems alone. And the entire proof of this proposition is actually an anologue of Proposition 4.5.
Proof.
Since $\sum_{i=1}^{s+t}r_{i}\in\mathbb{Z}_{>0}$, the infimum can always be reached. Take a decomposition with type-$2$ of $\Phi(n+1)$:
$\Phi(n+1)=\Phi_{1}\cup\cdots\cup\Phi_{s}\sqcup\Phi_{s+1}\sqcup\cdots\sqcup\Phi_{s+t}$
such that $\sum_{i=1}^{s+t}r_{i}=f_{2}\left(\Phi(n+1)\right)$. By deleting the set of extra roots $S$ of $\Phi(n+1)$ with respect to $A_{n}$, we obtain
(4.8) $A_{n}=\Phi_{1}^{\prime}\cup\cdots\cup\Phi_{s}^{\prime}\sqcup\Phi_{s+1}^{\prime}\sqcup\cdots\sqcup\Phi_{s+t}^{\prime}\quad,\quad\Phi_{k}^{\prime}:=\Phi_{k}\setminus S\,,\,\Phi_{s+j}^{\prime}:=\Phi_{s+j}\setminus S.$
Each $\Phi_{i}^{\prime}$ is either empty or a root subsystem of $A_{n}$. Then there are two possibilities:
  • There exits $i_{0}\in\left\{1,\cdots,s+t\right\}$ such that $\Phi_{i_{0}}^{\prime}=A_{n}$. Then the extra roots are contained in $\bigcup_{i\neq i_{0}}\Phi_{i}$, applying Lemma 4.6 we get:
    $f_{2}\left(\Phi(n+1)\right)\geq r_{i_{0}}+\sum_{i\neq i_{0}}r_{i}\geq n+n+1=2n+1.$
  • Each $\Phi_{i}^{\prime},i=1,\cdots,s+t$ is a proper root subsystem of $A_{n}$. Then it’s straight forward to verity the expression (4.8) after ignoring the empty sets is a decomposition with type-$2$ of $A_{n}$. Since the extra roots cannot be put into a same root subsystem, we have
    (4.9) $f_{2}\left(\Phi(n+1)\right)\geq f_{2}\left(A_{n}\right)+2.$
Now we can prove the lower bound $12n-$ 1 for $\Phi(n)$ by induction on $n$:
Since $f_{2}\left(A_{3}\right)\geq 2+1=3=2\cdot 2-1$. And the assumption $f_{2}\left(A_{n}\right)\geq 2n-1$ implies that either $f_{2}\left(A_{n+1}\right)\geq 2n+1$ or $f_{2}\left(A_{n}\right)\geq f_{2}\left(A_{n}\right)+2\geq 2n-1+2=2n+1$. Thus by induction we have $f_{2}\left(A_{n}\right)\geq 2n-1$. Then by repeated arguments we have $f_{2}\left(\Phi(n)\right)\geq 2n-1$. ∎
Thus we obtain the desired lower bound of equation (1.7):
$2r-1\leq f_{2}(\Phi)\leq\sum_{k=1}^{s}r_{k}+\sum_{j=1}^{t}r_{s+j}\leq\sum_{u\in\mathbb{P}^{1}}\sharp\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}(u)\right\}.$
Finally, we complete the proof of Case (iii) in Theorem 1.2.
5. A Semi-Continuity
Recall the map $Q_{1}$ defined in the introduction:
$Q_{1}:\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$ $\to\mathbb{Z}$
$\zeta$$\mapsto\sharp\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}$
In this section we will prove Theorem 1.4, i.e.
$\liminf_{\zeta\to\zeta_{0}}Q_{1}(\zeta)\geq Q_{1}(\zeta_{0})\quad,\,\,\forall\zeta_{0}\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}.$
Proof.
We divide our proof into three cases depending on $\text{rank}(\zeta_{0})$:
  • (1)
    $\text{rank}(\zeta_{0})=1$. By Theorem 1.2, $Q_{1}(\zeta_{0})=2\leq Q_{1}(\zeta)$ $\forall\zeta$.
  • (2)
    $\text{rank}(\zeta_{0})=2$. Since rank is lower semi-continue function, the set $\{\zeta\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}\mid\text{rank}(\zeta)>1\}$ is an open set. Hence, there exists $\delta_{0}>0$ such that $\forall\,\,\zeta$ with $|\zeta-\zeta_{0}|<\delta_{0}$, $\text{rank}(\zeta)\geq 2$.
    • if $\text{rank}(\zeta)=2$. From subsection 4.2, there are two plane $L_{\zeta}$ and $L_{\zeta_{0}}$ in $\mathfrak{h}^{\mathbb{C}}$ which give us two decompositions of the root system $\Phi$:
      $\Phi$ $=\Phi_{1}\sqcup\cdots\sqcup\Phi_{s_{\zeta}}$
      $\Phi$ $=\Phi_{1}^{\prime}\sqcup\cdots\sqcup\Phi_{s_{\zeta_{0}}}^{\prime}$
      and then
      $Q_{1}(\zeta_{0})$ $=\sum_{k=1}^{s_{\zeta_{0}}}\sharp\,\{u_{k}^{1},u_{k}^{2}\}$
      $Q_{1}(\zeta)$ $=\sum_{j=1}^{s_{\zeta}}\sharp\,\{v_{j}^{1},v_{j}^{2}\}$
      First, we claim $s_{\zeta_{0}}\leq s_{\zeta}$ for $\zeta$ sufficiently closed to $\zeta_{0}$ : $\forall\,\,k=1,\cdots,s_{\zeta_{0}}$, we take $\alpha_{k}\in\Phi_{k}$. We choose for $k=1,\cdots,s_{\zeta_{0}}$ a number $m(k)\in\{1,\cdots,s_{\zeta}\}$ such that $\alpha_{k}\in\Phi^{\prime}_{m(k)}$. We prove by contraction that the map $m:k\mapsto m(k)$ is injective: Assume there exists $(k_{1},k_{2})\in\{1,\cdots,s_{\zeta_{0}}\}\crossproduct\{1,\cdots,s_{\zeta_{0}}\}$ such that $k_{1}\neq k_{2}$ and $m(k_{1})=m(k_{2})$, i.e. $\begin{cases}\ker\alpha_{k_{1}}\,\cap\,L_{\zeta_{0}}\neq\ker\alpha_{k_{2}}\,\cap\,L_{\zeta_{0}},\\ \ker\alpha_{k_{1}}\,\cap\,L_{\zeta}=\ker\alpha_{k_{2}}\,\cap\,L_{\zeta}.\end{cases}$. As $\zeta\to\zeta_{0}$, $L_{\zeta}\to L_{\zeta_{0}}$. Hence $\ker\alpha_{k_{i}}\,\cap L_{\zeta}\to\ker\alpha_{k_{i}}\,\cap L_{\zeta_{0}}$ for $i=1,2$ which implies $\ker\alpha_{k_{1}}\,\cap L_{\zeta_{0}}=\ker\alpha_{k_{2}}\,\cap\,L_{\zeta_{0}}$, which is a contraction. Therefore $m:k\mapsto m(k)$ is injective and then $s_{\zeta_{0}}\leq s_{\zeta}$.
[…]
      $Q_{1}(\zeta_{0})=\sum_{k}^{s_{\zeta_{0}}}\sharp\,\{u_{k}^{1},u_{k}^{2}\}\leq\sum_{j=1}^{s_{\zeta}}\sharp\,\{v_{j}^{1},v_{j}^{2}\}=Q_{1}(\zeta).$
    • If $\text{rank}(\zeta)=3$. Denote $C_{\zeta}$ the projective of the image of period map $\tilde{p_{\zeta}}$ which is a conic curve in $\mathbb{P}(\mathfrak{h}^{\mathbb{C}})$ as in subsection 4.3. Then as $\zeta\to\zeta_{0}$, $C_{\zeta}\to\mathbb{P}(L_{\zeta_{0}})\subset\mathbb{P}(\mathfrak{h}^{\mathbb{C}})$. We claim that: $\forall\,\alpha\in\Phi$, $\exists\,\epsilon(\alpha)>0$ such that $\forall\,\zeta$ with $|\zeta-\zeta_{0}|<\epsilon(\alpha)$, the projective hyperplane $\mathbb{P}(\ker\alpha)$ is not tangent to $C_{\zeta}$. We prove this claim by contradiction: if $\exists\,\alpha_{0}\in\Phi$ such that $\forall\,\epsilon>0$, $\exists\,\zeta(\epsilon)$ with $|\zeta(\epsilon)-\zeta_{0}|<\epsilon$, the $\mathbb{P}(\ker\alpha_{0})$ is tangent to $C_{\zeta(\epsilon)}$. Then as $\epsilon\to 0$, $\zeta(\epsilon)\to\zeta$, $C_{\zeta(\epsilon)}\to\mathbb{P}(L_{\zeta_{0}})$. Hence $\mathbb{P}(\ker\alpha_{0})$ tangents to $\mathbb{P}(L_{\zeta_{0}})$. This implies $\mathbb{P}(L_{\zeta_{0}})=\mathbb{P}(\ker\alpha_{0})$. This contradicts to the fact that the image of period map is not contained in any $\ker\alpha$, see subsection 4.2.
      Hence if we take $\epsilon<\min\{\epsilon(\alpha),\alpha\in\Phi\}$, and $\zeta$ with $|\zeta-\zeta_{0}|<\epsilon$, then for any $\alpha\in\Phi$, $\mathbb{P}(\ker\alpha)$ intersects $C_{\zeta}$ transversally. Therefore:
      $Q_{1}(\zeta)=\text{ sum of intersection points of $C_{\zeta}$ with $\cup_{\alpha\in\Phi}\ker\alpha$}\geq 2s_{\zeta_{0}}\geq\sum_{k}^{s_{\zeta_{0}}}\sharp\,\{u_{k}^{1},u_{k}^{2}\}=Q_{1}(\zeta_{0}).$
  • (3)
    $\text{rank}(\zeta_{0})=3$. Since $\{\zeta\mid\text{rank}(\zeta)>2\}$ is open, $\exists\,\delta_{3}>0$ such that $\forall\zeta$ with $|\zeta-\zeta_{0}|<\delta_{3}$, $\text{rank}(\zeta)=3$. As $\zeta\to\zeta_{0}$, $C_{\zeta}\to C_{\zeta_{0}}$ in $\mathbb{P}(\mathfrak{h}_{\mathbb{C}})$. The transitive points $\gamma_{k}$ of $C_{\zeta_{0}}$ (defined in subsection 4.3) perturbs to transitive points again. Hence, $Q_{1}(\zeta)\geq Q_{1}(\zeta_{0})$.
6. Proof of Corollary 1.6
Hitchin proved that for any hyperKähler ALE with group $\mathbb{Z}_{2}$, its twistor space cannot be Kählerian due to the existence of holomorphic curves fiberwised in twistor space [6]. Hence we need only consider the case where $\Gamma\neq\mathbb{Z}_{2}$. If $\mathrm{Z}(X_{\zeta})$ is Kählerian, then its canonical bundle $K_{\mathrm{Z}}$ would be negative established by Hitchin [6]. Hence, the line bundle $K_{\mathrm{Z}}$ would be of negative degree when restricted on any compact holomorphic curves in $\mathrm{Z}(X_{\zeta})$. The following lemma that $K_{\mathrm{Z}}\simeq\pi^{*}\mathcal{O}_{\mathbb{P}^{1}}(-4)$ and the property of pullback, imply the line bundle $K_{\mathrm{Z}}$ is trivial when restricted to any fiber of $\pi$. Theorem 1.2 implies that there is always at least one fiber containing a rational curve. The existence of such holomorphic curves leads to the contradiction. Therefore, $\mathrm{Z}(X_{\zeta})$ cannot be Kählerian.∎
Lemma 6.1.
Denote $\mathrm{Z}(M)$ as the twistor space of a hyperKähler manifold $\left(M,g,I,J,K\right)$. Recall the natural holomorphic projection $\pi:\mathrm{Z}(M)\to\mathbb{P}^{1}$. Then, the canonical bundle $K_{\mathrm{Z}(M)}$ of the complex threefold $\mathrm{Z}(M)$ is given by:
$K_{\mathrm{Z}(M)}\simeq\pi^{*}\mathcal{O}_{\mathbb{P}^{1}}(-4).$
Proof.
Denote the three Kähler forms of $\left(M,g,I,J,K\right)$ by $\left(\omega_{I},\omega_{J},\omega_{K}\right)$. The twisted vertical two-form, $\Omega\in H^{0}(\mathrm{Z}(X_{\zeta}),\,\Lambda^{2}T_{F}^{*}\otimes\pi^{*}\mathcal{O}_{\mathbb{P}^{1}}(2))$, as defined in [7], is given by
$\Omega(u)$ $=(\omega_{I}+\mathrm{i}\omega_{J})+u(2\omega_{I})-u^{2}(\omega_{J}-\mathrm{i}\omega_{K}),$
$\Omega(v)$ $=v^{2}(\omega_{I}+\mathrm{i}\omega_{J})+v(2\omega_{I})-(\omega_{J}-\mathrm{i}\omega_{K}),$
[…]
Example 6.2.
For $\Gamma=\mathbb{Z}_{2}$, by Corollary 3.1, the hyperKähler ALE structure is realized on the minimal resolution $\widetilde{\mathbb{C}^{2}/\mathbb{Z}_{2}}$. As complex manifolds, there is an isomorphism
$\widetilde{\mathbb{C}^{2}/\mathbb{Z}_{2}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)\simeq\mathcal{T}^{*}\,\mathbb{P}^{1}.$
The hyperKähler metric on $\mathcal{O}_{\mathbb{P}^{1}}(-2)$ was constructed by Eguchi-Hanson [3] and also by Calabi [Calabi:79/mj]. Its twistor space $\mathrm{Z}$, as explained in [5], is obtained by resolving the singularities on a singular hypersurface $\tilde{\mathrm{Z}}$ of a four-dimensional complex manifold:
$\tilde{\mathrm{Z}}:=\left\{xy=(z-p_{1})(z-p_{2})\right\}\subset\mathcal{O}_{\mathbb{P}^{1}}(2)\oplus\mathcal{O}_{\mathbb{P}^{1}}(2)\oplus\mathcal{O}_{\mathbb{P}^{1}}(2)$
where $x,y,z$ are parameters of the fibers of three line bundle, and $p_{1},p_{2}$ are some two distinct sections of $\mathcal{O}_{\mathbb{P}^{1}}(2)$.
$\mathbf{Acknowledgements.}$ The authors wished to thank Friedrich Knop for answering questions on MathOverflow and providing the useful reference [10]. We also thank Song Sun for appreciated conversations.