### v3

# Counting rational curves and standard complex structures on hyperKähler ALE 4-manifolds

Abstract.
All hyperKähler ALE 4-manifolds with a given non-trivial finite group $1\not=\Gamma<{\rm SU}(2)$ at infinity are parameterized by an open dense subset of a real linear space of dimension $3{\rm rank}\,\Phi$. Here, $\Phi$ denotes the root system associated with $\Gamma$ via the McKay correspondence. Such manifolds are diffeomorphic to the minimal resolution of a Kleinian singularity. By using the period map of the twistor space, we specify those points in the parameter space at which the hyperKählerian family of complex structures includes the complex structure of the minimal resolution. Furthermore, we count the rational curves lying on each hyperKähler ALE 4-manifold. As an application, we prove that the twistor space of any hyperKähler ALE cannot be Kählerian. In particular, we strengthen some results of Kronheimer (J. Differential Geom., 29(3):665–683, 1989) and provide examples of non-compact and non-Kählerian twistor spaces.
Keywords hyperKähler ALE, Kleinian singularity, twistor space, root system, rational curve.
2000 MR Subject Classification 53C28, 17B22
B.X. is supported in part by the Project of Stable Support for Youth Team in Basic Research Field, CAS (Grant No. YSBR-001) and NSFC (Grant Nos. 12271495 and 12071449).
Y.L. is the corresponding author.
1. Introduction
1.1. Background
A hyperKähler structure on a Riemannian manifold $\left(M,g\right)$ is, by definition, a triple $(I,J,K)$ of three parallel complex structures on $M$ that satisfies the quaternion relation $IJ=-JI=K$. The set $\left\{aI+bJ+cK\,|\,(a,b,c)\in\mathbb{S}^{2}\right\}$ is called the hyperKählerian family of complex structures on $(M,g)$. We say a 4-dimensional Riemannian manifold $(M,g)$ is asymptotically locally Euclidean (ALE) if there exist a compact set $K\subset M$, a finite subgroup $1\not=\Gamma<\mathrm{SO}(4)$, and a diffeomorphism $\Psi:M\backslash K\to(\mathbb{R}^{4}\backslash B_{R})/\Gamma$, where $B_{R}$ is the standard ball of radius $R$ in $\mathbb{R}^{4}$, such that for each non-negative integer $k$, there holds
$|\nabla^{k}(\Psi_{*}g-g_{\rm Euc})(x)|=O(|x|^{-4-k})\,,\qquad x\in(\mathbb{R}^{4}\backslash B_{R})/\Gamma\qquad{\rm and}\qquad|x|\to+\infty.$
Kronheimer [10, 11] classified all hyperKähler ALE 4-manifolds. He constructed explicitly for each finite group $1\not=\Gamma<\mathrm{SU}(2)$ a complete family of hyperKähler ALEs with group $\Gamma$ at infinity. We provide a brief overview as follows. Let’s fix a non-trivial finite group $\Gamma<\mathrm{SU}(2)$.
Denote $Q$ as the natural two-dimensional complex representation of $\Gamma$ and $R$ as the regular representation. Define a linear space
$M(\Gamma):=\mathrm{Hom}_{\Gamma}\bigl{(}R,Q\otimes R\bigr{)}$
equipped with a natural hyperKähler structure when we identify $Q=\mathbb{C}^{2}\simeq\mathbb{H}$. This space $M(\Gamma)$ is acted upon by a Lie group $\mathrm{U}(\Gamma)$, which preserves the hyperKähler structure on $M(\Gamma)$. Denote by $Z$ the dual of the center of the Lie algebra $\mathfrak{U}(\Gamma)$ of $\mathrm{U}(\Gamma)$. There exists a natural isomorphism
$\rho:Z\xrightarrow{\sim}\mathfrak{h},$
where $\mathfrak{h}$ is the Cartan subalgebra associated with $\Gamma$, see subsection 2.1 for $\mathfrak{h}$. With respect to the $\mathrm{U}(\Gamma)$-action, we can associate a hyperKähler moment map
$\mu_{\mathbb{H}}:M(\Gamma)\to\mathfrak{U}(\Gamma)^{*}\otimes\mathbb{R}^{3}$
and consider the hyperKähler quotient $X_{\zeta}:=\mu_{\mathbb{H}}^{-1}(\zeta)/\mathrm{U}(\Gamma)$ for each $\zeta\in Z\otimes\mathbb{R}^{3}$. For details on hyperKähler quotient, see [8]. These quotient manifolds may not be smooth. However, for $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$, $X_{\zeta}$ is a smooth ALE 4-manifold equipped with the quotient metric $g_{\zeta}$, where
$(Z\otimes\mathbb{R}^{3})^{\circ}:=(Z\otimes\mathbb{R}^{3})\backslash\bigcup_{\theta\in\Phi}(D_{\theta}\otimes\mathbb{R}^{3})\,,\qquad D_{\theta}:=\mathrm{Ker}(\theta\circ\rho)\subset Z$
and the union is taken over all roots in the root system $\Phi$ associated with $\mathfrak{h}$. By the property of hyperKähler quotient, the smooth quotient Riemannian manifold $\left(X_{\zeta},g_{\zeta}\right)$ admits a natural hyperKähler structure. We denote by $I_{\zeta},J_{\zeta},K_{\zeta}$ the three compatible complex structures on $X_{\zeta}$, and by $\omega_{\zeta}^{k}$, $k\in\{1,2,3\}$ the corresponding Kähler forms. The ordering of $I_{\zeta},J_{\zeta},K_{\zeta}$ is inherited from the ordering of complex structures on the hyperKähler manifold $M(\Gamma)$.
Furthermore, Kronheimer showed that each smooth manifold $X_{\zeta}$ for $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$ is diffeomorphic to the minimal resolution of $\mathbb{C}^{2}/\Gamma$. We denote the minimal resolution of the $\mathbb{C}^{2}/\Gamma$ by $\eta:\widetilde{\mathbb{C}^{2}/\Gamma}\to\mathbb{C}^{2}/\Gamma$ (see Definition 2.1). The set $\eta^{-1}(0)$ is called the exceptional locus, it is a union of $\mathrm{rank}\,\Phi$ rational curves. We call the complex structure on the minimal resolution $\widetilde{\mathbb{C}^{2}/\Gamma}$ the standard complex structure, which we denote by $I_{\mathrm{std}}$.
1.2. Main results
For $\zeta=(\zeta^{1},0,0)\in(Z\otimes\mathbb{R}^{3})^{\circ}$, Kronheimer [10, Proposition 3.10] showed that the complex manifold $X_{\zeta}$, with respect to the complex structures $I_{\zeta}$, is isomorphic to $\widetilde{\mathbb{C}^{2}/\Gamma}$. Our first main result characterizes all such $\zeta$ as follows. Denote by $\mathrm{rank}(\zeta)$ the dimension of $\mathrm{span}_{\mathbb{R}}\left\{\zeta^{1},\zeta^{2},\zeta^{3}\right\}$. It is clear that $1\leq\mathrm{rank}(\zeta)\leq 3$ for each $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$.
Theorem 1.1.
Let $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$. The hyperKählerian family of complex structures on $(X_{\zeta},g_{\zeta})$ includes $I_{\rm std}$ if and only if ${\rm rank}(\zeta)=1$, that is $\zeta$ is of the form $\left(x\zeta^{1},y\zeta^{1},z\zeta^{1}\right)$ for some non-zero vector $(x,y,z)\in\mathbb{R}^{3}\setminus\{0\}$ and $\zeta^{1}\in Z\setminus\bigcup_{\theta\in\Phi}D_{\theta}$. Under this context, the hyperKählerian family of $X_{\zeta}$ includes exactly two complex structures which are equivalent to $I_{\rm std}$, and all the other complex structures are mutually equivalent and isomorphic to a hypersurface in $\mathbb{C}^{3}$.
The twistor space [8, section 3.F] of the hyperKähler manifold $\left(X_{\zeta},g_{\zeta}\right)$ is a complex threefold, denoted by $\mathrm{Z}(X_{\zeta})$, equipped with a holomorphic projection $\pi=\pi_{\zeta}:\mathrm{Z}(X_{\zeta})\to\mathbb{P}^{1}$. For $u\in\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$ we define a complex structure $I_{\zeta}^{u}:=aI_{\zeta}+bJ_{\zeta}+cK_{\zeta}$, where
[…]
Note that equation (1.1) provides an identification between $\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$ and
$\mathbb{S}^{2}=\left\{(a,b,c)\in\mathbb{R}^{3}|a^{2}+b^{2}+c^{2}=1\right\}.$
The fiber $\pi^{-1}(u)$ over $u$ is a complex submanifold of $\mathrm{Z}(X_{\zeta})$ that is isomorphic to the complex manifold $X_{\zeta}$ with respect to the complex structure $I_{\zeta}^{u}$. Our next result is:
Theorem 1.2.
Fix a non-trivial finite subgroup $\Gamma$ in $\mathrm{SU(2)}$. Recall the linear space $Z$ and root system $\Phi$ associated with $\Gamma$. Given $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$, we could count rational curves fiberwise lying in the twistor space $\pi:\mathrm{Z}(X_{\zeta})\to\mathbb{P}^{1}$ as follows:
  • (i)
    If $\mathrm{rank(\zeta)}=1$, then there hold
    (1.2) #$\#\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}=2,$
    and
    (1.3) $\sum_{u\in\mathbb{P}^{1}}\#\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}(u)\right\}=2r,$
    where $r:=\mathrm{rank}\,\Phi$.
  • (ii)
    If $\mathrm{rank(\zeta)}=2$, then there hold
    (1.4) $4\leq\#\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}\leq\#\Phi,$
    and
    (1.5) $2r\leq\sum_{u\in\mathbb{P}^{1}}\#\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}(u)\right\}\leq\#\Phi.$
  • (iii)
    If $\mathrm{rank(\zeta)}=3$, then there hold
    (1.6) $3\leq\#\left\{u\in\mathbb{P}^{1}\,|\,\pi^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}\leq\#\Phi,$
    and
    (1.7) $2r-1\leq\sum_{u\in\mathbb{P}^{1}}\#\left\{\mathrm{rational\,\,curves\,\,in\,\,}\pi^{-1}(u)\right\}\leq\#\Phi.$
Remark 1.3.
The upper bounds for (1.4) and (1.6) were established by Kronheimer [11, Lemma 2.7]. All lower bounds presented here are new results. However, determining optimal lower bounds remains challenging, primarily due to the underlying combinatorial problem discussed in Section 4. We plan to address this optimization problem in subsequent work.
Considering all $\zeta\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$, we define a map as:
$Q_{1}:\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$ $\to\mathbb{Z}$
$\zeta$$\mapsto\#\left\{u\in\mathbb{P}^{1}\,|\,\pi_{\zeta}^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}$
Although an explicit expression for $Q_{1}$ appears difficult to obtain, we establish:
Theorem 1.4.
$Q_{1}$ is lower semi-continuous, that is, for every $\zeta_{0}\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$, there holds
$\liminf_{\zeta\to\zeta_{0}}Q_{1}(\zeta)\geq Q_{1}(\zeta_{0})$
Remark 1.5.
Verifying the semi-continuity property for rational curve counts remains challenging.
Combining Theorem 1.1 and Theorem 1.2, we extend certain results of Hitchin [7] concerning the non-compact non-Kählerian twistor spaces:
Corollary 1.6.
For any finite group $1\not=\Gamma<\mathrm{SU(2)}$ and any $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$, the twistor space $\mathrm{Z}(X_{\zeta})$ can not be Kählerian (that is non-Kählerian).
1.3. Outline
Section 2 reviews fundamental aspects of Kleinian singularities and their semi-universal deformations. The core technical tool, developed in Subsections 2.3 and 2.4, is the period map associated with the twistor space of hyperKähler ALE manifolds. Section 3 provides proofs for Theorems 1.1 and 1.2 in Case (i). Section 4 completes the proof of Theorem 1.2 by reformulating it as a decomposition problem for irreducible root systems. This formulation naturally leads to new questions regarding root systems in Lie theory. Section 5 establishes the lower semi-continuity of $Q_{1}$ and presents the proof of Theorem 1.4. The final section derives Corollary 1.6 from Hitchin’s foundational work [7].
1.4. Some basic notations
  • $\left(X_{\zeta},I_{\zeta}^{u}\right)$: The complex manifold with underlying differential manifold $X_{\zeta}$ and complex structure $I_{\zeta}^{u}$.
  • $\mathfrak{h}^{\mathbb{C}}$: The complexification of the real Cartan subalgebra $\mathfrak{h}$.
  • $\mathbb{P}(V)$ or $\mathbb{P}V$: The projective space associated with the complex vector space $V$.
  • #$\#P$: The cardinality of the set $P$.
  • $P_{1}\sqcup P_{2}$: The disjoint union of two sets $P_{1}$ and $P_{2}$.
  • $\mathrm{i}$: The imaginary unit, $\sqrt{-1}$.
2. The Kleinian singularities
Building on Kronheimer’s foundational work [10, 11] establishing key connections between hyperKähler ALE spaces and Kleinian singularities, this section reviews essential background and prepares for proving our main theorems. Specifically: Let $\Gamma$ be a non-trivial finite subgroup of $\mathrm{SU}(2)$. The singular complex surface $\mathbb{C}^{2}/\Gamma$ is termed a Kleinian singularity.
  • Subsection 2.1 revisits the ADE root system correspondence for Kleinian singularities.
  • Subsection 2.2 summarizes structural properties of semi-universal deformations and their simultaneous resolutions.
  • Subsection 2.3 details the hyperKähler ALE-Kleinian singularity relationship: The twistor space of a hyperKähler ALE manifold serves as a simultaneous resolution for deformations of the associated singularity (Proposition 2.12). We further examine the period map of this twistor space—a fundamental tool for our proofs (Lemma 2.13 and Corollary 2.14).
  • Subsection 2.4 establishes a crucial isomorphism in the parameter space $Z\otimes\mathbb{R}^{3}$ (Lemma 2.16), leveraging Kronheimer’s results to enable our proof of Theorem 1.1 in Section 3.
2.1. Kleinian singularities and root systmes
Definition 2.1.
[1, Chapter III] A minimal resolution of a Kleinian singularity $\mathbb{C}^{2}/\Gamma$ is a smooth complex surface $\widetilde{\mathbb{C}^{2}/\Gamma}$ together with a proper holomorphic map $\eta:\widetilde{\mathbb{C}^{2}/\Gamma}\to\mathbb{C}^{2}/\Gamma$ such that:
  • The restriction, $\eta:\widetilde{\mathbb{C}^{2}/\Gamma}\setminus\eta^{-1}(0)\to\mathbb{C}^{2}/\Gamma\setminus{0}$, is an isomorphism,
  • $\widetilde{\mathbb{C}^{2}/\Gamma}$ contains no $(-1)$-curves.
The set $\eta^{-1}(0)$ is called the exceptional locus of $\widetilde{\mathbb{C}^{2}/\Gamma}$.
Any Kleinian singularity admits a unique minimal resolution, denoted by $\eta:\widetilde{\mathbb{C}^{2}/\Gamma}\to\mathbb{C}^{2}/\Gamma$. The exceptional locus is a union of rational curves $\eta^{-1}(0)=\Sigma_{1}\cup\cdots\cup\Sigma_{r}$, where each $\Sigma_{i}\simeq\mathbb{P}^{1}$. Topologically, these 2-spheres $\Sigma_{i}$ ($i=1,\dots,r$) generate the second homology group $H_{2}(\widetilde{\mathbb{C}^{2}/\Gamma};\mathbb{Z})$.
We associate a vertex with each $\Sigma_{i}$ and connect two vertices with an edge if the corresponding curves intersect. This construction yields a diagram $\Delta(\Gamma)$ of $\eta^{-1}(0)$. It is well known that $\Delta(\Gamma)$ is an ADE-type Dynkin diagram. Associating $\Gamma$ with its diagram $\Delta(\Gamma)$ induces a bijection [12, p.111]:
[…]
$\Gamma\,\,$ $\mapsto\,\,\Delta(\Gamma).$
Thus, the Kleinian singularity $\mathbb{C}^{2}/\Gamma$ is of type A, D, or E according to whether $\Delta(\Gamma)$ is of type A, D, or E.
Denote by $\mathfrak{h}$ the real Cartan algebra associated with $\Delta(\Gamma)$, and by $\Phi$ the corresponding root system. The vertices of $\Delta(\Gamma)$ correspond to simple roots $\theta_{1},\dots,\theta_{r}\in\mathfrak{h}^{*}$. Let $\Delta$ denote the set of simple roots, which forms a base of $\Phi$. Under the minimal resolution $\widetilde{\mathbb{C}^{2}/\Gamma}$, the 2-spheres $\Sigma_{i}$ ($i=1,\dots,r$) correspond bijectively to the simple roots $\theta_{i}$. Since these 2-spheres generate $H_{2}(\widetilde{\mathbb{C}^{2}/\Gamma};\mathbb{Z})$, we obtain an embedding of the root system $\Phi\hookrightarrow H_{2}(\widetilde{\mathbb{C}^{2}/\Gamma};\mathbb{Z})$. Hence there are natural isomorphisms:
$H_{2}\big{(}\widetilde{\mathbb{C}^{2}/\Gamma};\mathbb{R}\big{)}\simeq\mathfrak{h}^{*},\quad H^{2}\big{(}\widetilde{\mathbb{C}^{2}/\Gamma};\mathbb{C}\big{)}\simeq\mathfrak{h}^{\mathbb{C}}.$
The Kleinian singularity $\mathbb{C}^{2}/\Gamma$ is said to be of type $\Phi$ at the singular point 00. This classification is intrinsically local.
2.2. Deformations of Kleinian singularities
Definition 2.2.
[5, p.7] A complex space is a pair $(X,\mathcal{O}_{X})$, where $X$ is a topological space $X$ and $\mathcal{O}_{X}$ is a structure sheaf, which satisfies the following properties:
  • (i)
    Every point $x\in X$ admits a neighborhood $U$ such that $U$ can be represented as $U=D\cap Z$, where $D$ is a domain in $\mathbb{C}^{n}$ for some positive integer $n$ and $Z$ is the zero set of some holomorphic functions $f_{1},\cdots,f_{k}$ on $D$.
  • (ii)
    And the restriction of the structure sheaf $\mathcal{O}_{X}$ to $U$, denoted by $\mathcal{O}_{X}(U)$, is isomorphic to the restriction of the quotient sheaf $\mathcal{O}_{D}/I$ to $U$, where $I$ is the ideal generated by $f_{1},\cdots,f_{k}$.
Morphisms between complex spaces are called holomorphic maps.
Definition 2.3.
[14, Introduction] A deformation of a complex space $X^{0}$ is a flat holomorphic map $\pi:\mathcal{X}\to T$ between two complex spaces $\mathcal{X}$ and $T$, together with an isomorphism $\xi:X^{0}\xrightarrow{\sim}\pi^{-1}(t_{0})$, where $t_{0}$ is a point in $T$. We call $t_{0}$ the distinguished point. This data is expressed by the following commutative diagram:
[…]
Here $\pi$ is a flat holomorphic map, and $\xi$ is an embedding. We say a deformation is a $\mathbb{C}^{*}$-deformation if the above diagram admits a $\mathbb{C}^{*}$-equivariant action. This means there exist respectively $\mathbb{C}^{*}$-actions on complex spaces $X^{0},\mathcal{X}$ and $T$ such that $\left\{t_{0}\right\}$ is a fixed point in $T$ and the morphisms $\xi$ and $\pi$ are equivariant.
Lemma 2.4.
[13, Section 8.7][10, p.681] There exists a $\mathbb{C}^{*}$-deformation of the Kleinian singularity $\mathbb{C}^{2}/\Gamma$, denoted by $f:\mathcal{Y}\to\mathcal{V}$, satisfying:
  • (1)
    $\mathcal{V}$ is isomorphic to $\mathfrak{h}^{\mathbb{C}}/W$, where $\mathfrak{h}^{\mathbb{C}}$ is the complexification of $\mathfrak{h}$, and the distinguished point $0\in\mathfrak{h}^{\mathbb{C}}/W$ satisfies $f^{-1}(0)\simeq\mathbb{C}^{2}/\Gamma$.
  • (2)
    The branch locus $\mathcal{D}\subset\mathcal{V}$ consists of all points $v$ where the fiber $f^{-1}(v)$ is singular.
  • (3)
    $\mathcal{D}$ equals the image under the quotient map $\mathfrak{h}^{\mathbb{C}}\to\mathfrak{h}^{\mathbb{C}}/W$ of $\bigcup_{\alpha\in\Phi}\ker\alpha$.
This deformation $f:\mathcal{Y}\to\mathcal{V}$ is the semi-universal deformation of $\mathbb{C}^{2}/\Gamma$.
Lemma 2.5.
According to [9, pp. 23-25], each fiber of map $f$ in Lemma 2.4 is a complex surface (possibly singular) in the affine space $\mathbb{C}^{3}$. Moreover, in [15, Section 1], Tjurina pointed out, with the ADE classification of Kleinian singularities stated in our subsection 2.1, the space $\mathcal{Y}$ in the semi-universal deformation $f:\mathcal{Y}\to\mathcal{V}$ is a hypersurface in a vector space $\mathbb{C}^{3}\crossproduct\mathbb{C}^{r}$ given by one of the following equations:
$\mathrm{Type}\,A_{r}:\,\,$ $xy+z^{r+1}+t_{1}z^{r-1}+t_{2}z^{r-2}+\cdots+t_{r-1}z+t_{r}=0;$
$\mathrm{Type}\,D_{r}:\,\,$ $x^{2}+y^{2}z+z^{r-1}+t_{1}z^{r-2}+\cdots+t_{r-1}+2t_{r}y=0;$
$\mathrm{Type}\,E_{6}:\,\,$ $x^{2}+y^{3}+z^{4}+t_{1}z^{2}+t_{2}z+t_{3}z+y(t_{4}z^{2}+t_{5}z+t_{6})=0;$
$\mathrm{Type}\,E_{7}:\,\,$ $x^{2}+y^{3}+yz^{3}+y(t_{1}z+t_{2})+t_{3}z^{4}+\cdots+t_{6}z+t_{7}=0;$
$\mathrm{Type}\,E_{8}:\,\,$ $x^{2}+y^{3}+z^{5}+y(t_{1}z^{3}+\cdots+t_{4})+t_{5}z^{3}+\cdots+t_{8}=0,$
and the projection $f$ is simply the projection to the parameter space $\{(t_{1},\cdots,t_{r})\in\mathbb{C}^{r}\}\simeq\mathcal{V}$. This explicit expression will not be used later in the text; we would only use the fact that each fiber of $f$ is in $\mathbb{C}^{3}$.
We denote by $\tilde{f}:\tilde{\mathcal{Y}}\to\mathfrak{h}^{\mathbb{C}}$ the deformation which is the pullback of the semi-universal deformation $f$ by the natural quotient map $q:\mathfrak{h}^{\mathbb{C}}\to\mathfrak{h}^{\mathbb{C}}/W\simeq\mathcal{V}$, that is,
(2.1) ${\widetilde{\mathcal{Y}}}$${\mathcal{Y}}$${\mathfrak{h}^{\mathbb{C}}}$${\mathcal{V}}$$\scriptstyle{\widetilde{f}}$$\scriptstyle{f}$$\scriptstyle{q}$
Then, we have the explicit description about the singularities in the singular fibers of $\tilde{f}$:
Lemma 2.6.
[13, Lemma 1, p.94] For any $\lambda\in\mathfrak{h}^{\mathbb{C}}$, denote $\Phi_{\lambda}:=\{\theta\in\Phi\,|\,\theta(\lambda)=0\}$, which is a root subsystem of the root system $\Phi$. We have:
  • (1)
    The fiber $\tilde{\mathcal{Y}}_{\lambda}$ has singularities if and only if $\lambda\in\bigcup_{\theta\in\Phi}\ker\ \theta$.
  • (2)
    Let $\Phi_{\lambda}=\Phi_{\lambda,1}\sqcup\cdots\sqcup\Phi_{\lambda,m}$ be the orthogonal decomposition of $\Phi_{\lambda}$ into irreducible root subsystems. Then the fiber $\tilde{\mathcal{Y}}_{\lambda}$ has exactly $m$ Kleinian singularities of type $\Phi_{\lambda,1},\cdots,\Phi_{\lambda,m}$ at $m$ points respectively.
As a consequence, any singular fiber $\tilde{\mathcal{Y}}_{\lambda}$ is a singular complex surface that has only finite isolated singularities.
Definition 2.7.
[1, Section 6. Chapter III] A minimal resolution of a singular complex surface $X$ with only isolated singularities at a finite set $S$ is a smooth complex surface $\widetilde{X}$ together with a proper holomorphic map $\eta:\widetilde{X}\to X$ such that:
  • The restriction to the smooth part, $\eta:\widetilde{X}\setminus\eta^{-1}(S)\to X\setminus S$, is an isomorphism as complex manifolds.
  • $\tilde{X}$ contains no $(-1)$-curves.
And we call the pre-image of singularities $\eta^{-1}(S)$ the exceptional locus of $\widetilde{X}$.
Lemma 2.8.
[1, Theorem 6.2, Chapter III] Every normal surface with finitely many singularities admits a unique minimal resolution. In particular, for each $\lambda\in{\mathfrak{h}}^{\mathbb{C}}$, the fiber $\mathcal{Y}_{q\left(\lambda\right)}\simeq\tilde{\mathcal{Y}}_{\lambda}$ is normal since all the singularities are Kleinian by Lemma 2.6, and then admits a unique minimal resolution.
Lemma 2.9.
[10, pp.680-681] [13, Section 8.7] There exists a commutative diagram of holomorphic maps:
${\mathfrak{Y}}$${\mathcal{Y}}$${\mathfrak{h}^{\mathbb{C}}}$${\mathcal{V}}$$\scriptstyle{\psi}$$\scriptstyle{\hat{f}}$$\scriptstyle{f}$$\scriptstyle{q}$
such that:
  • $\hat{f}$ is a holomorphic map between smooth complex manifolds.
  • The restriction of $\psi$ to each fiber $\mathfrak{Y}_{\lambda}$ of $\hat{f}$ is the minimal resolution of the fiber $\mathcal{Y}_{q(\lambda)}$ of $f$, where $\lambda\in\mathfrak{h}^{\mathbb{C}}$.
We call the map $\hat{f}$ is a simultaneous resolution of the semi-universal deformation $f$.
Lemma 2.10.
Any rational curve in the fiber $\mathfrak{Y}_{\lambda}$ of $\hat{f}$, $\lambda\in\mathfrak{h}^{\mathbb{C}}$, must lie in the exception locus of $\mathfrak{Y}_{\lambda}$.
Proof.
By Lemma 2.5, $\mathcal{Y}_{q\left(\lambda\right)}$ is a hypersurface in $\mathbb{C}^{3}$. Since any holomorphic map $\mathbb{P}^{1}\hookrightarrow\mathcal{Y}_{q\left(\lambda\right)}\subset\mathbb{C}^{3}$ must be constant, $\mathcal{Y}_{q\left(\lambda\right)}$ contains no rational curves. Suppose that a rational curve in $\mathfrak{Y}_{\lambda}$ is given by a holomorphic map $\mathbb{P}^{1}\hookrightarrow\mathfrak{Y}_{\lambda}$. Then its composition with the resolution $\mathfrak{Y}_{\lambda}\to\mathcal{Y}_{q\left(\lambda\right)}\subset\mathbb{C}^{3}$ is a constant map. Hence by the definition of minimal resolution, the rational curve must lie in the exceptional locus of $\mathfrak{Y}_{\lambda}$. ∎
Corollary 2.11.
For any $\lambda\in\mathfrak{h}^{\mathbb{C}}$, recall that $\Phi_{\lambda}=\Phi_{1,\lambda}\sqcup\cdots\sqcup\Phi_{m,\lambda}$ represents the type of singularities in Lemma 2.6. Then we have the counting of rational curves in $\mathfrak{Y}_{\lambda}$:
#$\#\left\{\mathrm{rational\,\,curves\,\,in\,\,}\mathfrak{Y}_{\lambda}\right\}=\mathrm{rank}\,\Phi_{\lambda}.$
Proof.
By Lemma 2.10, the number of rational curves in $\mathfrak{Y}_{\lambda}$ is the sum of the numbers of rational curves in the exceptional locus of each singularity, hence is given by the sum of the ranks of the types $\Phi_{1,\lambda},\cdots,\Phi_{m,\lambda}$ of these singularities in $\mathcal{Y}_{q\left(\lambda\right)}\simeq\tilde{\mathcal{Y}}_{\lambda}$. Since the decomposition is orthogonal, we have:
$\mathrm{rank}\ \Phi_{\lambda}=\mathrm{rank}\ \Phi_{1,\lambda}+\cdots+\mathrm{rank}\ \Phi_{m,\lambda}.$
2.3. Period map of twistor space
Recall the notations from Subsections 1.1 and 1.2. For any non-trivial finite subgroup $\Gamma<\mathrm{SU}(2)$, consider $\zeta=(\zeta^{1},\zeta^{2},\zeta^{3})\in(Z\otimes\mathbb{R}^{3})^{\circ}$ as defined in Section 1.1. The hyperKähler manifold $X_{\zeta}$ carries the following:
  • Three compatible complex structures $I_{\zeta}$, $J_{\zeta}$, $K_{\zeta}$;
  • Associated Kähler forms $\omega_{\zeta}^{k}$ ($k=1,2,3$);
  • A twistor space denoted by $\mathrm{Z}(X_{\zeta})$ with a canonical holomorphic projection $\pi:\mathrm{Z}(X_{\zeta})\to\mathbb{P}^{1}$.
Proposition 2.12.
[11, pp.691-692] Related to $\mathrm{Z}(X_{\zeta})$, there is a $\mathbb{C}^{*}$-deformation of $\mathbb{C}^{2}/\Gamma$:
$\phi:Y\to\mathbb{C}^{2}.$
where the $\mathbb{C}^{*}$-action on $\mathbb{C}^{2}$ is the natural diagonal action and $\phi^{-1}(0)\simeq\mathbb{C}^{2}/\Gamma$. We define a map $\pi^{s}:{\rm Z}^{s}\to\mathbb{P}^{1}$ by:
[…]
Then there exists a commutative diagram below:
(2.2) ${\mathrm{Z}(X_{\zeta})}$${\mathrm{Z}^{s}}$${\mathbb{P}^{1}}$${\mathbb{P}^{1}}$$\scriptstyle{\chi}$$\scriptstyle{\pi}$$\scriptstyle{\pi^{s}}$
such that the restriction of $\chi$ to each fiber of $\pi$ provides the minimal resolution of corresponding fiber of $\pi^{s}$. We call the projection $\pi:\mathrm{Z}(X_{\zeta})\to\mathbb{P}^{1}$ a simultaneous resolution of the map $\pi^{s}$.
As the twistor space of a hyperKähler manifold, $\mathrm{Z}(X_{\zeta})$ carries a twisted vertical 2-form $\Omega\in H^{0}\big{(}\mathrm{Z}(X_{\zeta}),\,\Lambda^{2}T_{F}^{*}\otimes\pi^{*}\mathcal{O}_{\mathbb{P}^{1}}(2)\big{)}$ by [8]. Defined on the affine chart $\mathbb{C}\subset\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$, it is given by:
$\Omega(u)=\omega^{2}_{\zeta}+\mathrm{i}\omega^{3}_{\zeta}+2u\,\omega^{1}_{\zeta}-u^{2}\left(\omega^{2}_{\zeta}-\mathrm{i}\omega^{3}_{\zeta}\right),$
[…]
$\operatorname{Re}\Omega(u)$ $=2u_{1}\omega_{\zeta}^{1}+(1-u_{1}^{2}+u_{2}^{2})\,\omega_{\zeta}^{2}-2u_{1}u_{2}\,\omega_{\zeta}^{3},$
$\operatorname{Im}\Omega(u)$ $=2u_{2}\,\omega_{\zeta}^{1}-2u_{1}u_{2}\,\omega_{\zeta}^{2}+(1+u_{1}^{2}-u_{2}^{2})\,\omega_{\zeta}^{3}.$
Note that $\Omega(u)$ is a holomorphic two-form on $(X_{\zeta},I_{\zeta}^{u})$, where $I_{\zeta}^{u}$ is defined in the paragraph containing equation (1.1). The cohomology class $[\Omega(u)]$ is called the period of $(X_{\zeta},I_{\zeta}^{u})$. Taking the cohomology class of $\Omega$ yields, under the identification $H^{2}(X_{\zeta},\mathbb{C})\simeq\mathfrak{h}^{\mathbb{C}}$, an element of $H^{0}\big{(}\mathbb{P}^{1},\mathfrak{h}^{\mathbb{C}}\otimes\mathcal{O}_{\mathbb{P}^{1}}(2)\big{)}$. Lifting to $\mathbb{C}^{2}$, we obtain the map:
(2.3) $\widetilde{p_{\zeta}}\colon\mathbb{C}^{2}$ $\to\mathfrak{h}^{\mathbb{C}},$
$(z_{1},z_{2})$ $\mapsto z_{1}^{2}[\omega^{2}_{\zeta}+\mathrm{i}\omega^{3}_{\zeta}]+2z_{1}z_{2}[\omega^{1}_{\zeta}]-z_{2}^{2}[\omega^{2}_{\zeta}-\mathrm{i}\omega^{3}_{\zeta}].$
[…]
Lemma 2.13.
[11, p. 693] Recall the semi-universal deformation $f:\mathcal{Y}\to\mathcal{V}$ of $\mathbb{C}^{2}/\Gamma$ in Lemma 2.4. Then the deformation $\phi$ is the pullback of the semi-universal deformation by the period map $p_{\zeta}$, that is, the following diagram holds:
${Y}$${\mathcal{Y}}$${\mathbb{C}^{2}}$${\mathcal{V}}$$\scriptstyle{\phi}$$\scriptstyle{f}$$\scriptstyle{p_{\zeta}}$
The singular fibers of (the pullback of) semi-universal deformation $f$ and the types of their singularities are characterized by Lemma 2.6. Combining these with Lemma 2.13 and Proposition 2.12, we have:
Corollary 2.14.
Recall the holomorphic map $\pi:\mathrm{Z}(X_{\zeta})\to\mathbb{P}^{1}$. Given $u\in\mathbb{P}^{1}$, then:
  • (1)
    The complex submanifold $\pi^{-1}(u)$ is isomorphic to the minimal resolution of the fiber $\phi^{-1}(\tilde{u})$ of $\phi$, where $\tilde{u}$ is any lifting of $u$ with respect to the natural quotient map $\mathbb{C}^{2}\setminus 0\to\mathbb{P}^{1}$.
  • (2)
    $\pi^{-1}(u)$ contains a rational curve if and only if the period $\widetilde{p_{\zeta}}(\tilde{u})\in\mathfrak{h}^{\mathbb{C}}$ lies in the kernel of some root in $\Phi$.
  • (3)
    The number of rational curves in $\pi^{-1}(u)$ is given by $\mathrm{rank}\,\Phi_{\lambda}$, where $\lambda:=\widetilde{p_{\zeta}}(\tilde{u})$.
Proof.
The statement (1) is immediately obtained from Proposition 2.12. By Lemma 2.13, the fiber $\phi^{-1}(\tilde{u})$ is isomorphic to the fiber $f^{-1}(p_{\zeta}(\tilde{u}))$ hence isomorphic to the fiber $\tilde{f}^{-1}(\widetilde{p_{\zeta}}(\tilde{u}))$. Thus, statement (2) is obtained from the Lemma 2.6. And statement (3) is a straightforward consequence of Corollary 2.11. ∎
2.4. An isomorphism between hyperKähler ALEs
Lemma 2.15.
[10, Proposition 4.1] Given $\zeta=\left(\zeta^{1},\zeta^{2},\zeta^{3}\right)\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$. There exists a linear isomorphism
$\sigma:Z\xrightarrow{\sim}H^{2}(X_{\zeta}^{\mathrm{Diff}},\mathbb{R})(\simeq\mathfrak{h}),$
such that $\sigma(\zeta^{k})=[\omega^{k}_{\zeta}]$, for $k=1,2,3.$ Consequently, the complexification $\sigma_{\mathbb{C}}:Z\otimes\mathbb{C}\to H^{2}(X_{\zeta}^{\mathrm{Diff}},\mathbb{C})$ is also an isomorphism.
Lemma 2.16.
Without loss of generality, we assume that $u\neq\infty$ and denote $u=u_{1}+\mathrm{i}u_{2}\in\mathbb{C}\hookrightarrow\mathbb{P}^{1}$. Recall the notations for complex structures: $I_{\zeta}^{u}=aI_{\zeta}+bI_{\zeta}+cI_{\zeta}$. For $\zeta=\left(\zeta^{1},\zeta^{2},\zeta^{3}\right)\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$, let
$\tilde{\zeta}^{2}$ $:=2u_{1}\zeta^{1}+(1-u_{1}^{2}+u_{2}^{2})\zeta^{2}-2u_{1}u_{2}\zeta^{3},$
$\tilde{\zeta}^{3}$ $:=2u_{2}\zeta^{1}-2u_{1}u_{2}\zeta^{2}+(1+u_{1}^{2}-u_{2}^{2})\zeta^{3}.$
Then the complex manifold $\left(X_{\zeta},I_{\zeta}^{u}\right)$ is isomorphic to the complex manifold $\left(X_{\tilde{\zeta}},I_{\tilde{\zeta}}\right)$, where $\tilde{\zeta}=\left(\tilde{\zeta}^{1},\tilde{\zeta}^{2},\tilde{\zeta}^{3}\right)$ for any $\tilde{\zeta}^{1}\in Z$ such that $\tilde{\zeta}\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$.
Proof.
For the complex manifold $\left(X_{\zeta},I_{\zeta}^{u}\right)$, there is a holomorphic two form $\Omega(u):=2u\,\omega_{\zeta}^{1}+\left(1-u^{2}\right)\omega_{\zeta}^{2}+\mathrm{i}\left(1+u^{2}\right)\omega_{\zeta}^{3}$. By the isomorphism in Lemma 2.15, the period $[\Omega(u)]$ of $\left(X_{\zeta},I_{\zeta}^{u}\right)$ can also be given by an element in $Z\otimes\mathbb{C}$ as follows:
$\sigma_{\mathbb{C}}^{-1}\,[\Omega(u)]=\zeta^{2}+\mathrm{i}\zeta^{3}+2u\zeta^{1}-u^{2}\left(\zeta^{2}-\mathrm{i}\zeta^{3}\right).$
By the definitions of $\tilde{\zeta}^{2}$ and $\tilde{\zeta}^{3}$, the period of $(X_{\tilde{\zeta}},I_{\tilde{\zeta}})$ in $Z\otimes\mathbb{C}$ is given by $\tilde{\zeta}^{2}+\mathrm{i}\tilde{\zeta}^{2}$ which also equals to $\zeta^{2}+\mathrm{i}\zeta^{3}+2u\zeta^{1}-u^{2}\left(\zeta^{2}-\mathrm{i}\zeta^{3}\right)$. Hence $\left(X_{\zeta},I_{\zeta}^{u}\right)$ and $(X_{\tilde{\zeta}},I_{\tilde{\zeta}})$ have the same period. Denote by $u^{\prime}\in\mathbb{P}^{1}$ the corresponding point of $(1,0,0)\in\mathbb{S}^{2}$, and $\tilde{u}^{\prime}\in\mathbb{C}^{2}$ any lifting of $u^{\prime}$. Then,
$\widetilde{p_{\zeta}}(\tilde{u})=\widetilde{p_{\tilde{\zeta}}}(\tilde{u}^{\prime})=:\lambda_{0}.$
As in the proof of Corollary 2.14, $\left(X_{\zeta},I_{\zeta}^{u}\right)$ is the minimal resolution of the fiber $\tilde{f}^{-1}(\lambda_{0})$ of $\tilde{f}$. Similarly, $(X_{\tilde{\zeta}},I_{\tilde{\zeta}})$ is also the minimal resolution of the fiber $\tilde{f}^{-1}(\lambda_{0})$ of $\tilde{f}$. Since the minimal resolution of a normal surface is unique, see Lemma 2.8, they are isomorphic to each other.
Remark 2.17.
When $u=\infty$, $I_{\zeta}^{u}=-I_{\zeta}$ and the complex manifold $\left(X_{\zeta},-I_{\zeta}\right)$ is isomorphic to $\left(X_{\tilde{\zeta}},I_{\tilde{\zeta}}\right)$, where $\tilde{\zeta}:=\left(\zeta^{1},-\zeta^{2},\zeta^{3}\right)$.
3. HyperKähler ALEs as minimal resolution of Kleinian singularities
This section initiates the proofs of our main theorems. Specifically:
  • Subsection 3.1 establishes Theorem 1.1;
  • Subsection 3.2 addresses case (i) of Theorem 1.2;
  • The remaining cases of Theorem 1.2 are deferred to Section 4.
3.1. Proof of Theorem 1.1:
The core idea involves comparing the periods of $I_{\mathrm{std}}$ and $I_{\zeta}$.
Proof of the ”if” part. Suppose $I_{\mathrm{std}}\in\{aI_{\zeta}+bJ_{\zeta}+cK_{\zeta}\mid(a,b,c)\in\mathbb{S}^{2}\}$ for some $\zeta=(\zeta^{1},\zeta^{2},\zeta^{3})\in(Z\otimes\mathbb{R}^{3})^{\circ}$. First assume $I_{\mathrm{std}}\simeq I_{\zeta}$, so that $(X_{\zeta},I_{\zeta})\simeq\widetilde{\mathbb{C}^{2}/\Gamma}$. If $(\zeta^{2},\zeta^{3})\neq 0$, Lemma 2.15 implies that the cohomology class of the holomorphic 2-form $\omega_{\mathbb{C}}=\omega_{2}+\mathrm{i}\omega_{3}$ on $(X_{\zeta},I_{\zeta})$ is non-zero:
$[\omega_{\mathbb{C}}]=[\omega_{2}]+\mathrm{i}[\omega_{3}]=\sigma(\zeta^{2}+\mathrm{i}\zeta^{3})\neq 0.$
Since $H_{2}(X_{\zeta},\mathbb{Z})$ is generated by the exceptional 2-spheres $\{\Sigma_{i}\}_{i=1}^{r}$, there exists some $[\Sigma_{i_{0}}]\in H_{2}(X_{\zeta},\mathbb{Z})$ satisfying $\langle\omega_{\mathbb{C}},[\Sigma_{i_{0}}]\rangle\neq 0$. Consequently, $\Sigma_{i_{0}}$ cannot be holomorphic in $(X_{\zeta},I_{\zeta})$, a contradiction. Therefore, when $(X_{\zeta},I_{\zeta})\simeq\widetilde{\mathbb{C}^{2}/\Gamma}$, we must have $\zeta=(\zeta^{1},0,0)$. Kronheimer’s result confirms that $\zeta=(\zeta^{1},0,0)\in(Z\otimes\mathbb{R}^{3})^{\circ}$ implies $(X_{\zeta},I_{\zeta})\simeq\widetilde{\mathbb{C}^{2}/\Gamma}$. Thus the equivalence holds precisely for such $\zeta$.
Now assume $I_{\mathrm{std}}\simeq I_{\zeta}^{u}$ for some $u\in\mathbb{C}\subset\mathbb{P}^{1}$. Choose $\widetilde{\zeta}=(\widetilde{\zeta}^{1},\widetilde{\zeta}^{2},\widetilde{\zeta}^{3})$ as defined in Lemma 2.16, so that $(X_{\zeta},I_{\zeta}^{u})\simeq(X_{\widetilde{\zeta}},I_{\widetilde{\zeta}})$. Since $I_{\mathrm{std}}\simeq I_{\zeta}^{u}$, we have $(X_{\widetilde{\zeta}},I_{\widetilde{\zeta}})\simeq\widetilde{\mathbb{C}^{2}/\Gamma}$. The preceding argument implies $\widetilde{\zeta}^{2}=\widetilde{\zeta}^{3}=0$, leading to the system:
(3.1) $\begin{cases}2u_{1}\zeta^{1}+(1-u_{1}^{2}+u_{2}^{2})\zeta^{2}-2u_{1}u_{2}\zeta^{3}=0\\ 2u_{2}\zeta^{1}-2u_{1}u_{2}\zeta^{2}+(1+u_{1}^{2}-u_{2}^{2})\zeta^{3}=0.\end{cases}$
These equations force $\mathrm{rank}(\zeta)=1$.
Proof of the ”only if” part. Given $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$ with $\mathrm{rank}(\zeta)=1$, we may write $\zeta=\zeta^{1}(x,y,z)$ where $\zeta^{1}\in Z\setminus\bigcup_{\theta\in\Phi}D_{\theta}$ and $(x,y,z)\in\mathbb{R}^{3}\setminus\{0\}$. To find $u$ such that $I_{\mathrm{std}}\simeq I_{\zeta}^{u}$, we solve system (3.1), which holds iff the period of $(X_{\zeta},I_{\zeta}^{u})$ vanishes. Lifting $u$ to $(z_{1},z_{2})\in\mathbb{C}^{2}$, we solve:
$z_{1}^{2}(y\zeta^{1}+\mathrm{i}z\zeta^{1})+2z_{1}z_{2}x\zeta^{1}-z_{2}^{2}(y\zeta^{1}-\mathrm{i}z\zeta^{1})=0.$
Since $\zeta^{1}\neq 0$, this reduces to finding non-trivial solutions to:
$z_{1}^{2}(y+\mathrm{i}z)+2z_{1}z_{2}x-z_{2}^{2}(y-\mathrm{i}z)=0.$
Such solutions exist iff there are $(w_{1},w_{2},w_{3})\in\mathbb{C}^{3}\setminus\{0\}$ with $w_{1}w_{3}=w_{2}^{2}$ satisfying:
(3.2) $w_{1}(y+\mathrm{i}z)+w_{2}(2x)+w_{3}(-y+\mathrm{i}z)=0.$
Projectively, this requires the line (3.2) in $\mathbb{P}^{2}$ to intersect the conic $w_{1}w_{3}=w_{2}^{2}$. Since this intersection is non-empty, solutions always exist. This completes the proof of Theorem 1.1(i).
We now complete the proof of Theorem 1.1.
First, note that when $I_{\zeta}^{u}\simeq I_{\mathrm{std}}$, the period of $(X_{\zeta},I^{u}_{\zeta})$ vanishes. This follows from Lemma 2.6, since the exceptional locus has type $\Phi$.
Given $\mathrm{rank}(\zeta)=1$, the period map $p_{\zeta}$ has image in a one-dimensional subspace of $\mathfrak{h}^{\mathbb{C}}$. By [11, p.693-p.694], this image cannot be contained in the branch locus $\mathcal{D}$. Since $\mathcal{D}$ is a finite union of hyperplanes (Lemma 2.4), $\mathrm{Im}\,p_{\zeta}$ intersects $\mathcal{D}$ only at the origin. Thus for any $s\in\mathrm{Im}\,p_{\zeta}\setminus\{0\}$, the fiber $\phi^{-1}(s)$ is smooth by Lemma 2.13.
The $\mathbb{C}^{*}$-equivariance of $\phi$ implies that $\phi^{-1}(s)$ and $\phi^{-1}(\gamma s)$ are biholomorphic for $\gamma\in\mathbb{C}^{*}$. Consequently, all complex structures in the hyperKähler family distinct from $I_{\mathrm{std}}$ are equivalent. Moreover, Lemma 2.5 ensures that each $\phi^{-1}(s)$ is a smooth affine hypersurface in $\mathbb{C}^{3}$.
Therefore, if $(X_{\zeta},I^{u}_{\zeta})$ contains rational curves, then necessarily $I_{\zeta}^{u}\simeq I_{\mathrm{std}}$. The precise count of such complex structures (which equals 2) will be established in the next subsection. ∎
3.2. Proof of Theorem 1.2(i)
By Theorem 1.1, the cardinality of $\{u\in\mathbb{P}^{1}\mid(X_{\zeta},I_{\zeta}^{u})\text{ contains rational curves}\}$ equals that of $\{u\in\mathbb{P}^{1}\mid I_{\zeta}^{u}\simeq I_{\mathrm{std}}\}$. As established in the proof of Theorem 1.1, this corresponds to the number of intersection points between the projective line (3.2) and the conic $w_{1}w_{3}=w_{2}^{2}$.
Elementary algebra shows that tangency occurs iff:
$x^{2}+|y+\mathrm{i}z|^{2}=0,$
which implies $(x,y,z)=(0,0,0)$. This would contradict $\mathrm{rank}(\zeta)=1$ since $\zeta=\zeta^{1}(0,0,0)=0$. Therefore, the line and conic intersect transversely, yielding exactly two intersection points. Consequently:
(3.3) #$\#\left\{u\in\mathbb{P}^{1}:\pi^{-1}(u)\text{ contains a rational curve}\right\}=2.$
Equation (1.3) follows immediately from the $r$ rational curves in $\widetilde{\mathbb{C}^{2}/\Gamma}$:
(3.4) $\sum_{u\in\mathbb{P}^{1}}\#\left\{\text{rational curves in }\pi^{-1}(u)\right\}=2r.$
When $\Gamma=\mathbb{Z}_{2}$, we have $\dim Z=1$, so $\mathrm{rank}(\zeta)=1$ for all $\zeta\in(Z\otimes\mathbb{R}^{3})^{\circ}$. Theorems 1.1 and 1.2(i) imply:
Corollary 3.1.
In the hyperKähler family of any ALE space asymptotic to $\mathbb{C}^{2}/\mathbb{Z}_{2}$, exactly two complex structures are isomorphic to the minimal resolution $\widetilde{\mathbb{C}^{2}/\mathbb{Z}_{2}}$.
Remark 3.2.
Further properties of the twistor space are discussed in Example 6.2 (Section 5).
4. Counting rational curves
This section completes the proof of Theorem 1.2. Recall the three cases for the parameter $\zeta$: (i) $\operatorname{rank}(\zeta)=1$; (ii) $\operatorname{rank}(\zeta)=2$; (3) $\operatorname{rank}(\zeta)=3$. Case (i) was established in Section 3. We now prove Cases (ii) and (iii) in Subsections 4.2–4.3, with Corollary 1.6 addressed in Subsection 4.4. As a technical preparation for Case (ii), Subsection 4.1 introduces a decomposition of irreducible root systems and presents key combinatorial results of independent interest.
Proof strategy overview. By Corollary 2.14 and Lemma 2.6, the image of the period map $\widetilde{p_{\zeta}}$ encodes information about:
  1. (1)
    Complex structures containing rational curves
  2. (2)
    The number of such curves
To bound the two cardinalities
(4.1) #$\#\left\{u\in\mathbb{P}^{1}\mid\pi^{-1}(u)\text{ contains a rational curve}\right\}$
(4.2) $\sum_{u\in\mathbb{P}^{1}}\#\left\{\text{rational curves in }\pi^{-1}(u)\right\}$
we analyze the geometric position of $\operatorname{Im}\widetilde{p_{\zeta}}$. This position is determined by its intersection with root kernels, which induces a root system decomposition. The decomposition in Subsection 4.1, designed for Case (ii), also provides a framework for Case (iii). We develop distinct approaches for these cases in subsequent subsections.
4.1. Decomposition of irreducible root systems
For a nontrivial finite subgroup $\Gamma\leq\mathrm{SU}(2)$, we recall from Subsection 2.1 that $\mathfrak{h}^{\mathbb{C}}$, $\Phi$, and $\Delta$ denote respectively the complexified Cartan subalgebra, the root system, and the set of simple roots. Throughout, we assume $\operatorname{rank}\Phi\geq 2$, i.e., $\Gamma\neq\mathbb{Z}_{2}$.
Fix a two-dimensional subspace $L\subset\mathfrak{h}^{\mathbb{C}}$ not contained in any root kernel $\ker\theta$ for $\theta\in\Phi$. This induces for each root $\theta\in\Phi\subset\operatorname{Hom}(\mathfrak{h},\mathbb{R})\hookrightarrow\operatorname{Hom}(\mathfrak{h}^{\mathbb{C}},\mathbb{C})$ a one-dimensional kernel $l_{\theta}:=\ker\theta|_{L}$, yielding the map:
$\mu_{L}\colon\Phi$ $\to\mathbb{P}(L)$
$\theta$ $\mapsto[l_{\theta}].$
The image $\operatorname{Im}\mu_{L}=\{p_{1},\dots,p_{s}\}$ is finite, inducing a disjoint decomposition:
(4.3) $\Phi=\bigsqcup_{k=1}^{s}\Phi_{k},\quad\Delta=\bigsqcup_{k=1}^{s}\Delta_{k},$
where $\Phi_{k}:=\mu_{L}^{-1}(p_{k})$ and $\Delta_{k}:=\Phi_{k}\cap\Delta$. Each $\Phi_{k}$ is a root subsystem satisfying $\Phi_{k}=\Phi_{k}\cap\,\operatorname{span}_{\mathbb{R}}\Phi_{k}$. Crucially, this decomposition is generally non-orthogonal and $\Delta_{k}$ may not be a base of $\Phi_{k}$. Since $\bigcap_{\theta\in\Phi}\ker\theta=\{0\}\subset\mathfrak{h}$, we have $\bigcap_{\theta\in\Phi}l_{\theta}=\{0\}\subset L$, implying $s\geq 2$. Moreover:
Lemma 4.1.
$3\leq s\leq\frac{1}{2}|\Phi|$.
Proof.
The upper bound $s\leq\frac{1}{2}|\Phi|$ is immediate. For the lower bound, suppose toward contradiction that $s=2$. Then $\operatorname{Im}\mu_{L}=\{[l_{\theta_{1}}],[l_{\theta_{2}}]\}$ for distinct roots $\theta_{1},\theta_{2}$, inducing:
$\Phi=\Phi_{1}\sqcup\Phi_{2},\quad\Delta=\Delta_{1}\sqcup\Delta_{2}.$
Both $\Delta_{1}$ and $\Delta_{2}$ are non-empty: If $\Delta_{1}=\emptyset$, then $\Delta_{2}=\Delta$ implies $\Phi_{2}=\Phi$, forcing $\Phi_{1}=\emptyset$. Since $\Phi$ is irreducible, there exist $\alpha_{1}\in\Delta_{1}$ and $\alpha_{2}\in\Delta_{2}$ with $(\alpha_{1},\alpha_{2})\neq 0$ in the ambient Euclidean space. Thus $\alpha_{1}\pm\alpha_{2}\in\Phi$ for some choice of sign.
Assume $\beta:=\alpha_{1}+\alpha_{2}\in\Phi$. The kernel condition gives:
$l_{\beta}$ $=\ker(\alpha_{1}+\alpha_{2})|_{L}$
$=\ker\alpha_{1}|_{L}\cap\ker\alpha_{2}|_{L}$
$=l_{\alpha_{1}}\cap l_{\alpha_{2}}.$
Since $\mu_{L}(\alpha_{1})=[l_{\theta_{1}}]\neq[l_{\theta_{2}}]=\mu_{L}(\alpha_{2})$, we have $l_{\alpha_{1}}\neq l_{\alpha_{2}}$. Thus $\dim(l_{\alpha_{1}}\cap l_{\alpha_{2}})<\dim l_{\alpha_{1}}=1$, implying $l_{\beta}=\{0\}$. But $\beta\in\Phi$ requires $\ker\beta|_{L}=l_{\beta}$ to be one-dimensional, contradiction.
Therefore $s\geq 3$ as required. ∎
Now we define $r:=\operatorname{rank}\Phi$, $r_{k}:=\operatorname{rank}\Phi_{k}$, and $\delta_{k}:=|\Delta_{k}|$. Note that $\delta_{k}\leq r_{k}$ and $r\leq\sum_{k=1}^{s}r_{k}$. Moreover:
Lemma 4.2.
(4.4) $r+1\leq\sum_{k=1}^{s}r_{k}\leq\frac{1}{2}|\Phi|.$
Proof.
The upper bound follows since finer decompositions increase the rank sum. For the lower bound, suppose $r=\sum_{k=1}^{s}r_{k}$. As $r=\sum_{k=1}^{s}\delta_{k}$ and $\delta_{k}\leq r_{k}$, we have $r_{k}=\delta_{k}$ for each $k$. Thus $\Delta_{k}$ forms a base for $\Phi_{k}$. Let $\theta_{\max}$ be the maximal root of $\Phi$ relative to $\Delta$, with expression $\theta_{\max}=\sum_{\alpha\in\Delta}k_{\alpha}\alpha$ ($k_{\alpha}>0$). Since each $\Delta_{k}$ is a proper subset of $\Delta$ (as $s\geq 3$ by Lemma 4.1), $\theta_{\max}$ involves roots from multiple $\Delta_{k}$. Hence $\theta_{\max}\notin\Phi_{k}$ for any $k$, contradicting $\Phi=\bigsqcup_{k}\Phi_{k}$. ∎
To achieve the stronger lower bound $12r-$ 1 (improving Lemma 4.2), we introduce a refined class of root system decompositions:
Definition 4.3.
A decomposition $\Phi=\bigsqcup_{k=1}^{s}\Phi_{k}$ is called type-1 if:
  1. (1)
    $s\geq 1$,
  2. (2)
    Each $\Phi_{k}$ is a non-empty proper root subsystem satisfying $\Phi_{k}=\Phi\cap\operatorname{span}_{\mathbb{R}}\Phi_{k}$,
  3. (3)
    The subsystems are pairwise disjoint.
The space of all type-1 decompositions is denoted $\mathfrak{D}^{1}(\Phi)$.
Remark 4.4.
While the $L$-induced decomposition (4.3) is type-1, $\mathfrak{D}^{1}(\Phi)$ contains decompositions not arising from any subspace $L$. Thus $\mathfrak{D}^{1}(\Phi)$ provides a coarser classification than the $L$-induced decomposition space.
For any type-1 decomposition $\Phi=\bigsqcup_{k=1}^{s}\Phi_{k}$, define $r_{k}:=\operatorname{rank}\Phi_{k}$. The minimal rank sum $f_{1}(\Phi)$ is defined as:
$f_{1}(\Phi):=\inf_{\mathfrak{D}^{1}(\Phi)}\left\{\sum_{k=1}^{s}r_{k}\right\}.$
By Lemma 4.2, whose proof depends only on $\Phi$’s irreducibility and properties (1)–(2), we have $f_{1}(\Phi)\geq r+1$. A sharper lower bound holds for ADE-type root systems:
Proposition 4.5.
Let $\Phi(n)$ denote an irreducible ADE root system of rank $n\geq 2$, i.e., $\Phi(n)=A_{n}$ ($n\geq 2$), $D_{n}$ ($n\geq 3$), or $E_{n}$ ($n=6,7,8$). Then:
$2n-1\leq f_{1}(\Phi(n)).$
This is proved by induction on $n$. We first establish a key lemma. Consider $\Phi(n+1)$ (an ADE root system of rank $n+1$) embedded in a certain Euclidean space $\mathrm{E}^{n+1}$. Fix an embedding $A_{n}\hookrightarrow\Phi(n+1)$ (to be specified in the proof). The set of extra roots of $\Phi(n+1)$ with respect to $A_{n}$ is defined as $S:=\Phi(n+1)\setminus A_{n}$.
Lemma 4.6.
The extra roots in $S$ span $\mathrm{E}^{n+1}$, that is, $\dim_{\mathbb{R}}\operatorname{span}_{\mathbb{R}}S=n+1$.
Proof.
The proof involves explicitly calculating for each type of root systems.
[…]
    $\alpha_{1}=\epsilon_{1}-\epsilon_{2}$
    $\cdots$
    $\alpha_{i}=\epsilon_{i}-\epsilon_{i+1}$
    $\cdots$
    $\alpha_{n}=\epsilon_{n}-\epsilon_{n+1}$
    $\alpha_{n+1}=\epsilon_{n+1}-\epsilon_{n+2}$
    Then we can write the positive roots as linear combinations of these simple roots as:
[…]
    The first $n$ simple roots $\alpha_{1},\cdots,\alpha_{n}$ determine the embedding of $A_{n}\hookrightarrow A_{n+1}$. The ’extra’ positive roots with respect to $A_{n}$ are: $S^{+}:=\left\{\epsilon_{i}-\epsilon_{n+2},1\leq i\leq n+1\right\}$. By elementary computation, we see:
    $\mathrm{span}_{\mathbb{R}}\,S^{+}=\mathrm{E}.$
    Thus, $\dim\mathrm{span}_{\mathbb{R}}S=n+1$.
[…]
    $\alpha_{1}=\epsilon_{1}-\epsilon_{2}$
    $\cdots$
    $\alpha_{i}=\epsilon_{i}-\epsilon_{i+1}$
    $\cdots$
    $\alpha_{n}=\epsilon_{n}-\epsilon_{n+1}$
    $\alpha_{n+1}=\epsilon_{n}+\epsilon_{n+1}$
    Then we draw the Dynkin diagram of $D_{n+1}$ together with an embedding of $A_{n}$ as:
[…]
    We also write the positive roots as linear combinations of these simple roots:
    $\epsilon_{i}-\epsilon_{j}$ $=\sum_{i<k<j}\alpha_{k}$ $1\leq i<j\leq n+1;$
    $\epsilon_{i}+\epsilon_{j}$ $=\sum_{i\leq k<j}\alpha_{k}+2\left(\sum_{j\leq k<n}\alpha_{k}\right)+\alpha_{n}+\alpha_{n+1}$ $1\leq i<j<n+1;$
    $\epsilon_{i}+\epsilon_{n+1}$ $=\sum_{i\leq k\leq n-1}\alpha_{k}+\alpha_{n+1}$ $1\displaystyle 1\leq i<n+$ 1 .
    The simple roots $\alpha_{1},\cdots,\alpha_{n}$ determine an embedding of $A_{n}\hookrightarrow D_{n+1}.$ The ’extra’ (positive) roots with respect to $A_{n}$ are: $S^{+}:=\left\{\epsilon_{i}+\epsilon_{j},1\leq i<j\leq n+1\right\}$. By elementary computation, we see:
    $\mathrm{span}_{\mathbb{R}}\,S=\mathrm{E}.$
    Thus, $\dim\mathrm{span}_{\mathbb{R}}S=n+1$.
  • For $\Phi(n+1)=E_{6},E_{7},E_{8}$. We only show here the case $\Phi(n+1)=E_{6}$, since the other cases are basically the same. The Dynkin diagram of $E_{6}$ together with an embedding of $A_{5}$ is
[…]
    where we choose a root base $\left\{\alpha_{1},\cdots,\alpha_{6}\right\}$. The simple roots $\alpha_{1},\alpha_{3},\cdots,\alpha_{6}$ determine an embedding of $A_{5}\hookrightarrow E_{6}$. We denote a root $a\alpha_{1}+b\alpha_{2}+c\alpha_{3}+d\alpha_{4}+e\alpha_{5}+f\alpha_{6}$ in the root system $E_{6}$ by a diagram $\begin{pmatrix}&&b&&\\ a&c&d&e&f\end{pmatrix}$. Then we cite from [2] that the following roots are part of the extra roots of $E_{6}$ with respect to $A_{5}$:
    $\begin{pmatrix}&&1&&\\ 0&1&2&1&0\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&1&2&1&0\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 0&1&2&1&1\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&2&2&1&0\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&1&2&1&1\end{pmatrix}$
    $\begin{pmatrix}&&1&&\\ 0&1&2&2&1\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&2&2&1&1\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&1&2&2&1\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&2&2&2&1\end{pmatrix}$ $\begin{pmatrix}&&1&&\\ 1&2&3&2&1\end{pmatrix}$
    $\begin{pmatrix}&&2&&\\ 1&2&3&2&1\end{pmatrix}$
    As an observation, the $\mathbb{R}$-span of these ’extra’ roots of $E_{6}$ with respect to $A_{5}$ contains all the simple roots $\alpha_{1},\cdots,\alpha_{6}$. For example, the difference of the first two roots above is $\alpha_{1}$, hence $\alpha_{1}$ lies in the span of extra roots. Thus, $\dim\mathrm{span}_{\mathbb{R}}S=6$.
Proof of Proposition 4.5: Since $\Phi(n+1)$ is a finite set, the set of all decompositions is also finite. Hence, we can take a decomposition of $\Phi(n+1)$
$\Phi(n+1)=\Phi_{1}\sqcup\cdots\sqcup\Phi_{s}\quad,\quad r_{k}:=\mathrm{rank}\,\Phi_{k}.$
such that $\sum_{k=1}^{s}r_{k}=f_{1}(\Phi(n+1))$. By deleting the extra roots, we obtain the following:
(4.5) $A_{n}=\Phi_{1}^{\prime}\sqcup\cdots\sqcup\Phi_{s}^{\prime}\quad,\quad\Phi_{k}^{\prime}:=\Phi_{k}\setminus S=\Phi_{k}\cap A_{n}\,,\,r_{k}^{\prime}=\mathrm{rank}\,\Phi_{k}^{\prime}.$
By definition, each set $\Phi_{k}^{\prime}$ for $k=1,\cdots,s$ is either empty or a root subsystem of $A_{n}$. Set $s^{\prime}:=\left\{k\mid\Phi_{k}^{\prime}\neq\emptyset\right\}$, then we have two possibilities:
  • $s^{\prime}=1$. Then $\Phi_{k_{0}}=A_{n}$ for some $k_{0}$ and the other $\Phi_{k}$, $k\neq k_{0}$, contains only extra roots. By Lemma 4.6:
    $f_{1}\left(\Phi(n+1)\right)$ $=\sum_{k=1}^{s}r_{k}$
    $=r_{k_{0}}+\sum_{k\neq k_{0}}r_{k}$
    $\geq n+n+1=2n+1.$
  • $s^{\prime}>1$. Then the disjoint union (4.5) is a decomposition of $A_{n}$ of type-1 (ignoring the empty sets). By Lemma 4.6, these extra roots cannot be put into a single root subsystem, hence there exist at least two root subsystems $\Phi_{k_{i}},i=1,2$ such that $r_{k_{i}}\geq r_{k_{i}}^{\prime}+1$. Therefore,
    $f_{1}\left(\Phi(n+1)\right)$ $=\sum_{k=1}^{s}r_{k}$
    $\geq\sum_{k=1}^{s}r_{k}^{\prime}+2$
    $\geq f_{1}\left(A_{n}\right)+2.$
Now we can prove the lower bound $12n-$ 1 for $A_{n}$ in the proposition by induction on $n$:
Lemma 4.2 implies that $f_{1}\left(A_{2}\right)\geq 3=2\cdot 2-1$. And the assumption $f_{1}\left(A_{n}\right)\geq 2n-1$ implies that either $f_{1}\left(A_{n+1}\right)\geq 2n+1$ or $f_{1}\left(A_{n+1}\right)\geq f_{1}\left(A_{n}\right)+2\geq 2n-1+2=2n+1$. Thus, by induction we have $f_{1}\left(A_{n}\right)\geq 2n-1$. And then it is a repeated argument to show $f_{1}\left(\Phi_{n}\right)\geq 2n-1$. ∎
4.2. Proof of Theorem 1.2(ii)
Assume $\operatorname{rank}(\zeta)=2$. Without loss of generality, $\zeta^{1}\in\operatorname{span}\{\zeta^{2},\zeta^{3}\}$. The period map:
$\widetilde{p_{\zeta}}\colon\mathbb{C}^{2}\to\mathfrak{h}^{\mathbb{C}},\quad(z_{1},z_{2})\mapsto z_{1}^{2}(\zeta^{2}+\mathrm{i}\zeta^{3})+2z_{1}z_{2}\zeta^{1}-z_{2}^{2}(\zeta^{2}-\mathrm{i}\zeta^{3})$
factors through $\mathbb{P}^{1}$ via:
$p_{\zeta}^{\prime}\colon\mathbb{P}^{1}$ $\to\mathbb{P}(\operatorname{Im}\widetilde{p_{\zeta}})$
$[z_{1}:z_{2}]$ $\mapsto[w_{1}:w_{2}]:=[z_{1}^{2}+2cz_{1}z_{2}:-z_{2}^{2}+2\bar{c}z_{1}z_{2}]$
where $\zeta^{1}=c(\zeta^{2}+\mathrm{i}\zeta^{3})+\bar{c}(\zeta^{2}-\mathrm{i}\zeta^{3})$ for some $c\in\mathbb{C}$. This is a degree-2 branched covering with exactly two branch points.
Since $\operatorname{rank}(\zeta)=2$, $\operatorname{Im}\widetilde{p_{\zeta}}$ is a 2-dimensional subspace satisfying the $L$-condition from Subsection 4.1 (by [11, p.693-694]). Setting $L:=\operatorname{Im}\widetilde{p_{\zeta}}$, we obtain the decomposition:
$\Phi=\bigsqcup_{k=1}^{s}\Phi_{k},\quad\Delta=\bigsqcup_{k=1}^{s}\Delta_{k}$
with $s\geq 3$ and $2r-1\leq\sum_{k=1}^{s}r_{k}\leq\frac{1}{2}|\Phi|$. Let $\{l_{k}\}_{k=1}^{s}$ be the corresponding lines in $L$.
By Corollary 2.14:
(4.6) #$\#\left\{u\in\mathbb{P}^{1}:\pi^{-1}(u)\text{ contains a rational curve}\right\}=\#(p_{\zeta}^{\prime})^{-1}\big{(}\{[l_{1}],\dots,[l_{s}]\}\big{)}.$
The upper bound $|\Phi|$ is known [11, Lemma 2.7]. For the lower bound, note that each $[l_{k}]$ has two preimages $\{u_{k}^{1},u_{k}^{2}\}\subset\mathbb{P}^{1}$, coinciding iff $[l_{k}]\in\{[1:0],[0:1]\}$. Since $s\geq 3$:
(4.7) #$\#(p_{\zeta}^{\prime})^{-1}\big{(}\{[l_{1}],\dots,[l_{s}]\}\big{)}\geq 4.$
Each $(X_{\zeta},u_{k}^{i})$ ($i=1,2$) is the minimal resolution of $\widetilde{f}^{-1}(v)$ for $v\in l_{k}\setminus\{0\}$ (see Diagram 2.1). By Corollary 2.11, $\widetilde{f}^{-1}(v)$ contains $r_{k}$ rational curves. Proposition 4.5 yields:
$2r\leq\sum_{k=1}^{s}r_{k}\cdot\#\{u_{k}^{1},u_{k}^{2}\}$ $=\sum_{u\in\mathbb{P}^{1}}\#\{\text{rational curves in }\pi^{-1}(u)\}\leq|\Phi|.$
This completes the proof for Case (ii).
4.3. Proof of Theorem 1.2(iii)
Assume $\operatorname{rank}(\zeta)=3$. The proof strategy begins with analyzing the period map’s image. Since $\dim V_{\zeta}=3$ where $V_{\zeta}:=\operatorname{span}_{\mathbb{C}}\{\zeta^{1},\zeta^{2},\zeta^{3}\}\subset\mathfrak{h}^{\mathbb{C}}$, the image $\operatorname{Im}\widetilde{p_{\zeta}}$ is a hypersurface in $V_{\zeta}$.
[…]
$p_{\zeta}^{\prime}\colon\mathbb{P}^{1}$ $\xrightarrow{\sim}C\subset\mathbb{P}(V_{\zeta})\simeq\mathbb{P}^{2}$
$[z_{1}:z_{2}]$ $\mapsto[z_{1}^{2}:z_{1}z_{2}:z_{2}^{2}]$
Here $C=\mathbb{P}(\operatorname{Im}\widetilde{p_{\zeta}})$ is the smooth conic $\{xy=z^{2}\}\subset\mathbb{P}^{2}$.
As $\operatorname{Im}\widetilde{p_{\zeta}}\not\subset\ker\theta$ for any $\theta\in\Phi$, define for each root: $H_{\theta}:=\ker\theta|_{V_{\zeta}}$ (a 2-dimensional subspace); $\mathbb{P}H_{\theta}\simeq\mathbb{P}^{1}$ (projective line in $\mathbb{P}(V_{\zeta})$). Each $\mathbb{P}H_{\theta}$ intersects $C$ either transversely (at two points) or tangentially (at one point). Hence, for any intersection point $\gamma\in C$ we define
Definition 4.7.
An intersection point $\gamma$ is called a tangential point if all $\mathbb{P}H_{\theta}$ through $\gamma$ are tangent to $C$; $\gamma$ is called a transversal point if some $\mathbb{P}H_{\theta}$ intersects $C$ transversely at $\gamma$.
The intersection locus decomposes as:
$\bigcup_{\theta\in\Phi}(\mathbb{P}H_{\theta}\cap C)=\{\gamma_{1},\dots,\gamma_{s},\gamma_{s+1},\dots,\gamma_{s+t}\}$
where $\{\gamma_{k}\}_{k=1}^{s}$ are transversal points and $\{\gamma_{s+j}\}_{j=1}^{t}$ are tangential points. By the bijectivity of $p_{\zeta}^{\prime}$, each $\gamma_{l}$ for $l=1,\cdots,s+t$ corresponds to a unique complex structure $u_{l}\in\mathbb{P}^{1}$. Thus:
(4.8) $s+t=\#\left\{u\in\mathbb{P}^{1}:\pi^{-1}(u)\text{ contains a rational curve}\right\}$
This induces the decomposition:
(4.9) $\Phi=\bigcup_{l=1}^{s+t}\Phi_{l},\quad\Delta=\bigcup_{l=1}^{s+t}\Delta_{l}$
where for $l=1,\dots,s+t$:
$\Phi_{l}:=\{\theta\in\Phi:\gamma_{l}\in\mathbb{P}H_{\theta}\cap C\},\quad\Delta_{l}:=\Phi_{l}\cap\Delta$
Each $\Phi_{l}$ for $l=1,\cdots,s+t$ is a root subsystem satisfying
(4.10) $\Phi\cap\operatorname{span}_{\mathbb{R}}\Phi_{l}=\Phi_{l}.$
Lemma 4.8.
The subsystems $\{\Phi_{s+j}\}_{j=1}^{t}$ are pairwise disjoint, and $\bigcup_{k=1}^{s}\Phi_{k}$ is disjoint from $\bigcup_{j=1}^{t}\Phi_{s+j}$.
Proof.
For any tangential point $\gamma_{s+j}$, every line $\mathbb{P}H_{\theta}$ through it is tangent to $C$. As $C$ is a conic, such lines intersect $C$ only at $\gamma_{s+j}$. Thus:
  • $\Phi_{s+j}\cap\Phi_{s+j^{\prime}}=\emptyset$ for $j\neq j^{\prime}$ (distinct tangential points)
  • $\Phi_{s+j}\cap\Phi_{k}=\emptyset$ (no line through $\gamma_{s+j}$ passes through $\gamma_{k}$)
Denote $r_{k}:=\operatorname{rank}\Phi_{k}$, $\delta_{k}:=|\Delta_{k}|$, $r_{s+j}:=\operatorname{rank}\Phi_{s+j}$, and $\delta_{s+j}:=|\Delta_{s+j}|$. Equation (1.7) then becomes:
(4.11) $\sum_{k=1}^{s}r_{k}+\sum_{j=1}^{t}r_{s+j}=\sum_{u\in\mathbb{P}^{1}}\#\left\{\text{rational curves in }\pi^{-1}(u)\right\}.$
The upper bound follows by refining the decomposition (4.9) into disjoint subsystems: First,
$\Phi_{1}\cup\cdots\cup\Phi_{s}=\left(\bigsqcup_{k=1}^{s}\Phi_{kk}\right)\sqcup\left(\bigsqcup_{1\leq k_{1}<k_{2}\leq s}\Phi_{k_{1}k_{2}}\right)$
where
$\Phi_{kk}$ :=$:=\left\{\theta\in\Phi:\mathbb{P}H_{\theta}\text{ tangent to }C\text{ at }\gamma_{k}\right\}\subset\Phi_{k}$ $(k$ $=1,\dots,s)$
$\Phi_{k_{1}k_{2}}$ :=$:=\left\{\theta\in\Phi:\mathbb{P}H_{\theta}\ni\gamma_{k_{1}},\gamma_{k_{2}}\right\}=\Phi_{k_{1}}\cap\Phi_{k_{2}}$ (1(1( 1 $\leq k_{1}\neq k_{2}\leq s).$
Since $\Phi_{k}=\Phi_{kk}\cup\bigcup_{k^{\prime}\neq k}\Phi_{kk^{\prime}}$, we see $r_{k}\leq r_{kk}+\sum_{k^{\prime}\neq k}r_{kk^{\prime}}$, implying $\sum_{k=1}^{s}r_{k}\leq\sum_{k_{1},k_{2}}r_{k_{1}k_{2}}$. Consequently:
$\sum_{k=1}^{s}r_{k}+\sum_{j=1}^{t}r_{s+j}$ $\leq\sum_{k}r_{kk}+\sum_{k_{1}\neq k_{2}}r_{k_{1}k_{2}}+\sum_{j=1}^{t}r_{s+j}$
$\leq 2\left(\sum_{k=1}^{s}r_{kk}+\sum_{1\leq k_{1}<k_{2}\leq s}r_{k_{1}k_{2}}+\sum_{j=1}^{t}r_{s+j}\right)$
$\leq|\Phi|$
where the last inequality follows from Lemma 4.2 since the union $\bigsqcup_{k=1}^{s}\Phi_{kk}\sqcup\bigsqcup_{1\leq k_{1}<k_{2}\leq s}\Phi_{k_{1}k_{2}}\sqcup\bigsqcup_{j=1}^{t}\Phi_{s+j}$ forms a type-1 decomposition of $\Phi$ (implied by Lemma 4.8).
To establish the lower bound $s+t\geq 3$ for (4.8), we consider three cases:
  1. (1)
    $s=0$ (all tangential points)
  2. (2)
    $t=0$ (all transversal points)
  3. (3)
    $s>0$ and $t>0$ (mixed types)
For lower bound of (4.11), we introduce in Subsection 4.3 a refined decomposition of $\Phi$ analogous to that in Subsection 4.1 to prove $2r-1\leq\sum_{l=1}^{s+t}r_{l}$
4.3.1. Case $s=0$: Impossibility
We prove this case cannot occur.
Lemma 4.9.
In decomposition (4.9), $s\neq 0$.
Proof.
Suppose $s=0$. By Lemma 4.8, the decomposition simplifies to:
$\Phi=\bigsqcup_{j=1}^{t}\Phi_{j},\quad\Delta=\bigsqcup_{j=1}^{t}\Delta_{j}$
where we re-index $\Phi_{0+j}$ as $\Phi_{j}$. Lemma 4.1 implies $t\geq 3$. By irreducibility, we may assume $\Phi_{1}\neq\emptyset$, $\Phi_{2}\neq\emptyset$, and $(\Phi_{1},\Phi_{2})\neq 0$. Thus there exist $\alpha\in\Phi_{1}$, $\beta\in\Phi_{2}$ with $\alpha\pm\beta\in\Phi$. Assume $\gamma:=\alpha+\beta\in\Phi_{3}$ without loss of generality.
The projective intersections satisfy:
$\mathbb{P}H_{\alpha}\cap\mathbb{P}H_{\beta}=\mathbb{P}(H_{\alpha}\cap H_{\beta})=\mathbb{P}H_{\alpha}\cap\mathbb{P}H_{\gamma}=\mathbb{P}H_{\beta}\cap\mathbb{P}H_{\gamma}$
Thus $\mathbb{P}H_{\alpha}$, $\mathbb{P}H_{\beta}$, $\mathbb{P}H_{\gamma}$ are concurrent at some $p\in\mathbb{P}(V_{\zeta})$. By definition, each is tangent to $C$ at distinct points $\gamma_{1},\gamma_{2},\gamma_{3}$. But a conic cannot have three distinct tangent lines concurrent, a contradiction. ∎
4.3.2. Case $t=0$: Transversal points only
Lemma 4.10.
For decomposition (4.9) with $t=0$, we have $s\geq 3$.
Proof.
Assume $t=0$. Since $s$ counts transversal points, minimality implies $s\geq 2$. Suppose toward contradiction $s=2$, so all intersection points are $\{\gamma_{1},\gamma_{2}\}$. The decomposition becomes:
$\Phi$ $=\Phi_{1}\cup\Phi_{2},$
$\Delta$ $=\Delta_{1}\cup\Delta_{2}.$
Refine this into disjoint subsystems:
$\Phi$ $=\Phi^{\prime}_{1}\sqcup\Phi^{\prime}_{2}\sqcup\Phi^{\prime}_{3},$
$\Delta$ $=\Delta^{\prime}_{1}\sqcup\Delta^{\prime}_{2}\sqcup\Delta^{\prime}_{3},$
where
$\Phi^{\prime}_{1}$ :=$:=\{\theta:\mathbb{P}H_{\theta}\text{ tangent to }C\text{ at }\gamma_{1}\}$
$\Phi^{\prime}_{2}$ :=$:=\{\theta:\mathbb{P}H_{\theta}\text{ tangent to }C\text{ at }\gamma_{2}\}$
$\Phi^{\prime}_{3}$ :=$:=\{\theta:\mathbb{P}H_{\theta}\text{ contains }\gamma_{1}\text{ and }\gamma_{2}\}$
These are root subsystems with $\Phi^{\prime}_{3}\neq\emptyset$ and $\Phi^{\prime}_{1},\Phi^{\prime}_{2}\neq\emptyset$ (since $\bigcap_{\theta}H_{\theta}=\{0\}$). By irreducibility, they are not pairwise orthogonal. Without loss of generality, take $\alpha\in\Phi^{\prime}_{1}$, $\beta\in\Phi^{\prime}_{2}$ with $\gamma:=\alpha+\beta\in\Phi^{\prime}_{3}$.
Projectively, there holds
$\mathbb{P}H_{\alpha}\cap\mathbb{P}H_{\gamma}=\mathbb{P}H_{\beta}\cap\mathbb{P}H_{\gamma}=\mathbb{P}H_{\alpha}\cap\mathbb{P}H_{\beta}$
Thus $\mathbb{P}H_{\alpha}$, $\mathbb{P}H_{\beta}$, $\mathbb{P}H_{\gamma}$ are concurrent at some $p\in\mathbb{P}(V_{\zeta})$. But we see
$\mathbb{P}H_{\alpha}\text{ tangent to }C\text{ at }\gamma_{1}$
$\mathbb{P}H_{\beta}\text{ tangent to }C\text{ at }\gamma_{2}$
$\mathbb{P}H_{\gamma}\text{ transverse to }C\text{ at both }\gamma_{1}\text{ and }\gamma_{2}$
This configuration is geometrically impossible in $\mathbb{P}^{2}$. Contradiction. ∎
Thus for $t=0$:
(4.12) $3\leq s=\#\left\{u\in\mathbb{P}^{1}:\pi^{-1}(u)\text{ contains a rational curve}\right\}$
4.3.3. Case $s>0$ and $t>0$: Mixed intersection types
We have $s\geq 2$ and $t\geq 1$, so:
$s+t\geq 3=\#\{u_{k},u_{s+j}:k=1,\dots,s;\,j=1,\dots,t\}=\#\left\{u\in\mathbb{P}^{1}:\pi^{-1}(u)\text{ contains a rational curve}\right\}.$
Moreover:
Lemma 4.11.
For decomposition (4.9) with $s=2$, we have $t\geq 2$.
Proof.
Assume $s=2$. By Lemma 4.10, $t\geq 1$.
Claim: $\Phi_{1}\cup\Phi_{2}$ is a root subsystem. Suppose not. Then there exist $\alpha,\beta\in\Phi_{1}\cup\Phi_{2}$ and an integer $m$ such that $\gamma:=\alpha+m\beta\in\Phi$ but $\gamma\notin\Phi_{1}\cup\Phi_{2}$. Thus $\gamma\in\Phi_{2+j_{0}}$ for some $j_{0}$. Since $\Phi_{1}\cap\Phi_{2}$ is a root subsystem, $\alpha$ and $\beta$ cannot both lie in $\Phi_{1}\cap\Phi_{2}$. By symmetry, consider two cases:
  1. (P1)
    $\alpha\in\Phi_{1}\setminus\Phi_{2}$, $\beta\in\Phi_{2}\setminus\Phi_{1}$. Then $\mathbb{P}H_{\alpha}$, $\mathbb{P}H_{\beta}$, and $\mathbb{P}H_{\gamma}$ are concurrent. But
    $\mathbb{P}H_{\alpha}\text{ tangent to }C\text{ at }\gamma_{1}$
    $\mathbb{P}H_{\beta}\text{ tangent to }C\text{ at }\gamma_{2}$
    $\mathbb{P}H_{\gamma}\text{ tangent to C at }\gamma_{2+j_{0}}(\text{ since}\gamma\in\Phi_{2+j_{0}})$
    Three distinct tangent lines cannot be concurrent in $\mathbb{P}^{2}$—contradiction.
  2. (P2)
    $\alpha\in\Phi_{1}\cap\Phi_{2}$, $\beta\in\Phi_{2}\setminus\Phi_{1}$. Then
    $\mathbb{P}H_{\alpha}\text{ contains both }\gamma_{1}\text{ and }\gamma_{2}(\text{ since }\alpha\in\Phi_{1}\cap\Phi_{2})$
    $\mathbb{P}H_{\beta}\text{ tangent at }\gamma_{2}$
    $\mathbb{P}H_{\gamma}\text{ must contain }\gamma_{2}(\text{ since }\gamma_{2}\text{ is transversal })$
    But $\gamma\in\Phi_{2+j_{0}}$ implies $\mathbb{P}H_{\gamma}$ is tangent at $\gamma_{2+j_{0}}\neq\gamma_{2}$, contradiction.
Thus $\Phi_{1}\cup\Phi_{2}$ is a root subsystem satisfying:
(4.13) $\Phi\cap\operatorname{span}_{\mathbb{R}}(\Phi_{1}\cup\Phi_{2})=\Phi_{1}\cup\Phi_{2}.$
Proof of $t\geq 2$: Suppose $t=1$. Then $\Phi=(\Phi_{1}\cup\Phi_{2})\sqcup\Phi_{3}$. By irreducibility, $(\Phi_{1}\cup\Phi_{2},\Phi_{3})\neq 0$. Take $\alpha\in\Phi_{1}\cup\Phi_{2}$, $\theta\in\Phi_{3}$ with $(\alpha,\theta)\neq 0$. Assume $\eta:=\alpha+\theta\in\Phi$. Conditions (4.10) and (4.13) imply $\eta\notin\Phi_{1}\cup\Phi_{2}\cup\Phi_{3}$, contradiction. ∎
Thus for $\operatorname{rank}(\zeta)=3$:
(4.14) $3\leq\#\left\{u\in\mathbb{P}^{1}:\pi^{-1}(u)\text{ contains a rational curve}\right\}.$
We now establish the lower bound for (1.7) or (4.11):
(4.15) $2r-1\leq\sum_{k=1}^{s}r_{k}+\sum_{j=1}^{t}r_{s+j}\leq\sum_{u\in\mathbb{P}^{1}}\#\left\{\text{rational curves in }\pi^{-1}(u)\right\}.$
To achieve $12r-$ 1, we introduce a refined decomposition class:
Definition 4.12.
A type-2 decomposition of $\Phi$ is an expression
$\Phi=\left(\bigcup_{k=1}^{s}\Phi_{k}\right)\sqcup\left(\bigsqcup_{j=1}^{t}\Phi_{s+j}\right)$
satisfying:
  1. (1)
    $s,t\geq 0$ with $s+t\geq 2$
  2. (2)
    Each $\Phi_{k}$ is a non-empty root subsystem such that $\Phi_{k}=\Phi\cap\operatorname{span}_{\mathbb{R}}\Phi_{k}$.
  3. (3)
    The $\Phi_{s+j}$’s are pairwise disjoint, and $(\bigcup_{k=1}^{s}\Phi_{k})\cap(\bigcup_{j=1}^{t}\Phi_{s+j})=\emptyset$
The space of all type-2 decompositions is denoted by $\mathfrak{D}^{2}(\Phi)$.
For any type-2 decomposition of $\mathfrak{D}^{2}(\Phi)$, define $r_{l}:=\operatorname{rank}\Phi_{l}$ and the minimal rank sum:
$f_{2}(\Phi):=\inf_{\mathfrak{D}^{2}(\Phi)}\left\{\sum_{l=1}^{s+t}r_{l}\right\}.$
Similar to Lemma 4.2, we have $f_{2}(\Phi)\geq r+1$.
Remark 4.13.
The conic-induced decomposition (4.9) is type-2, so $f_{2}(\Phi)$ bounds (4.11) or (1.7). When $s=0$, type-1 decompositions embed into $\mathfrak{D}^{2}(\Phi)$. However, $\mathfrak{D}^{2}(\Phi)$ is too coarse for obtaining the optimal bounds.
Proposition 4.14.
For irreducible ADE root systems $\Phi(n)$ of rank $n\geq 2$:
$2n-1\leq f_{2}(\Phi(n)).$
Proof.
The proof is basically the same as Proposition 4.5. Since the sum $\sum_{l=1}^{s+t}r_{l}$ is a positive integer, the infimum is attained. Fix a minimal type-2 decomposition:
$\Phi(n+1)=\left(\bigcup_{k=1}^{s}\Phi_{k}\right)\sqcup\left(\bigsqcup_{j=1}^{t}\Phi_{s+j}\right).$
Let $S$ be the extra roots in $\Phi(n+1)$ relative to $A_{n}$ (as in Lemma 4.6). By deleting $S$ from $\Phi(n+1)$, we obtain a decomposition for $A_{n}$:
(4.16) $A_{n}=\left(\bigcup_{k=1}^{s}\Phi^{\prime}_{k}\right)\sqcup\left(\bigsqcup_{j=1}^{t}\Phi^{\prime}_{s+j}\right),\quad\Phi^{\prime}_{l}:=\Phi_{l}\setminus S=\Phi_{l}\cap A_{n}.$
Each $\Phi^{\prime}_{l}$ is either empty or a root subsystem. Two cases arise:
Case 1: Some $\Phi^{\prime}_{l_{0}}=A_{n}$. Then:
$f_{2}(\Phi(n+1))\geq\operatorname{rank}A_{n}+\sum_{l\neq l_{0}}\operatorname{rank}\Phi_{l}\geq n+(n+1)=2n+1.$
Case 2: All $\Phi^{\prime}_{l}$ are proper subsystems. Then (4.16) induces a type-2 decomposition of $A_{n}$ (ignoring empty set). Since $S$ spans $\mathbb{E}^{n+1}$:
$f_{2}(\Phi(n+1))\geq f_{2}(A_{n})+2.$
Induction: For $A_{3}$ there holds $f_{2}(A_{3})\geq 3=2\cdot 2-1$. Assume $f_{2}(A_{n})\geq 2n-1$. Then:
$f_{2}(A_{n+1})\geq\min(2n+1,(2n-1)+2)=2n+1.$
Thus $f_{2}(A_{n})\geq 2n-1$. The analogous arguments extend to $D_{n}$ and $E_{n}$. ∎
Therefore:
(4.17) $2r-1\leq f_{2}(\Phi)\leq\sum_{k=1}^{s}r_{k}+\sum_{j=1}^{t}r_{s+j}\leq\sum_{u\in\mathbb{P}^{1}}\#\left\{\text{rational curves in }\pi^{-1}(u)\right\}.$
This completes the proof of Theorem 1.2(iii).
5. The semi-continuity of $Q_{1}$
Recall the map $Q_{1}$ defined in the introduction:
$Q_{1}:\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$ $\to\mathbb{Z}$
$\zeta$$\mapsto\#\left\{u\in\mathbb{P}^{1}\,|\,\pi_{\zeta}^{-1}(u)\mathrm{\,\,contains\,\,a\,\,rational\,\,curve}\right\}$
In this section we prove Theorem 1.4, which states that
$\liminf_{\zeta\to\zeta_{0}}Q_{1}(\zeta)\geq Q_{1}(\zeta_{0})\quad,\,\,\forall\zeta_{0}\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}.$
Proof.
We proceed by three cases with respect to $\mathrm{rank}(\zeta_{0})$. The proof arise from studying how the plane $L_{\zeta}$ and the conic $C_{\zeta}$ behave under perturbation of $\zeta$:
  • (1)
    $\text{rank}(\zeta_{0})=1$. Theorem 1.2 implies that for all $\zeta\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}$, $Q_{1}(\zeta_{0})=2\leq Q_{1}(\zeta)$.
  • (2)
    $\text{rank}(\zeta_{0})=2$. Since rank is a lower semi-continuous function, the set $\{\zeta\in\left(Z\otimes\mathbb{R}^{3}\right)^{\circ}\mid\text{rank}(\zeta)>1\}$ is open. Thus there exists $\delta_{0}>0$ such that for all $\zeta$ with $|\zeta-\zeta_{0}|<\delta_{0}$, we have $\text{rank}(\zeta)\geq 2$.
    • Case: $\text{rank}(\zeta)=2$. From subsection 4.2, there exists two planes $L_{\zeta}$ and $L_{\zeta_{0}}$ in $\mathfrak{h}^{\mathbb{C}}$ inducing two decompositions of the root system $\Phi$:
      $\Phi$ $=\Phi_{1}\sqcup\cdots\sqcup\Phi_{s_{\zeta}}$
      $\Phi$ $=\Phi_{1}^{\prime}\sqcup\cdots\sqcup\Phi_{s_{\zeta_{0}}}^{\prime}$
      which yield
      $Q_{1}(\zeta_{0})$ $=\sum_{k=1}^{s_{\zeta_{0}}}\#\,\{u_{k}^{1},u_{k}^{2}\}$
      $Q_{1}(\zeta)$ $=\sum_{j=1}^{s_{\zeta}}\#\,\{v_{j}^{1},v_{j}^{2}\}$
      First we show that $s_{\zeta_{0}}\leq s_{\zeta}$ for $\zeta$ sufficiently closed to $\zeta_{0}$ : For each $k=1,\cdots,s_{\zeta_{0}}$, select any $\alpha_{k}\in\Phi_{k}$ and denote the unique number $m(k)\in\{1,\cdots,s_{\zeta}\}$ such that $\alpha_{k}\in\Phi^{\prime}_{m(k)}$. We prove by contradiction that the map (depending on choice of $\alpha_{k}$’s) $m:k\mapsto m(k)$ is injective: suppose there exists distinct $(k_{1},k_{2})\in\{1,\cdots,s_{\zeta_{0}}\}\crossproduct\{1,\cdots,s_{\zeta_{0}}\}$ with $m(k_{1})=m(k_{2})$, that is $\begin{cases}(\ker\alpha_{k_{1}})\cap\,L_{\zeta_{0}}\neq(\ker\alpha_{k_{2}})\cap\,L_{\zeta_{0}},\\ (\ker\alpha_{k_{1}})\cap\,L_{\zeta}=(\ker\alpha_{k_{2}})\cap\,L_{\zeta}.\end{cases}$. As $\zeta\to\zeta_{0}$, it’s obvious that $L_{\zeta}\to L_{\zeta_{0}}$, so $(\ker\alpha_{k_{i}})\cap L_{\zeta}\to(\ker\alpha_{k_{i}})\cap L_{\zeta_{0}}$ for $i=1,2$, leading to the contradiction $(\ker\alpha_{k_{1}})\cap L_{\zeta_{0}}=(\ker\alpha_{k_{2}})\cap\,L_{\zeta_{0}}$. Therefore $m:k\mapsto m(k)$ is injective and then $s_{\zeta_{0}}\leq s_{\zeta}$.
[…]
      Since $p^{\prime}_{\zeta(n)}(T(n))=[l_{m(k),\zeta(n)}]$ and $[l_{m(k),\zeta}]\to[l_{k}]$ when $\zeta\to\zeta_{0}$,
      $p^{\prime}_{\zeta_{0}}(t_{0})=\lim_{n\to\infty}p^{\prime}_{\zeta(n)}(t_{n})=\lim_{n\to\infty}[l_{m(k),\zeta(n)}]=[l_{k}].$
      It follows that $[l_{k}]$ is a ramification value of $p^{\prime}_{\zeta_{0}}$, a contradiction. Therefore, we see for each $k$, whether $[l_{k}]$ is a ramification value or not, there always holds
      #$\#\{u_{k}^{1},u_{k}^{2}\}\leq\#\{v_{m(k)}^{1},v_{m(k)}^{2}\}$
      and hence:
      $Q_{1}(\zeta_{0})=\sum_{k}^{s_{\zeta_{0}}}\#\,\{u_{k}^{1},u_{k}^{2}\}\leq\sum_{j=1}^{s_{\zeta}}\#\,\{v_{j}^{1},v_{j}^{2}\}=Q_{1}(\zeta).$
    • Case: $\text{rank}(\zeta)=3$. Let $C_{\zeta}$ denote the projectivization of the image of period map $\tilde{p_{\zeta}}$, which is a conic curve in $\mathbb{P}(\mathfrak{h}^{\mathbb{C}})$, see subsection 4.3. Then as $\zeta\to\zeta_{0}$, $C_{\zeta}\to\mathbb{P}(L_{\zeta_{0}})\subset\mathbb{P}(\mathfrak{h}^{\mathbb{C}})$. We claim that for every $\alpha\in\Phi$, there exists $\epsilon(\alpha)>0$ such that $\mathbb{P}(\ker\alpha)$ is not tangent to $C_{\zeta}$ whenever $|\zeta-\zeta_{0}|<\epsilon(\alpha)$: If not, that is $\exists\alpha_{0}\in\Phi$ such that $\forall\,\epsilon>0$, $\exists\,\zeta(\epsilon)$ with $|\zeta(\epsilon)-\zeta_{0}|<\epsilon$, the $\mathbb{P}(\ker\alpha_{0})$ is tangent to $C_{\zeta(\epsilon)}$. Then as $\epsilon\to 0$, $\zeta(\epsilon)\to\zeta$, $C_{\zeta(\epsilon)}\to\mathbb{P}(L_{\zeta_{0}})$. Hence $\mathbb{P}(\ker\alpha_{0})$ tangents to $\mathbb{P}(L_{\zeta_{0}})$. This implies $\mathbb{P}(L_{\zeta_{0}})=\mathbb{P}(\ker\alpha_{0})$. This contradicts to the fact that the image of period map is not contained in any $\ker\alpha$, see subsection 4.2.
      Hence if we take $\epsilon<\min\{\epsilon(\alpha),\alpha\in\Phi\}$, and $\zeta$ with $|\zeta-\zeta_{0}|<\epsilon$, then for any $\alpha\in\Phi$, $\mathbb{P}(\ker\alpha)$ intersects $C_{\zeta}$ transverse. Therefore:
      $Q_{1}(\zeta)$ $=\text{the number of intersection points of $C_{\zeta}$ with the set $\bigcup_{\alpha\in\Phi}\mathbb{P}(\ker\alpha)$}$
      $\geq 2s_{\zeta_{0}}\geq\sum_{k}^{s_{\zeta_{0}}}\#\,\{u_{k}^{1},u_{k}^{2}\}=Q_{1}(\zeta_{0}).$
  • (3)
    $\text{rank}(\zeta_{0})=3$. Since $\{\zeta\mid\text{rank}(\zeta)>2\}$ is open, $\exists\,\delta_{3}>0$ such that $\forall\zeta$ with $|\zeta-\zeta_{0}|<\delta_{3}$, $\text{rank}(\zeta)=3$. As $\zeta\to\zeta_{0}$, $C_{\zeta}\to C_{\zeta_{0}}$ in $\mathbb{P}(\mathfrak{h}_{\mathbb{C}})$. The transverse points $\gamma_{k}$ of $C_{\zeta_{0}}$ (defined in subsection 4.3) perturbs to transverse points again. Hence, $Q_{1}(\zeta)\geq Q_{1}(\zeta_{0})$.
6. Proof of Corollary 1.6
Hitchin proved that for any hyperKähler ALE with group $\mathbb{Z}_{2}$, its twistor space cannot be Kählerian due to the existence of holomorphic curves fiberwised in twistor space [7]. Hence we need only consider the case where $\Gamma\neq\mathbb{Z}_{2}$. If $\mathrm{Z}(X_{\zeta})$ is Kählerian, then its canonical bundle $K_{\mathrm{Z}}$ would be negative established by Hitchin [7, p.150]. Hence, the line bundle $K_{\mathrm{Z}}$ would be of negative degree when restricted on any compact holomorphic curves in $\mathrm{Z}(X_{\zeta})$. The following lemma that $K_{\mathrm{Z}}\simeq\pi^{*}\mathcal{O}_{\mathbb{P}^{1}}(-4)$ and the property of pullback, imply the line bundle $K_{\mathrm{Z}}$ is trivial when restricted to any fiber of $\pi$. Theorem 1.2 implies that there is always at least one fiber containing a rational curve. The existence of such holomorphic curves leads to the contradiction. Therefore, $\mathrm{Z}(X_{\zeta})$ cannot be Kählerian.∎
Lemma 6.1.
Denote $\mathrm{Z}(M)$ as the twistor space of a hyperKähler manifold $\left(M,g,I,J,K\right)$. Recall the natural holomorphic projection $\pi:\mathrm{Z}(M)\to\mathbb{P}^{1}$. Then, the canonical bundle $K_{\mathrm{Z}(M)}$ of the complex threefold $\mathrm{Z}(M)$ is given by:
$K_{\mathrm{Z}(M)}\simeq\pi^{*}\mathcal{O}_{\mathbb{P}^{1}}(-4).$
Proof.
Denote the three Kähler forms of $\left(M,g,I,J,K\right)$ by $\left(\omega_{I},\omega_{J},\omega_{K}\right)$. The twisted vertical two-form, $\Omega\in H^{0}(\mathrm{Z}(X_{\zeta}),\,\Lambda^{2}T_{F}^{*}\otimes\pi^{*}\mathcal{O}_{\mathbb{P}^{1}}(2))$, as defined in [8], is given by
$\Omega(u)$ $=(\omega_{I}+\mathrm{i}\omega_{J})+u(2\omega_{I})-u^{2}(\omega_{J}-\mathrm{i}\omega_{K}),$
$\Omega(v)$ $=v^{2}(\omega_{I}+\mathrm{i}\omega_{J})+v(2\omega_{I})-(\omega_{J}-\mathrm{i}\omega_{K}),$
[…]
Example 6.2.
For $\Gamma=\mathbb{Z}_{2}$, by Corollary 3.1, the hyperKähler ALE structure is realized on the minimal resolution $\widetilde{\mathbb{C}^{2}/\mathbb{Z}_{2}}$. As complex manifolds, there is an isomorphism
$\widetilde{\mathbb{C}^{2}/\mathbb{Z}_{2}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)\simeq\mathcal{T}^{*}\,\mathbb{P}^{1}.$
The hyperKähler metric on $\mathcal{O}_{\mathbb{P}^{1}}(-2)$ was constructed by Eguchi-Hanson [4] and also by Calabi [3]. Its twistor space $\mathrm{Z}$, as explained in [6], is obtained by resolving the singularities on a singular hypersurface $\tilde{\mathrm{Z}}$ of a four-dimensional complex manifold:
$\tilde{\mathrm{Z}}:=\left\{xy=(z-p_{1})(z-p_{2})\right\}\subset\mathcal{O}_{\mathbb{P}^{1}}(2)\oplus\mathcal{O}_{\mathbb{P}^{1}}(2)\oplus\mathcal{O}_{\mathbb{P}^{1}}(2)$
where $x,y,z$ are parameters of the fibers of three line bundle, and $p_{1},p_{2}$ are some two distinct sections of $\mathcal{O}_{\mathbb{P}^{1}}(2)$.
$\mathbf{Acknowledgements.}$ Both authors express their profound gratitude to Friedrich Knop for his insightful responses to our inquiries on MathOverflow and for generously sharing the invaluable reference [13]. The first author is particularly indebted to Song Sun for the enlightening discussions that substantially enriched this research.